### v1

# ReDiFine: Reusable Diffusion Finetuning for Mitigating Degradation in the Chain of Diffusion

Abstract
Diffusion models have achieved tremendous improvements in generative modeling for images, enabling high-quality generation that is indistinguishable by humans from real images. The qualities of images have reached a threshold at which we can reuse synthetic images for training machine learning models again. This attracts the area as it can relieve the high cost of data collection and fundamentally solve many problems in data-limited areas. In this paper, we focus on a practical scenario in which pretrained text-to-image diffusion models are iteratively finetuned using a set of synthetic images, which we call the Chain of Diffusion. Finetuned models generate images that are used for the next iteration of finetuning. We first demonstrate how these iterative processes result in severe degradation in image qualities. Thorough investigations reveal the most impactful factor for the degradation, and we propose finetuning and generation strategies that can effectively resolve the degradation. Our method, Reusable Diffusion Finetuning (ReDiFine), combines condition drop finetuning and CFG scheduling to maintain the qualities of generated images throughout iterations. ReDiFine works effectively for multiple datasets and models without further hyperparameter search, making synthetic images reusable to finetune future generative models.
1 Introduction
Can state-of-the-art AI models learn from their own outputs and improve themselves? As generative AI (e.g., GPT [1], Diffusion [2, 3]) can now churn out uncountable synthetic texts and images, this question piqued curiosity from many researchers in the past couple of years [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. While some show positive results of self-improving [4, 15], most report an undesirable “model collapse” [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 16, 17, 18]—a phenomenon where a model’s performance degrades when it goes through multiple cycles of training with the self-generated data. When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text that has a lot of repetitions [19], and its linguistic diversity declines rapidly [8, 16]; image models also show quality degradation [7, 10] and loss of diversity [12, 20]. This phenomenon has even been compared to mad cow disease, a fatal neurodegenerative disease caused by feeding cattle with the carcasses of other cattle [12].
Our work examines model collapse in a common use case of generative AI: customizing pretrained text-to-image generative models through finetuning. While previous works [5, 12, 14, 19] suggest including more real images in the training set to avoid model collapse, this is challenging for several reasons. First, the number of artworks generated by a single artist is often limited. Even highly productive artists create at most a few thousands of pieces during their lifetime (e.g., Monet produced 1,983 paintings over 61 years). Secondly, during image collection, it is easy to be deceived by realistic AI-generated images and accidentally include them in the training set. Therefore, we need to understand mode collapse behavior when the proportion of real data is limited and seek solutions beyond simply increasing the number of real images. In this paper, we propose a novel concept of generating “reusable” images, which refers to their potential for finetuning without causing catastrophic model collapse.
(a) Self-Consuming Chain of Diffusion.
(b) Image degradation with the Chain of Diffusion.
Figure 1: (a) Overall pipeline of the Chain of Diffusion. Given a pretrained text-to-image diffusion model $M_{0}$ and a prompt set $P$, a finetuned model $M_{k}$ is trained using $D_{k-1}$ generated at the previous iteration $k-1$. Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$. Chain of diffusion begins with the original dataset $D_{0}$. (b) Image degradation universally occurs during Chain of Diffusion across four datasets. As the Chain of Diffusion progresses, the severity of image degradation intensifies, exhibiting universal patterns of highly saturated images containing repetitive high-frequency patterns. This consistently holds across four datasets and $10$ scenarios that we comprehensively investigate in Section 3.2.
Our goal in this paper is two-fold: (1) to conduct an in-depth empirical analysis of model collapse behavior in the self-consuming chain of diffusion finetuning (referred to as Chain of Diffusion)—a setting already widely used by artists—and to examine which factors in finetuning and sampling affect the reusability of diffusion-generated images; (2) to develop a strategy that can improve the reusability of synthetic images. Our contributions toward these goals are summarized below:
  • We consider a setting widely used for style transfer in diffusion models: finetuning Stable Diffusion [3] with low-rank adaptation (LoRA) finetuning [21] using a moderately-sized training set($\sim$1000100010001000 images). We examine the evolution of images in the Chain of Diffusion across four datasets: two digital art datasets (Pokemon [22] and Kumapi [23]) and two natural image datasets (Butterfly [24] and CelebA-1k [25]). Through extensive analysis, we show that model collapse universally occurs in all datasets regardless of training set size, real image mixing (up to 90%), and a wide range of scenarios with different configurations (see Figure 11(b) and Table 1).
  • We discover that among all factors in the Chain of Diffusion, only one substantially impacts the speed of model collapse: classifier-free guidance (CFG) scale. With low CFG, we observe low-frequency degradation where images gradually become unrecognizably blurry. With high CFG scales, high-frequency degradation occurs, where some features of the images become amplified beyond the normal range with disturbing color saturation. There exists a medium CFG range that can significantly slow down model collapse (see Figure 22(b)). To the best of our knowledge, we are the first to demonstrate such a role of CFG in the Chain of Diffusion.
  • Further, we reveal that the default CFG scale $7.5$ widely used for Stable Diffusion performs very poorly in terms of reusability metric (defined in Eqn (4)) despite its excellent image quality in the first iteration (measured by FID). This highlights the trade-off between first-iteration image quality and future reusability (see Figure 23).
  • We develop a new finetuning strategy, Reusable Diffusion Finetuning (ReDiFine), to mitigate degradation inspired by our analysis of how CFG modulates the model collapse pattern in latent diffusion models. We combine condition drop (for finetuning) and CFG scheduling (for generation) in ReDiFine. While an optimal CFG scale can sufficiently slow down model collapse, it is impractical to sweep through CFG scales over multiple iterations for each dataset to find an optimal trade-off. With our proposed ReDiFine approach, we show that even the default CFG scale (CFG=7.5) can generate reusable images across all four datasets without the need for parameter tuning (see Figure 5 and 6).
2 Related Work
The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [11], has been observed in both language and image generative models. Empirical studies on LLMs [8, 15, 16] reveal that linguistic diversity collapses, especially in high-entropy tasks [8], although this can be mitigated with data accumulation [15]. In image generation, several works [5, 7, 10, 12, 18, 20] note image degradation when diffusion models are recursively trained with self-generated data. The authors of [5] and [12] provide theoretical insights into diversity reduction, demonstrating that the covariance in a toy Gaussian model converges to zero as iterations increase in the self-consuming training loop. Among theoretical studies on this topic [13, 14, 17, 19], [17] is particularly relevant. They use dynamical system tools to show that generative distributions either collapse to a small set of outputs or become uniform over a large set, depending on the temperature setting. This finding is analogous to our observations with modulating CFG scales. Connecting our empirical analysis on finetuning Stable Diffusion with their theoretical framework could be an interesting future direction.
Regarding mitigation strategies, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [5, 12, 14] or accumulating more data [15]. The only mitigation strategy beyond altering the training data composition is proposed by [6], who suggests a self-correcting self-consuming loop using an expert model to correct synthetic outputs. While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not always be feasible for finetuning customized image generative models. In our work, we explore an alternative solution through reusable image generation. By making small improvements in finetuning and sampling steps, we can substantially enhance the reusability of images, allowing them to be used for many more iterations before the model starts to deteriorate. In a tangentially related work [9], the authors discuss bias amplification in self-consuming chains, noting minimal amplification over iterations.
3 Model Collapse in Self-Consuming Chain of Diffusion Finetuning
Table 1: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse. Dash (--) in table entries indicates that the default value was used. All experiments were conducted on Pokemon dataset except for the dataset size. For dataset size, we use CelebA since its original dataset is bigger than Pokemon and we can subsample $500$, $1000$, $2000$ images. More details in Appendix A.
Candidates Default 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
Learning Rate 1e-4 1e-5 1e-3 -- -- -- -- -- -- -- -- -- -- -- -- -- -- -- --
Finetune Both -- -- Unet Text -- -- -- -- -- -- -- -- -- -- -- -- -- --
Noise No -- -- -- -- Yes -- -- -- -- -- -- -- -- -- -- -- -- --
Prompt [^1] C, W, S, and L for Combine, Waifu, Short, and Long, respectively. We concatenate BLIP and Waifu captions as default setting, referred as Combine. Short and Long prompts are BLIP captions generated with limitations in the lengths of captions. More details can be found in Appendix C.4. C -- -- -- -- -- W S L -- -- -- -- -- -- -- -- -- --
CLIP skip $2$- -- -- -- -- -- -- -- $1$ $3$- -- -- -- -- -- -- --
Diffusion Steps $30$- -- -- -- -- -- -- -- -- -- $100$- -- -- -- -- -- --
Epochs $100$- -- -- -- -- -- -- -- -- -- -- $50$ $200$- -- -- -- --
Real mixing $0.0$- -- -- -- -- -- -- -- -- -- -- -- -- $0.5$ $0.9$- -- --
Img / Prompt $1$- -- -- -- -- -- -- -- -- -- -- -- -- -- -- $5$- --
Dataset Size $1000$- -- -- -- -- -- -- -- -- -- -- -- -- -- -- -- $500$ $2000$
3.1 Problem setting & Experimental Setup
Chain of Diffusion
We begin with formally defining the self-consuming chain of diffusion finetuning. Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$. $M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a common adaptation case of a pretrained generative model with a personally collected dataset. Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
$M_{k+1}=\text{Finetune}(M_{0},D_{k},P),$ (1)
$D_{k+1}=\text{Generate}(M_{k+1},P).$ (2)
During the Chain of Diffusion, $M_{0}$ and $P$ are fixed across all the iterations. To maintain the same size of training dataset, we generate one image per text prompt for all iterations. The overall pipeline of the Chain of Diffusion is shown in Figure 11(a).
Model and datasets.
We use Stable Diffusion v1.5 [3] as the pretrained model $M_{0}$ and apply LoRA [21] to finetune the Stable Diffusion at each iteration. We build our implementation on [26] and perform experiments on four datasets: Pokemon [22], Kumapi [23], Butterfly [24], and CelebA-1k [25] to investigate various domains including animation, handwriting, and real pictures. During each iteration, we finetune the pretrained Stable Diffusion $M_{0}$ for 100 epochs. More implementation details can be found in Appendix A.1 and A.2.
(a) Reusability-FID trade-off [^3] We do not display scenarios that change the training dataset size, such as Img/Prompt and Dataset Size, as varying sizes result in FID scores on different scales. The related results are presented in the Appendix C, where we observe similar image degradation..
(b) Image degradation across different CFGs.
Figure 2: (a) Reusability-FID trade-off plot (Pokemon) shows that all scenarios other than changing CFG scales show poor reusability, amplifying the FID score $\sim$3x in 6 iterations of the chain. The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the reusability defined in equation (4). For both $\text{FID}_{1}$ and reusability, lower is better. reusability (500 $\sim$ 1100) is about two times larger than $\text{FID}_{1}$ (300 $\sim$ 600), indicating $\text{FID}_{6}$ three times larger than $\text{FID}_{1}$. (b) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chains of Diffusion. CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down. More images in Appendix B.
Evaluation metrics.
We use CLIP score [27] to quantify text-to-image alignment and Frechet Inception Distance (FID) score [28] to measure image fidelity. Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (3)
to evaluate each iteration $k$. SFD overcomes the problem of FID being sensitive to the number of images to compare. Following [29], we use DiNOv2 [30] as a feature extractor for both FID and SFD since it is more consistent with our visual inspection than Inception-V3 network [31]. With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 features as the FID score.
New evaluation metric: Reusability.
In addition to evaluation metrics for generative models presented above, we propose a new metric to quantify the reusability of images generated in the Chain of Diffusion. We define a reusability metric as the performance discrepancy between the first iteration and the $K$-th iteration in the Chain of Diffusion:
$K\text{-reusability}={\rm FID}_{K}-{\rm FID}_{1},$ (4)
where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set. We use FID as a performance metric here, but this can be CLIP, SFD, or any other performance metric of interest. Note that a low $K$-reusability indicates more reusable images since it means that the model does not degrade much in $K$ iterations. In the rest of the paper, $K=6$, and we simply call it reusability.
(a) t-SNE of images.
(b) CLIP.
(c) FID.
(d) SFD.
Figure 3: (a) t-SNE plot visualizes how the distribution of generated images shifts from the original distribution (black) in the Chain of Diffusion and shows distinct paths for three different CFG scales for Pokemon dataset. Different CFG scales and iterations are differentiated with different colors and transparency. Arrows indicate how the distributions of generated images of chains move for different CFG scales. While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (1.0 and 7.5) deviate fast, indicating image degradation. (b-d) Qualitative comparison for different Chains using CLIP $\uparrow$, FID $\downarrow$, and SFD $\downarrow$ for Pokemon dataset. CFG $2.5$ achieves the most robust performance for all three scores. CFG $1.0$ begins to gradually degrade from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in Figure 22(b). Additional results for different datasets can be found in Appendix D
3.2 Model Collapse in the Chain of Diffusion
In this section, we make a series of observations regarding the model collapse behavior in the Chain of Diffusion. We conduct extensive investigations to reveal the most impactful factor in the model collapse and analyze how this factor contributes to the Chain of Diffusion.
Observation 1: Model collapse is universal in the Chain of Diffusion.
We observe significant image degradation in all four datasets (see Figure 11(b)) in the Chain of Diffusion. The quality of generated images begins to clearly deteriorate in the third iteration and it drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations. We investigated a variety of different scenarios (summarized in Table 1) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon. We tested various dataset sizes for $D_{0}$ and $D_{k}$ (Appendix C.1), increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt (Appendix C.2), mixing real images from $D_{0}$ to $D_{k}$ (Appendix C.3), changing the descriptiveness of prompts (Appendix C.4), freezing U-Net or text encoder during finetuning (Appendix C.5), and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip) (Appendix C.6, C.7, C.8, and C.9). We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability (Appendix C.10), and investigated if the degradation happens for larger Stable Diffusion (Appendix C.11). In all settings we tested, image degradation was universally present. We plot the results for Pokemon dataset on Figure 23 where y-axis is reusability and x-axis is the ${\rm FID}_{1}$. While adding noise and mixing $90\%$ real images show the best reusability (the lower, the better), they still exhibit significant degradation (FID score has been doubled in $6$ iterations). More visual inspections and metric plots can be found in Appendix C.
Observation 2: CFG is the only factor that impacts the model collapse.
Throughout all our experiments, classifier-free guidance (CFG) had the biggest impact in the speed of model collapse. CFG scale was first introduced in [32] to modulate the strength between unconditional and conditional scores at each diffusion step as follows:
[…]
High CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images. On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images. For those familiar with temperature sampling [33], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
In Figure 23, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$), which is expected from our understanding of CFG. A more surprising part is that this comes at the cost of a worse reusability metric (increase in ${\rm FID}_{6}-{\rm FID}_{1}$). Also, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and reusability. Similarly, when the CFG scale is too low—below $2.0$—the improvement in reusability plateaus and both ${\rm FID}_{1}$ and reusability begin to decline. Interestingly, there is an optimal region of CFG values (near $2.5$, specific to Pokemon dataset), where we achieve good reusability while maintaining a good quality in the first iteration as well. Figure 3 presents quantitative results for Pokemon dataset to demonstrate how different CFG scales affect the performance in CLIP, FID, and SFD. CFG scale $2.5$ achieves the most robust performance for all metrics.
(a) Probability density function (PDF) of values of latent vectors at the last diffusion step.
(b) Norm of differences between conditional and unconditional scores during diffusion steps.
Figure 4: Latent analysis to investigate how different CFG scales affect the latent of Stable Diffusion (Pokemon dataset). (a) Latent distribution shrinks over iteration for low CFG and expands with high CFG. Larger values in latent vectors are more likely to occur with high CFG, gradually increasing the tail of the distribution. (b) Differences between conditional and unconditional scores increase as the training set is more degraded. Especially, high differences in the later diffusion steps can be a cause of high-frequency degradation.
Observation 3: High CFG scales cause high-frequency degradation and low CFG scales cause low-frequency degradation.
CFG scale does not only affect the speed of model collapse, but also the pattern of model collapse. As shown in Figure 22(b), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to low-frequency degradation. On the other hand, for CFG $7.5$ how images degrade looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated. The t-SNE plot in Figure 33(a) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales. These patterns were consistent in all four datasets and detailed results can be found in Appendix B.
We further analyze this phenomenon by examining the latent space. First, we show a histogram of the final latent vectors before decoding in Figure 44(a). For a CFG of $1.0$, the latent distribution quickly evolves into a Gaussian-like shape, with its variance decreasing over iterations. This behavior aligns with previous work [5, 12, 19], which theoretically predicted that the self-consuming loop would crop out the tails of the distribution, reducing output diversity until it collapses to a single mode. We believe that these narrowing distributions in the latent space lead to blurrier and more homogeneous outputs in the image space. For a CFG of $7.5$, the evolution of the latent distribution is completely opposite, creating longer tails and approaching a more uniform distribution over space. For a CFG scale of $2.5$, which has the best reusability among the three, the latent distribution is better preserved over iterations.
[…]
Implications of our observations.
Through extensive investigations, we gained a comprehensive understanding of CFG scales’ effects on the Chain of Diffusion. Even though a high CFG scale of $7.5$ is a common choice to generate visually appealing images, it significantly sacrifices reusability to achieve slightly better FID. Moreover, it is noteworthy that CLIP score increases for the early three iterations and drops rapidly in the latter when the CFG scale is $7.5$ (Figure 33(b)). This shows that Stable Diffusion tends to generate images favored by CLIP encoder, which can contain artifacts easily amplified by further finetuning. Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [12] and they provide FID plots with varying levels of sampling bias by changing the CFG scale. While they reported a monotonic decrease in reusability as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$. It shows an intriguing trade-off between perceptual quality and reusability. This suggests that if developers of diffusion models are mindful of the reusability metric, can improve future generations of images substantially by carefully choosing CFG.
4 ReDiFine
In the previous section, we showed that choosing a good CFG scale can considerably improve reusability in the Chain of Diffusion. However, such optimal CFG scales vary from dataset to dataset. To mitigate model collapse solely with CFG tuning, we have to do multiple rounds of finetuning and sampling to evaluate the reusability-FID trade-off. This is too computationally burdensome to be a practical solution. In this section, we propose ReDiFine, Reusable Diffusion Finetuning, which can mitigate model collapse without any parameter tuning. ReDiFine combines two ideas—condition drop finetuning and CFG scheduling—and effectively improves the reusability-FID trade-off in all four datasets (comparable to optimal CFG scales).
4.1 Condition drop finetuning.
We propose condition drop finetuning, which randomly drops the text condition during finetuning to update both conditional and unconditional scores. Even though condition drop was suggested in the original CFG paper [32], it is not commonly done during finetuning since the first-iteration image qualities are outstanding without updating unconditional scores. However, the small difference between conditional and unconditional scores can accumulate over iterations and create a large gap as the iteration progresses in the Chain of Diffusion if we do not explicitly finetune the unconditional scores as shown in Figure 44(b). Inspired by this observation, we drop text conditions with probability $0.2$ during finetuning. We illustrate the effect of condition drop with CFG $7.5$ in Figure 44(b). Note that the norm of Diff does not grow as fast even with a high CFG of $7.5$. While this prevents a potential cause of the high-frequency degradation, the Diff value still spikes up in iteration 6. We conjecture that this is because the $D_{5}$ already contains severe artifacts that conditional scores have to grow substantially to fit the data.
4.2 CFG Scheduling.
Based on the analysis showing that a high CFG scale amplifies the discrepancy between the condition and uncondition scores during iterative fine-tuning (see Figure 44(b)), we propose gradually reducing the CFG scale during diffusion steps. This CFG scheduling mechanism mitigates the negative impacts of overemphasizing the conditional score in later stages while preserving the initial conditional information in the early stages. Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as
$s=s_{0}\times e^{-\alpha\times t/T}$ (6)
where $T$ is $30$ and $s_{0}$ is $7.5$. $\alpha$ denotes the degree of decrease and we use $2$ throughout this paper. Comparison to different $\alpha$ and linear decreasing can be found in Appendix E.
Our method gradually decreases the CFG scale during diffusion steps, leveraging the benefits of high CFG scale in the early stage and low CFG scale in the later stage. This adaptive approach aligns with observations from [34] that different steps contribute uniquely to the generation process. In addition, our CFG scheduling takes advantage of using a small CFG scale without the need for extensive hyperparameter searches for each dataset, providing a flexible and robust choice for the Chain of Diffusion.
5 Reusable Image Generation
Figure 5: ReDiFine effectively mitigates image degradation in the Chain of Diffusion. ReDiFine successfully preserves the characteristics and features without further dataset-specific hyperparameter search. Artifacts observed in high-frequency degradation do not exist for all four datasets.
(a) Pokemon.
(b) CelebA-1k.
(c) Kumapi.
(d) Butterfly.
Figure 6: ReDiFine effectively mitigates Reusability-FID trade-off across all four datasets. While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low reusability and FID at the same time (lower is better). While CFG scheduling alone works well for Pokemon dataset, it is still beneficial to have condition drop finetuning for CelebA-1k and Kumapi datasets. Note that the differences in FID are relatively smaller than those in reusability, supporting the necessity to evaluate reusability in the Chain of Diffusion.
In this section, we present generated images and qualitative metrics for ReDiFine. We use the CFG scale $s_{0}=7.5$ for all datasets and experiments except for the ablation study. Note that ReDiFine can be further improved with hyperparameter tuning, e.g., CFG scale, condition drop probability, and CFG scheduling function. However, our focus in this work is to demonstrate that ReDiFine can robustly stabilize iterative finetuning regardless of datasets and hyperparameters. This robustness provides a great advantage when applying our ReDiFine to finetune Stable Diffusion and generate images using any personally collected image sets.
Visual comparison.
Comparing Figure 5 with Figure 11(b), ReDiFine effectively slows down and mitigates the image degradation in the Chain of Diffusion compared to using the default CFG scale $7.5$ across four datasets. In addition, ReDiFine successfully achieves the same level of high-quality image generation in the Chain of Diffusion compared to using the optimal CFG scales for each dataset, which requires extensive hyperparameter tuning. This is demonstrated by visually comparing the generated images in Figure 5 with the ones generated using the optimal CFG scales for Pokemon and CelebA dataset in Figure 22(b). Additional comparisons for Kumapi and Butterfly datasets can be found in Appendix B.
Mitigate Reusability-FID trade-off.
In addition to the visual comparison, we also provide quantitative results in Figure 6. We can see that ReDiFine consistently helps mitigate Reusability-FID trade-off in the Chain of Diffusion across all four datasets. In contrast, a fixed CFG scale that works best on one dataset fails on others. For example, the optimal CFG scale $2.5$ for Pokemon dataset achieves high (poor) ${\rm Reusability}$ for CelebA-1k and Butterfly datasets. Moreover, the optimal CFG scale $1.5$ for CelebA-1k and Butterfly datasets results in the high (poor) ${\rm FID}_{1}$ for Pokemon dataset. In addition to its ability to generalize across different datasets, ReDiFine demonstrates robustness when initialized with different CFG scales. Illustrated in Figure 66(a), ReDiFine achieves comparable performance with CFG scales of $5.0$ and $10.0$ as CFG $=7.5$.
Ablation study.
We provide an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine. As shown in Figure 6, CFG scheduling alone achieves comparable reusability in the Pokemon dataset, but condition drop finetuning is essential for CelebA-1k and Kumapi. Combining these two, ReDiFine consistently achieves the best trade-off between ${\rm Reusability}$ and ${\rm FID}_{1}$ across all four datasets. Additionally, the differences in ${\rm FID}_{1}$ are relatively small—using different CFG scales can generate high-quality images in the first iteration except for very low CFG scales, e.g., CFG = $1.0$. In contrast, the variance of ${\rm Reusability}$ is much larger. This supports the necessity of evaluating our proposed ${\rm Reusability}$ metric for the Chain of Diffusion. Additional results can be found in Appendix E.
More iterations.
In addition to the comparison over 6 iterations, Figure 7 illustrates the evolution of image degradation over 12 iterations. Notably, as fine-tuning iterations increase, high-frequency degradation emerges with the optimal CFG, highlighting the necessity for additional hyperparameter search to sustain image quality across longer diffusion lengths. Conversely, ReDiFine preserves image quality without high-frequency degradation.
Figure 7: Comparison of the optimal CFG scale and ReDiFine on 12 iterations for Pokemon dataset.
(a) Histogram of latent values.
(b) Norm of differences.
(c) Power spectrum of latents.
Figure 8: Comparison of latent analysis at the 6th iteration for ReDiFine and baselines. (a) ReDiFine preserves the latent distribution. (b) Conditional and unconditional score differences do not surge in the $6$-th iteration, showing dramatic improvements from baselines. (c) ReDiFine maintains similar power spectrum to CFG scale $2.5$ in the $6$-th iteration.
Latent analysis.
Figure 8 demonstrates that ReDiFine can successfully alleviate the symptoms we observed in the latent space in Section 3. When we plot latent histograms, we can notice that ReDiFine preserves the original latent distribution at the 6th iteration as compared to other baselines with different CFG values. More importantly, the conditional-unconditional score differences (Diff) are kept at bay all the way until the 6th iteration, as opposed to the baseline CFG $7.5$, showing a severe discrepancy between the unconditional and conditional scores. Additionally, the power spectrum of latents in the $6$th iteration increases or decreases for CFG scales $1.0$ and $7.5$ while ReDiFine and CFG scale $2.5$ preserve density distribution. More analysis and comparison for all iterations are provided in Appendix F.1.
(a) Autocorrelation.
(b) Average power spectra.
Figure 9: The fingerprints of the original training set and synthetic sets show clear differences, and ReDiFine produces fingerprints similar to CFG scale 2.5. (a) Autocorrelation of image fingerprints. Horizontal and vertical lines gradually disappear for CFG scales $1.0$ and $7.5$ while they are maintained for CFG scale $2.5$ and ReDiFine. (b) Average power spectra of images. Central regions are amplified or diminished for CFG scales $1.0$ and $7.5$, demonstrating low and high-frequency degradation.
Frequency fingerprint.
Several works [35, 36] aim to identify fingerprints of synthetic images. High-quality synthetic images from different generative models have clearly distinct fingerprints, showing the potential to be used for synthetic image detection. We analyze fingerprints of synthetic images for different CFG scales and iterations, and compare them to fingerprints of the original training set. Both power spectra and autocorrelation show clear differences between the original training set and synthetic images as shown in Figure 9. Moreover, fingerprints of synthetic images from ReDiFine are similar to those of images from CFG scale $2.5$, while other CFG scales ($1.0$ and $7.5$) make fingerprints different from the first iteration as iterations proceed. Horizontal and vertical lines in autocorrelation gradually disappear and central regions in power spectra vary for further iterations. Also, the varying central regions in power spectra imply that low frequency features increase and decrease for CFG scale $1.0$ and $7.5$, respectively, aligning with visual inspections. Generating images with fingerprints similar to those of the original real images can be an interesting future direction to reduce the degradation in the Chain of Diffusion. Additional analysis with power spectrum density of images is provided in Appendix F.2.
6 Conclusion and Discussion
Motivated by the prominent usage of diffusion models by artists and the trend of the Internet being filled with synthetic data, we investigate how image quality degrades as we feed diffusion-generated synthetic images to finetune another diffusion model. In our self-consuming chain of diffusion finetuning setting, we observe image degradation in all four datasets we tested. We reveal that CFG plays a crucial role in the trade-off between image reusability and its perceptual quality in the first iteration. We propose a novel idea of generating reusable iamges and propose the ReDiFine strategy that combines two small techniques (condition drop finetuning and CFG scheduling) that can be easily implemented in the diffusion finetuning pipeline with no requirement for additional parameter tuning.
We started this paper with a question: can current AI models learn from their own output and improve themselves? This is not only an imminent question as synthetic data rapidly outnumbers human-generated data [37, 38], but also a long-term existential question as it relates to reaching AI singularity—the point where AI learns from its own output to continuously enhance its intelligence and far surpasses human intelligence. Our paper shows a glimpse that widely-used text-to-image models are not ready to improve from their own creation quite yet. However, we believe that there are many open questions on how we develop AI in a world dominated by AI data. We present one angle by emphasizing the importance of generating reusable data for future AI developers. Another pathway could be developing a new learning algorithm that can discern real and synthetic data and take a different training technique for each.
Supplementary Material
Appendix A Experimental setup
A.1 Hyperparameters
We use Stable Diffusion v1.5 [3] as a pretrained text-to-image diffusion model and apply LoRA [21] to fine-tune it at each iteration. We utilize ft-MSE [39] as a fixed VAE to project images into latent space. We augment images with horizontal flip, and no additional augmentation is used. Our implementation is based on [26], with the environment built on PyTorch v2.2.2 [40] and its accompanying torchvision 0.17.2, utilizing CUDA 12.4 on NVIDIA A-100 and L40S GPUs. All default hyperparameters used throughout our paper can be found in Table 2.
Table 2: Default hyperparameters used for the Chain of Diffusion.
Hyperparameter Value
Optimizer AdamW
Learning Rate - Unet $0.0001$
Learning Rate - CLIP $0.00005$
LoRA Weight Scaling $8$
LoRA Rank $32$
Batch Size $6$
Max Epochs $100$
CLIP Skip $2$
Noise Offset $0.0$
Mixed Precision fp16
Loss Function MSE
Min SNR gamma $5.0$
Max Gradient Norm Clipping $1.0$
Caption Dropout Rate $0.0$
Sampler Euler A
Classifier-Free Guidance Scale $7.5$
Number of Diffusion Steps $30$
Number of Images per Prompt $1$
A.2 Datasets
We use four image datasets to demonstrate that the degradation occurs universally: Pokemon [22], Kumapi [23], Butterfly [24], and CelebA-1k [25] to investigate various domains including animation, handwriting, and real pictures. All images are resized to $512\times 512$ pixels. BLIP captioner [41] and Waifu Diffusion v1.4 tagger [42] are used to generate text prompts. Samples images and prompts can be found in Table 3.
A.2.1 Pokemon
The Pokemon dataset [22] contains $1008$ images corresponding to the Pokemon indexing from No.1 to No.1008. Prompts are constructed by concatenating captions from Waifu Diffusion tagger and BLIP captioner. We restrict the length of BLIP captions between $50$ and $75$.
Table 3: Samples images and prompts for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets.
Pokemon
a green pokemon with red eyes and a leaf on the back of its head and tail, an image of the pokemon character with a red eye and big green tail, all set up to look like it is holding a lea, ultra-detailed, high-definition, high quality, masterpiece, sugimori ken (style), solo, smile, open mouth, simple background, red eyes, white background, standing, full body, pokemon (creature), no humans, fangs, transparent background, claws, Bulbasaur [Uncaptioned image]
a very cute looking pokemon with a big leaf on its back and a big leaf on its head, a very cute little pokemon character with leaves in the back ground around his chest and hea, white background, ultra-detailed, high-definition, high quality, masterpiece, sugimori ken (style), solo, red eyes, closed mouth, standing, full body, pokemon (creature), no humans, fangs, transparent background, claws, outline, white outline, animal focus, fangs out [Uncaptioned image]
CelebA-1k
a woman with brown hair smiling and posing for a picture in front of a mirror and gold and white stripes [Uncaptioned image]
a woman with a very long red hair smiles and laughs on a city street and other people in the background [Uncaptioned image]
Kumapi
solo, simple background, food, donut, grey background, no humans, food focus, still life, Kumapi style [Uncaptioned image]
solo, looking at viewer, cute yellow figure, two tiny hands and feet, simple background, black dot eyes, white background, grey background, no humans, Kumapi style [Uncaptioned image]
Butterfly
a crimson patched longwing butterfly with a red and black stripe on its wings, wings are long, narrow, rounded, black, crossed on fore wing by broad crimson patch, and on hind wing by narrow yellow line [Uncaptioned image]
a Common Buckeye butterfly is sitting on a flower in the sun, wings scalloped and rounded except at drawn-out fore wing tip, on hind wing, 1 large eyespot near upper margin and 1 small eyespot below it. Eyespots are black, yellow-rimmed, with iridescent blue and lilac irises, on fore wing, 1 very small near tip and 1 large eyespot in white fore wing bar. [Uncaptioned image]
A.2.2 CelebA-1k
We subsample the first $1000$ images from CelebA dataset [25] to construct CelebA-1k dataset. Prompts are generated using BLIP captioner with length between $25$ and $50$.
A.2.3 Kumapi
The Kumapi dataset [23] consists of $391$ images with handwriting styles. Prompts are generated using Waifu tagger, and we manually modify some prompts that do not reflect the images well.
A.2.4 Butterfly
Butterfly dataset [24] consists of $832$ images with $8$ species categories. We combine captions generated using the BLIP captioner and descriptions for species from the dataset. The length of captions generated by the BLIP captioner is between $5$ and $75$.
Appendix B Different CFG Scales
This section presents images from Chains of Diffusion for different CFG scales on four datasets outlined above. Figure 10, 11, 12, and 13 are for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets, respectively. For each dataset, results for $5$ different CFG scales are provided to cover high, low, and medium CFG scales. We found $2.5$ to be the optimal medium CFG scale for Pokemon and Kumapi datasets, and $1.5$ for CelebA-1k and Butterfly datasets. With CFG scales lower than the optimal value, low-frequency degradations are observed throughout the four datasets we demonstrate. On the other hand, high-frequency degradations with saturated color and repetitive patterns are universally observed with CFG scales higher than the optimal value. Notably, the optimal value for Pokemon and Kumapi causes the Chain of Diffusion for CelebA-1k and Butterfly datasets to degrade severely.
Figure 10: Chain of Diffusion with different CFG scales for Pokemon dataset. CFG scale $2.5$ is the optimal value we found.
Figure 11: Chain of Diffusion with different CFG scales for CelebA-1k dataset. CFG scale $1.5$ is the optimal value we found.
Figure 12: Chain of Diffusion with different CFG scales for Kumapi dataset. CFG scale $2.5$ is the optimal value we found.
Figure 13: Chain of Diffusion with different CFG scales for Butterfly dataset. CFG scale $1.5$ is the optimal value we found.
Appendix C Hyperparameter investigations to unveil the most impactful factor of degradation
In this section, we present experiment results to reveal the most impactful factor of degradation observed in the Chain of Diffusion. We conduct extensive investigations to check the effects of each hyperparameter outlined in Table 1. Based on the default hyperparameter setting in Table 2, we change each candidate hyperparameter and investigate whether any of them can meaningfully mitigate or affect the degradation. CFG scale is $7.5$ for all cases in this section.
C.1 Training set size
We used CelebA-1k dataset to investigate how the size of the training set (both $D_{0}$ and $D_{k}$) affects the degradation observed in the Chain of Diffusion. We chose this dataset as we can increase or decrease the training set size easily by subsampling from a larger original dataset. We changed the size of the training set to $100$, $250$, $500$, and $2000$. Degradations are observed regardless of the size of the training set, as shown in Figure 14. Degradations occur at earlier iterations when a smaller number of images are used, but it is hardly distinguished in $6$th iteration. We fix the number of parameter updates.
Figure 14: Chain of Diffusion with varying training set size on CelebA-1k dataset. Degradations begin to occur faster with smaller sizes, but all of them results in unrecognizable degree of degradations in $6$th iteration.
C.2 Number of images per prompt
Generating more than one image per prompt to increase the diversity of the training set is a straightforward method to mitigate the degradation in the Chain of Diffusion. We conducted an experiment to investigate how increasing the training set size by generating more images affects the degradation. Figure 15 demonstrates that $5$ times larger images cannot mitigate the degradation. It seems the degradation is slowed down by one iteration ($6$th iteration for the default case and $5$th iteration for the target case look similar), but at the cost of high computations, which makes it an infeasible solution.
Figure 15: Chain of Diffusion when more than one images are generated per prompt on Pokemon dataset. Increasing the training set size by generating more than one image per prompt cannot mitigate the degradation.
C.3 Mixing real images to synthetic sets
Many previous works suggested augmenting the training set with real images to mitigate the degradation observed in iterative training. We investigated if mixing images from the original training set to a synthetic set at each iteration can relieve the degradation. Images are randomly sampled at each iteration and replaced by corresponding real images in the original training set. Figure 16 and 17 demonstrate how the degradation in the Chain of Diffusion differs when $50\%$ and $90\%$ images at each iteration are replaced for Pokemon and CelebA-1k datasets, respectively. It is noteworthy that $10\%$ is far enough to cause the degradation, and $50\%$ rarely slows down the degradation. It is interesting that CelebA-1k dataset is much more vulnerable to degradation.
Figure 16: Chain of Diffusion when real images are randomly sampled and replace synthetic images in every iteration on Pokemon dataset. $50\%$ of real images rarely slow down the degradation, and $10\%$ is enough to commence the degradation.
Figure 17: Chain of Diffusion when real images are randomly sampled and replace synthetic images in every iteration on CelebA-10k dataset. The degradation is rarely slowed down even when only $10\%$ of synthetic images are used for finetuning. Natural images undergo more serious degradation compared to animation images in Figure 16.
C.4 Prompt Set
We hypothesized that the descriptiveness of prompts can affect the degradation of the Chain of Diffusion. Accordingly, we examined different sets of prompts for Pokemon and CelebA-1k datasets. The default prompt set used for Pokemon dataset is the concatenations of Waifu and BLIP captions. BLIP captions are generated to be between $50$ and $75$. We investigated how Waifu prompt set alone and different lengths of BLIP captions can affect the degradation in Figure 18. It is interesting that slightly different styles of high-frequency degradation are observed and short prompt set can relieve repetitive patterns at the cost of decreased diversity. Different lengths of BLIP captions in CelebA-1k dataset demonstrate similar degrees of degradation, but insufficient or excessive descriptiveness of prompt makes images to be different from the original ones as shown in Figure 19.
Figure 18: Chain of Diffusion when different prompts are used on Pokemon dataset. The default prompts are concatenations of BLIP caption of length between $50$ and $75$, and Waifu caption. We compare the Chain of Diffusion with Waifu captions, BLIP captions (used for default prompts), and short BLIP captions. Short captions are generated to have lengths smaller than $25$.
Figure 19: Chain of Diffusion when different prompts are used on CelebA-1k dataset. The default prompts have lengths between $25$ and $50$. We compare the Chain of Diffusion with the lengths of prompts larger than $50$ (Long) and smaller than $25$ (Short).
C.5 U-Net and Text-Encoder
Figure 20 shows the Chains of Diffusion when one of U-Net or text encoder is finetuned. Similar degradation happens when the text encoder is not updated (second row), and the degradation pattern is different when the U-Net of Stable Diffusion is not updated since image qualities that models can generate do not change in this case. However the contents of images are not preserved as text encoder is updated.
Figure 20: Chain of Diffusion when one of U-net and text encoder is finetuned on Pokemon dataset.
C.6 Number of diffusion steps
We examined if the number of diffusion steps during generation is insufficient and causes the degradation. Figure 21 demonstrates that the number of diffusion steps cannot improve the Chain.
C.7 Number of Epochs
We also tested if the training is insufficient or excessive for our default setting. However, images from the first iterations look similar in quality, as shown in Figure 21, and the Chains result in similar degradations. We finetune models for $100$ epochs as the default setting since loss values still decrease in $50$ epochs.
Figure 21: Chain of Diffusion when the number of diffusion steps and training epochs differ on Pokemon dataset. Both more diffusion steps and different training epochs cannot mitigate the degradation, and similar patterns are observed.
C.8 Learning Rate
Similarly, we examined in Figure 22 if finetuning is properly done with our default setting. We multiplied the learning rate for U-Net and text encoder by $10$ and $0.1$. Images of the first iterations indicate that the default values are appropriate for finetuning, and the degradation constantly occurs for different learning rate.
Figure 22: Chain of Diffusion when learning rate differs and Gaussian noise is added to the original training set on Pokemon dataset.
C.9 CLIP Skip
We examined if CLIP skip hyperparameter can affect the degradation. CLIP skip determines what intermediate feature from CLIP text encoder is used as text embedding for conditional generation. Smaller value means the feature closer to the output of CLIP text encoder is used, and larger values for features closer to the input text. Figure 23 shows that this hyperparameter rarely changes the degradation patterns.
Figure 23: Chain of Diffusion when different CLIP skip hyperparameter is used on Pokemon dataset. CLIP skip hyperparameter has negligible effect to the degradation.
C.10 Adding Gaussian noise to the original training set
We inspected if there are any differences between real and synthetic images, and if they can be cause the degradation. Random Gaussian noise is added to the original training set $D_{0}$. Figure 22 shows that the characteristics of the original training set have some effects, but the degradation still happens. This aligns with our observations that the degradations are universal to real, animation, and hand-written images.
C.11 Stable Diffusion XL
We investigated how image degradation occurs for different Stable Diffusion models. We note here that optimal hyperparameters for finetuning SDXL using LoRA have not been investigated enough yet, so we use the same hyperparameters to Stable Diffusion v1.5 and it can be suboptimal. Due to the space complexity, we reduce batch size to $2$. We also use the same resolution $512\times 512$ for this model as the first iteration images are impressively high-quality.
Figure 24: Chain of Diffusion of SDXL on Pokemon dataset. Optimal hyperparameters for finetuning SDXL have not been investigated enough yet, so our experiments are largely based on hyperparameters for Stable Diffusion v1.5. We note that images for Pikachu are provided to align with other cases, but SDXL has a huge bias in generating Pikachu well while it fails for other indices. Nonetheless, there are clear changes with different CFG scales.
Appendix D ReDiFine
Figure 25, 26, 27, and 28 show how robust ReDiFine is to different CFG scales. It successfully mitigates the high-frequency degradations for an extensive range of CFG scales.
Figure 25: Chain of Diffusion of ReDiFine with different CFG scales on Pokemon dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
Figure 26: Chain of Diffusion of ReDiFine with different CFG scales on CelebA-1k dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
Figure 27: Chain of Diffusion of ReDiFine with different CFG scales on Kumapi dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
Figure 28: Chain of Diffusion of ReDiFine with different CFG scales on Butterfly dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
In addition to visual inspections for images generated using our ReDiFine, we also compare the quantitative results of ReDiFine to baselines with different CFG scales on FID, CLIP, and SFD. Results are shown in Figure 29. For different datasets and metrics, ReDiFine achieves comparable performance to the optimal CFG scale ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly).
(a) Pokemon, CLIP.
(b) CelebA-1k, CLIP.
(c) Kumapi, CLIP.
(d) Butterfly, CLIP.
(e) Pokemon, FID.
(f) CelebA-1k, FID.
(g) Kumapi, FID.
(h) Butterfly, FID.
(i) Pokemon, SFD.
(j) CelebA-1k, SFD.
(k) Kumapi, SFD.
(l) Butterfly, SFD.
Figure 29: Quantitative results of ReDiFine and baselines (different CFG scales).
Appendix E Ablation study
This section provides an ablation study to understand how condition drop finetuning and CFG scheduling contribute to the success of ReDiFine.
E.1 Condition drop finetuning
We conducted an ablation study to understand how the probability of dropping text embedding during finetuning affects the image quality in the Chain of Diffusion. Based on the probability used for training Stable Diffusion ($0.1$ or $0.2$), we tested $0.1$, $0.2$, and $0.4$. For both Pokemon and CelebA-1k datasets, a probability of 0.2 seems to work the best, as shown in Figure 30 and Figure 31, respectively. Interestingly, condition drop finetuning helps to mitigate the color saturation problem, but its effect decreases with a higher probability. For both of these datasets, condition drop finetuning can mitigate image degradation to some degree, but still, there is a large quality degradation that needs to be improved.
Figure 30: Chain of Diffusion with condition drop finetuning on Pokemon dataset.
Figure 31: Chain of Diffusion with condition drop finetuning on CelebA-1k dataset.
E.2 CFG scheduling
We also tested how different decreasing strategies for CFG scale affect image degradation in the Chain of Diffusion. We used two different exponential orders for exponential decreasing, and also compared a linear decreasing strategy. Figure 32 shows that CFG scheduling is effective for Pokemon dataset, generating high-quality images comparable to those of ReDiFine. However, it fails to improve image quality on CelebA-1k dataset, as shown in Figure 33. This demonstrates the necessity of condition drop finetuning to improve the Chain of Diffusion universally for various datasets.
Figure 32: Chain of Diffusion with CFG scheduling on Pokemon dataset.
Figure 33: Chain of Diffusion with CFG scheduling on CelebA-1k dataset.
Appendix F Analysis
In this section, we provide various analyses of images generated through the Chain of Diffusion. Specifically, we provide frequency analysis of latents during diffusion steps and fingerprints of synthetic images for forensic analysis.
F.1 Latent analysis
(a) Histogram of latent vectors.
(b) Diffs.
(c) Power spectrum density.
Figure 34: Fingerprints of the original training set and synthetic sets for Pokemon dataset. There are clear differences between fingerprints of natural images (Original) and synthetic images. ReDiFine makes similar fingerprints to CFG scale $2.5$. (a) Autocorrelation. Horizontal and vertical lines gradually disappear for CFG scales $1.0$ and $7.5$ while they are maintained for CFG scale $2.5$ and ReDiFine. (b) Average power spectra. Central regions are amplified or diminished for CFG scales $1.0$ and $7.5$, demonstrating low and high-frequency degradation.
Figure 34 demonstrates how latent distribution evolves for different iterations and diffusion steps on Pokemon dataset. First, the histogram of the final latent vectors before decoding is shown in Figure 3434(a). For a CFG of $1.0$, the latent distribution quickly evolves into a Gaussian-like shape, with its variance decreasing over iterations. This behavior aligns with previous work [5, 12, 19], which theoretically predicted that the self-consuming loop would crop out the tails of the distribution, reducing output diversity until it collapses to a single mode. We believe that these narrowing distributions in the latent space lead to blurrier and more homogeneous outputs in the image space. The evolution of the latent distribution is completely opposite for $7.5$, creating longer tails and approaching a more uniform distribution over space. For a CFG scale of $2.5$, which has the best reusability among the three, the latent distribution is better preserved over iterations. ReDiFine better preserves the histogram from the first to the last iteration, achieving both first iteration fidelity and reusability.
[…]
Figure 3434(c) demonstrates how the power spectrum density is maintained by ReDiFine. While high and low CFG scales cannot preserve the power spectrum of the first iteration as the Chain of Diffusion proceeds, ReDiFine can maintain the density distribution during iterations. This prevents images from being saturated to high or low-frequency degradation. Moreover, the power spectrum is more similar to the original training set compared to CFG scale $2.5$.
In summary, it is shown that ReDiFine can successfully alleviate the symptoms we observed in the latent space for baselines in Section 3. We can notice that ReDiFine preserves the original latent distribution at the $6$th iteration as compared to other baselines with different CFG values. More importantly, the conditional-unconditional score differences (Diff) are kept at bay all the way until the 6th iteration, as opposed to the baseline CFG $7.5$, showing a severe discrepancy between the unconditional and conditional scores.
F.2 Fingerprints for forensic analysis
(a) Radial spectrum power density.
(b) Angular spectrum power density.
Figure 35: Power spectrum density of the original training set and synthetic sets for Pokemon dataset. Images generated by ReDiFine maintain power density distribution during Chain of Diffusion while baselines fail. Even CFG scale $2.5$ cannot maintain the distribution for the last iteration. (a) Radial spectrum power density. ReDiFine shows a density distribution similar to that of the original training set. (b) Angular spectrum power density. Power density of generated images by ReDiFine remains during the iterations while baselines cannot maintain angular distribution.
Fingerprints of synthetic images have distinct characteristics that differ according to the types of generative models, architectures, and training images. This provides unique patterns that can be used for synthetic image detection. Figure 9 shows how fingerprints of the original training set and synthetic images differ and how fingerprints of synthetic images change in the Chain of Diffusion. We use $1008$ images from Pokemon dataset. Following visual inspection, ReDiFine makes similar fingerprints to CFG scale $2.5$. Moreover, fingerprints for ReDiFine and baseline with CFG scale $2.5$ are maintained through the Chain of Diffusion. On the other hand, low or high CFG scales remain different patterns to the synthetic images as iterations proceed. Fingerprints gradually change for low or high CFG scales, showing the effects of high or low-frequency degradation.
Figure 35 demonstrates the radial and angular spectrum power density of the original and synthetic images. It is clearly observed that ReDiFine maintains a similar radial spectrum power density compared to the original training set, while even CFG scale $2.5$ fails. Moreover, it shows stable angular spectrums during the Chain of Diffusion despite it being distinct from that of the original training set.

### v2

@@
-ReDiFine: Reusable Diffusion Finetuning for Mitigating Degradation in the Chain of Diffusion
+Model Collapse in the Self-Consuming Chain of Diffusion Finetuning: A Novel Perspective from Quantitative Trait Modeling
 Abstract
-Diffusion models have achieved tremendous improvements in generative modeling for images, enabling high-quality generation that is indistinguishable by humans from real images.
-The qualities of images have reached a threshold at which we can reuse synthetic images for training machine learning models again.
-This attracts the area as it can relieve the high cost of data collection and fundamentally solve many problems in data-limited areas.
-In this paper, we focus on a practical scenario in which pretrained text-to-image diffusion models are iteratively finetuned using a set of synthetic images, which we call the Chain of Diffusion.
-Finetuned models generate images that are used for the next iteration of finetuning.
-We first demonstrate how these iterative processes result in severe degradation in image qualities.
-Thorough investigations reveal the most impactful factor for the degradation, and we propose finetuning and generation strategies that can effectively resolve the degradation.
-Our method, Reusable Diffusion Finetuning (ReDiFine), combines condition drop finetuning and CFG scheduling to maintain the qualities of generated images throughout iterations.
-ReDiFine works effectively for multiple datasets and models without further hyperparameter search, making synthetic images reusable to finetune future generative models.
+The success of generative models has reached a unique threshold where their outputs are indistinguishable from real data, leading to the inevitable contamination of future data collection pipelines with synthetic data.
+While their potential to generate infinite samples initially offers promise for reducing data collection costs and addressing challenges in data-scarce fields, the severe degradation in performance has been observed when iterative loops of training and generation occur—known as “model collapse.” This paper explores a practical scenario in which a pretrained text-to-image diffusion model is finetuned using synthetic images generated from a previous iteration, a process we refer to as the “Chain of Diffusion.” We first demonstrate the significant degradation in image quality caused by this iterative process and identify the key factor driving this decline through rigorous empirical investigations.
+Drawing an analogy between the Chain of Diffusion and biological evolution, we then introduce a novel theoretical analysis based on quantitative trait modeling.
+Our theoretical analysis aligns with empirical observations of the generated images in the Chain of Diffusion.
+Finally, we propose Reusable Diffusion Finetuning (ReDiFine), a simple yet effective strategy inspired by genetic mutations.
+ReDiFine mitigates model collapse without requiring any hyperparameter tuning, making it a plug-and-play solution for reusable image generation.
 1 Introduction
 Can state-of-the-art AI models learn from their own outputs and improve themselves?
-As generative AI (e.g., GPT [1], Diffusion [2, 3]) can now churn out uncountable synthetic texts and images, this question piqued curiosity from many researchers in the past couple of years [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
-While some show positive results of self-improving [4, 15], most report an undesirable “model collapse” [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 16, 17, 18]—a phenomenon where a model’s performance degrades when it goes through multiple cycles of training with the self-generated data.
-When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text that has a lot of repetitions [19], and its linguistic diversity declines rapidly [8, 16]; image models also show quality degradation [7, 10] and loss of diversity [12, 20].
-This phenomenon has even been compared to mad cow disease, a fatal neurodegenerative disease caused by feeding cattle with the carcasses of other cattle [12].
-Our work examines model collapse in a common use case of generative AI: customizing pretrained text-to-image generative models through finetuning.
-While previous works [5, 12, 14, 19] suggest including more real images in the training set to avoid model collapse, this is challenging for several reasons.
-First, the number of artworks generated by a single artist is often limited.
-Even highly productive artists create at most a few thousands of pieces during their lifetime (e.g., Monet produced 1,983 paintings over 61 years).
-Secondly, during image collection, it is easy to be deceived by realistic AI-generated images and accidentally include them in the training set.
-Therefore, we need to understand mode collapse behavior when the proportion of real data is limited and seek solutions beyond simply increasing the number of real images.
-In this paper, we propose a novel concept of generating “reusable” images, which refers to their potential for finetuning without causing catastrophic model collapse.
-(a) Self-Consuming Chain of Diffusion.
-(b) Image degradation with the Chain of Diffusion.
+As generative AI models (e.g., GPT, Diffusion) now churn out uncountable synthetic texts and images, this question piqued the curiosity of many researchers in the past couple of years.
+While some show positive results of self-improving [1, 2], most report an undesirable “model collapse”—a phenomenon where a model’s performance degrades when it undergoes multiple cycles of training with the self-generated data [3, 4, 5, 6, 7, 8, 9, 10].
+When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text with many repetitions [11], and its linguistic diversity declines rapidly [12, 13]; image models also show quality degradation [14, 15] and loss of diversity [16, 17].
+The goal of this paper is to investigate model collapse in the practical scenario of finetuning pretrained text-to-image diffusion models.
+An end user of text-to-image diffusion models often wants to finetune the latest model to generate images with a very specific style (e.g., creating new characters in the style of Pokémon).
+In fact, hundreds of new finetuned diffusion models are regularly uploaded on platforms like CivitAI [^1] https://civitai.com/, each designed to produce different styles of images.
+When users scrape the internet to collect images in the style they want, it becomes almost inevitable that synthetic images will be included in their datasets.
+This is because the number of real images is limited, while synthetic images can be generated in massive quantities and dominate online sources.
+As a result, users will feel compelled to include more synthetic images in their datasets to meet the data demands of these ever-growing, data-hungry models.
+To this end, we conduct a thorough investigation into how various hyperparameters commonly used during diffusion finetuning (e.g., learning rate, diffusion steps, prompts) impact model collapse.
+From our extensive empirical analysis, we make a crucial observation: the classifier-free guidance (CFG) scale is the most significant factor influencing the rate of model collapse.
+Moreover, we observe a fascinating phenomenon that the direction of degradation varies with different CFG values—low CFG results in low-frequency degradation in images, while high CFG leads to high-frequency degradation.
+The critical role of CFG in determining both the rate and direction of model collapse is a novel insight previously unrecognized in the literature.
+To gain a deeper theoretical understanding of this phenomenon, we draw upon the concept of quantitative trait modeling from statistical genetics.
+Unlike existing studies that attribute model collapse to limited sample sizes and conclude that it leads to zero variance [16, 3, 6], we propose that the underlying cause is a truncation-based selection process modulated by CFG.
+This theoretical model accurately describes our experimental observations.
+Moreover, we show that other theoretical work on model collapse can also be connected to statistical genetics models, such as random genetic drift and its variants [18, 19, 20, 21].
+This fresh angle of drawing parallels with statistical genetics offers a new framework to understand the mechanism of model collapse.
+Finally, inspired by our empirical results and theoretical insights, we introduce the Reusable Diffusion Finetuning (ReDiFine) method.
+Our experiments show that selecting an appropriate CFG can significantly slow down model collapse.
+However, finding the optimal CFG is computationally expensive, requiring numerous iterations of finetuning.
+To address this, we propose ReDiFine as a plug-and-play solution that achieves a comparable reduction in collapse rate without the need for any hyperparameter tuning.
+The default ReDiFine setting effectively mitigates model collapse across all four datasets we tested.
+Moreover, ReDiFine is effective at mitigating model collapse not only in a worst-case scenario of having a fully synthetic dataset, but also in a more practical case, where we have a mix of real and synthetic images in the training set.
+Our contributions can be summarized as follows:
+  • We performed thorough and extensive empirical investigations into model collapse when we finetune a diffusion model with diffusion-generated images.
+We tested an exhaustive list of experimental parameters on four datasets (two digital art & two natural images) to identify what affects model collapse.
+Our analysis reveals that CFG scale is the most influential factor that not only controls the rate of model collapse but also dictates the type of image degradation (Section 3).
+  • We provide a novel theoretical analysis of model collapse based on quantitative trait modeling that can accurately predict how power spectra of generated images evolve over iterations (Section 4).
+  • We propose a simple yet effective strategy to mitigate model collapse, named ReDiFine, which combines condition drop finetuning and CFG scheduling.
+Across all four datasets, ReDiFine successfully generates more “reusable” images, which can be reused to finetune a diffusion model without causing a severe model collapse (Section 5).
+2 Related Work
+The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [15, 10, 5, 16, 14, 12, 3, 7, 13, 4, 8, 9].
+Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [6], has been observed in both language and image generative models.
+Empirical studies on LLMs [13] reveal that linguistic diversity collapses, especially in high-entropy tasks [12], although this can be mitigated with data accumulation [2].
+In image generation, several works [15, 10, 16, 17, 14, 3] note image degradation when diffusion models are recursively trained with self-generated data.
+We conduct extensive empirical experiments to identify the most significant factor causing model collapse in text-to-image diffusion models.
+Our findings reveal that while model collapse is universally observed across various datasets and scenarios, it manifests in three distinct types of image degradation.
+(a) Self-Consuming Chain of Diffusion Finetuning.
+(b) Image degradation in the Chain of Diffusion.
 Figure 1: (a) Overall pipeline of the Chain of Diffusion.
 Given a pretrained text-to-image diffusion model $M_{0}$ and a prompt set $P$, a finetuned model $M_{k}$ is trained using $D_{k-1}$ generated at the previous iteration $k-1$.
-Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$.
-Chain of diffusion begins with the original dataset $D_{0}$.
+Then, $M_{k}$ generates $D_{k}$ using the same prompt set $P$, building a fully synthetic loop.
+Chain of diffusion begins with the original real dataset $D_{0}$.
 (b) Image degradation universally occurs during Chain of Diffusion across four datasets.
@@
 This consistently holds across four datasets and $10$ scenarios that we comprehensively investigate in Section 3.2.
-Our goal in this paper is two-fold: (1) to conduct an in-depth empirical analysis of model collapse behavior in the self-consuming chain of diffusion finetuning (referred to as Chain of Diffusion)—a setting already widely used by artists—and to examine which factors in finetuning and sampling affect the reusability of diffusion-generated images; (2) to develop a strategy that can improve the reusability of synthetic images.
-Our contributions toward these goals are summarized below:
-  • We consider a setting widely used for style transfer in diffusion models: finetuning Stable Diffusion [3] with low-rank adaptation (LoRA) finetuning [21] using a moderately-sized training set($\sim$1000100010001000 images).
-We examine the evolution of images in the Chain of Diffusion across four datasets: two digital art datasets (Pokemon [22] and Kumapi [23]) and two natural image datasets (Butterfly [24] and CelebA-1k [25]).
-Through extensive analysis, we show that model collapse universally occurs in all datasets regardless of training set size, real image mixing (up to 90%), and a wide range of scenarios with different configurations (see Figure 11(b) and Table 1).
-  • We discover that among all factors in the Chain of Diffusion, only one substantially impacts the speed of model collapse: classifier-free guidance (CFG) scale.
-With low CFG, we observe low-frequency degradation where images gradually become unrecognizably blurry.
-With high CFG scales, high-frequency degradation occurs, where some features of the images become amplified beyond the normal range with disturbing color saturation.
-There exists a medium CFG range that can significantly slow down model collapse (see Figure 22(b)).
-To the best of our knowledge, we are the first to demonstrate such a role of CFG in the Chain of Diffusion.
-  • Further, we reveal that the default CFG scale $7.5$ widely used for Stable Diffusion performs very poorly in terms of reusability metric (defined in Eqn (4)) despite its excellent image quality in the first iteration (measured by FID).
-This highlights the trade-off between first-iteration image quality and future reusability (see Figure 23).
-  • We develop a new finetuning strategy, Reusable Diffusion Finetuning (ReDiFine), to mitigate degradation inspired by our analysis of how CFG modulates the model collapse pattern in latent diffusion models.
-We combine condition drop (for finetuning) and CFG scheduling (for generation) in ReDiFine.
-While an optimal CFG scale can sufficiently slow down model collapse, it is impractical to sweep through CFG scales over multiple iterations for each dataset to find an optimal trade-off.
-With our proposed ReDiFine approach, we show that even the default CFG scale (CFG=7.5) can generate reusable images across all four datasets without the need for parameter tuning (see Figure 5 and 6).
-2 Related Work
-The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
-Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [11], has been observed in both language and image generative models.
-Empirical studies on LLMs [8, 15, 16] reveal that linguistic diversity collapses, especially in high-entropy tasks [8], although this can be mitigated with data accumulation [15].
-In image generation, several works [5, 7, 10, 12, 18, 20] note image degradation when diffusion models are recursively trained with self-generated data.
-The authors of [5] and [12] provide theoretical insights into diversity reduction, demonstrating that the covariance in a toy Gaussian model converges to zero as iterations increase in the self-consuming training loop.
-Among theoretical studies on this topic [13, 14, 17, 19], [17] is particularly relevant.
-They use dynamical system tools to show that generative distributions either collapse to a small set of outputs or become uniform over a large set, depending on the temperature setting.
-This finding is analogous to our observations with modulating CFG scales.
-Connecting our empirical analysis on finetuning Stable Diffusion with their theoretical framework could be an interesting future direction.
-Regarding mitigation strategies, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [5, 12, 14] or accumulating more data [15].
+Theoretical studies on model collapse have largely focused on the reduction of diversity, typically framed as either decreasing covariance in continuous domains [16, 6, 3] or shrinking support in discrete domains [11, 9].
+The finite number of generated samples per iteration has been identified as a key cause of model collapse by several authors [3, 6, 22, 11, 8], while others [16, 23, 9] emphasize that sampling bias reduces the generative model’s effective distribution.
+In contrast, we present a novel theoretical framework based on quantitative trait modeling, shifting the focus from variance reduction to mean drift induced by the selection process.
+We further demonstrate that many existing theories of model collapse can be understood as extensions of classical results from statistical genetics.
+Regarding mitigation strategies for model collapse, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [16, 3, 8, 23] or accumulating additional data [2].
 The only mitigation strategy beyond altering the training data composition is proposed by [4], who suggests a self-correcting self-consuming loop using an expert model to correct synthetic outputs.
-While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not always be feasible for finetuning customized image generative models.
-In our work, we explore an alternative solution through reusable image generation.
-By making small improvements in finetuning and sampling steps, we can substantially enhance the reusability of images, allowing them to be used for many more iterations before the model starts to deteriorate.
-In a tangentially related work [9], the authors discuss bias amplification in self-consuming chains, noting minimal amplification over iterations.
-3 Model Collapse in Self-Consuming Chain of Diffusion Finetuning
-Table 1: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse.
-Dash (--) in table entries indicates that the default value was used.
-All experiments were conducted on Pokemon dataset except for the dataset size.
-For dataset size, we use CelebA since its original dataset is bigger than Pokemon and we can subsample $500$, $1000$, $2000$ images.
-More details in Appendix A.
-Candidates Default 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
-Learning Rate 1e-4 1e-5 1e-3 -- -- -- -- -- -- -- -- -- -- -- -- -- -- -- --
-Finetune Both -- -- Unet Text -- -- -- -- -- -- -- -- -- -- -- -- -- --
-Noise No -- -- -- -- Yes -- -- -- -- -- -- -- -- -- -- -- -- --
-Prompt [^1] C, W, S, and L for Combine, Waifu, Short, and Long, respectively.
-We concatenate BLIP and Waifu captions as default setting, referred as Combine.
-Short and Long prompts are BLIP captions generated with limitations in the lengths of captions.
-More details can be found in Appendix C.4.
-C -- -- -- -- -- W S L -- -- -- -- -- -- -- -- -- --
-CLIP skip $2$- -- -- -- -- -- -- -- $1$ $3$- -- -- -- -- -- -- --
-Diffusion Steps $30$- -- -- -- -- -- -- -- -- -- $100$- -- -- -- -- -- --
-Epochs $100$- -- -- -- -- -- -- -- -- -- -- $50$ $200$- -- -- -- --
-Real mixing $0.0$- -- -- -- -- -- -- -- -- -- -- -- -- $0.5$ $0.9$- -- --
-Img / Prompt $1$- -- -- -- -- -- -- -- -- -- -- -- -- -- -- $5$- --
-Dataset Size $1000$- -- -- -- -- -- -- -- -- -- -- -- -- -- -- -- $500$ $2000$
+While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not be feasible and could be too costly for many real-world applications.
+In our work, we propose an alternative solution through reusable image generation, and show that it can mitigate model collapse more effectively than mixing more real data.
+3 Model Collapse in Self-Consuming Chain of Diffusion
 3.1 Problem Setting & Experimental Setup
-Chain of Diffusion
+Chain of Diffusion.
 We begin with formally defining the self-consuming Chain of Diffusion finetuning.
 Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$.
-$M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a common adaptation case of a pretrained generative model with a personally collected dataset.
+$M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a fully synthetic loop [16] for finetuning a pretrained generative model.
 Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
@@
 The overall pipeline of the Chain of Diffusion is shown in Figure 11(a).
+Figure 2: Left: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse.
+All experiments were conducted on Pokemon except for the dataset size.
+For dataset size, we use CelebA since its original dataset is bigger, and we can subsample $500$, $1000$, and $2000$ images.
+For prompt, we concatenate prompts with different lengths.44footnotemark: 4 Right: All hyperparameter settings other than changing CFG show a high collapse rate greater than $100$, i.e., FID score increases by $\sim$2x in 6 iterations, indicating severe image degradation.
+The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the collapse rate defined in Equation (3).
+For both $\text{FID}_{1}$ and collapse rate, lower is better.55footnotemark: 5
 Model and datasets.
@@
 More implementation details can be found in Appendix A.1 and A.2.
-(a) Reusability-FID trade-off [^3] We do not display scenarios that change the training dataset size, such as Img/Prompt and Dataset Size, as varying sizes result in FID scores on different scales.
-The related results are presented in the Appendix C, where we observe similar image degradation..
-(b) Image degradation across different CFGs.
-Figure 2: (a) Reusability-FID trade-off plot (Pokemon) shows that all scenarios other than changing CFG scales show poor reusability, amplifying the FID score $\sim$3x in 6 iterations of the chain.
-The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the reusability defined in equation (4).
-For both $\text{FID}_{1}$ and reusability, lower is better.
-reusability (500 $\sim$ 1100) is about two times larger than $\text{FID}_{1}$ (300 $\sim$ 600), indicating $\text{FID}_{6}$ three times larger than $\text{FID}_{1}$.
-(b) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chains of Diffusion.
-CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down.
-More images in Appendix B.
 Evaluation metrics.
-We use CLIP score [27] to quantify text-to-image alignment and Frechet Inception Distance (FID) score [28] to measure image fidelity.
-Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
-${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (3)
-to evaluate each iteration $k$.
-SFD overcomes the problem of FID being sensitive to the number of images to compare.
-Following [29], we use DiNOv2 [30] as a feature extractor for both FID and SFD since it is more consistent with our visual inspection than Inception-V3 network [31].
+We use Frechet Inception Distance (FID) [31] to measure image fidelity.
+Following [32], we use DiNOv2 [33] as a feature extractor for FID since it is more consistent with our visual inspection than Inception-V3 network [34].
 With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 features as the FID score.
-New evaluation metric: Reusability.
+Quantitative results for additional metrics including CLIP score [35] can be found in Appendix E.
 In addition to evaluation metrics for generative models presented above, we propose a new metric to quantify the reusability of images generated in the Chain of Diffusion.
-We define a reusability metric as the performance discrepancy between the first iteration and the $K$-th iteration in the Chain of Diffusion:
-$K\text{-reusability}={\rm FID}_{K}-{\rm FID}_{1},$ (4)
+We define collapse rate as the performance degeneration per iteration in the Chain of Diffusion:
+collapse $\text{collapse rate}=\Delta{\rm FID}=\frac{{\rm FID}_{K}-{\rm FID}_{1}}{K-1},$ (3)
 where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set.
-We use FID as a performance metric here, but this can be CLIP, SFD, or any other performance metric of interest.
-Note that a low $K$-reusability indicates more reusable images since it means that the model does not degrade much in $K$ iterations.
-In the rest of the paper, $K=6$, and we simply call it reusability.
-(a) t-SNE of images.
-(b) CLIP.
-(c) FID.
-(d) SFD.
-Figure 3: (a) t-SNE plot visualizes how the distribution of generated images shifts from the original distribution (black) in the Chain of Diffusion and shows distinct paths for three different CFG scales for Pokemon dataset.
-Different CFG scales and iterations are differentiated with different colors and transparency.
-Arrows indicate how the distributions of generated images of chains move for different CFG scales.
-While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (1.0 and 7.5) deviate fast, indicating image degradation.
-(b-d) Qualitative comparison for different Chains using CLIP $\uparrow$, FID $\downarrow$, and SFD $\downarrow$ for Pokemon dataset.
-CFG $2.5$ achieves the most robust performance for all three scores.
-CFG $1.0$ begins to gradually degrade from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in Figure 22(b).
-Additional results for different datasets can be found in Appendix D
+We use FID as a performance metric here, but this can be CLIP, or any other performance metric of interest.
+Note that a low collapse rate indicates more reusable images since it means that the model does not degrade much during the Chain of Diffusion and we have $K=6$ in the rest of the paper.
 3.2 Model Collapse in the Chain of Diffusion
@@
 We observe significant image degradation in all four datasets (see Figure 11(b)) in the Chain of Diffusion.
-The quality of generated images begins to clearly deteriorate in the third iteration and it drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations.
-We investigated a variety of different scenarios (summarized in Table 1) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon.
-We tested various dataset sizes for $D_{0}$ and $D_{k}$ (Appendix C.1), increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt (Appendix C.2), mixing real images from $D_{0}$ to $D_{k}$ (Appendix C.3), changing the descriptiveness of prompts (Appendix C.4), freezing U-Net or text encoder during finetuning (Appendix C.5), and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip) (Appendix C.6, C.7, C.8, and C.9).
-We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability (Appendix C.10), and investigated if the degradation happens for larger Stable Diffusion (Appendix C.11).
-In all settings we tested, image degradation was universally present.
-We plot the results for Pokemon dataset on Figure 23 where y-axis is reusability and x-axis is the ${\rm FID}_{1}$.
-While adding noise and mixing $90\%$ real images show the best reusability (the lower, the better), they still exhibit significant degradation (FID score has been doubled in $6$ iterations).
-More visual inspections and metric plots can be found in Appendix C.
-Observation 2: CFG is the only factor that impacts the model collapse.
+The quality of generated images begins to clearly deteriorate in the third iteration and drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations.
+We investigated a variety of different scenarios (summarized in Figure 2) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon.
+We tested various dataset sizes for $D_{0}$ and $D_{k}$, increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt, mixing real images from $D_{0}$ to $D_{k}$, changing the descriptiveness of prompts, freezing U-Net or text encoder during finetuning, and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip).
+We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability, and investigated if the degradation happens for larger Stable Diffusion.
+In all settings we tested, image degradation was universally present and very fast.
+We plot the results for Pokemon on Figure 2 where y-axis is collapse rate and x-axis is the ${\rm FID}_{1}$ (better when closer to the origin).
+While adding noise to images and mixing $90\%$ real images to every iteration as proposed by [16, 3, 8, 23, 2] show the lowest collapse rate, they still exhibit significant degradation as shown in quantitative values (FID score has been doubled in $6$ iterations).
+Visual inspection for images is provided in Appendix C.
+Observation 2: CFG is the most significant factor that impacts the model collapse.
 Throughout all our experiments, classifier-free guidance (CFG) had the biggest impact in the speed of model collapse.
@@
 […]
-High CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images.
+A higher CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images.
 On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images.
 For those familiar with temperature sampling [37], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
-In Figure 23, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$), which is expected from our understanding of CFG.
-A more surprising part is that this comes at the cost of a worse reusability metric (increase in ${\rm FID}_{6}-{\rm FID}_{1}$).
-Also, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and reusability.
-Similarly, when the CFG scale is too low—below $2.0$—the improvement in reusability plateaus and both ${\rm FID}_{1}$ and reusability begin to decline.
-Interestingly, there is an optimal region of CFG values (near $2.5$, specific to Pokemon dataset), where we achieve good reusability while maintaining a good quality in the first iteration as well.
-Figure 3 presents quantitative results for Pokemon dataset to demonstrate how different CFG scales affect the performance in CLIP, FID, and SFD.
-CFG scale $2.5$ achieves the most robust performance for all metrics.
-(a) Probability density function (PDF) of values of latent vectors at the last diffusion step.
-(b) Norm of differences between conditional and unconditional scores during diffusion steps.
-Figure 4: Latent analysis to investigate how different CFG scales affect the latent of Stable Diffusion (Pokemon dataset).
-(a) Latent distribution shrinks over iteration for low CFG and expands with high CFG.
-Larger values in latent vectors are more likely to occur with high CFG, gradually increasing the tail of the distribution.
-(b) Differences between conditional and unconditional scores increase as the training set is more degraded.
-Especially, high differences in the later diffusion steps can be a cause of high-frequency degradation.
+In Figure 2, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$ on x-axis), which is expected from our understanding of CFG.
+A more surprising part is that this comes at the cost of a worse collapse rate (an increase on the y-axis).
+Additionally, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and collapse rate.
+Similarly, when the CFG scale is too low—below $2.0$—the improvement in the collapse rate plateaus and both ${\rm FID}_{1}$ and collapse rate begin to increase.
+There is an optimal region of CFG values (near $2.5$, specific to Pokemon), where we achieve low collapse rate while maintaining a good quality in the first iteration as well.
+Moreover, Figure 33(c) presents FID for Pokemon to demonstrate how different CFG scales affect the performance during iterations.
+CFG scale $2.5$ achieves the most robust performance for Pokemon across iterations.
+Interestingly, optimal CFG scales differ for different styles: $2.5$ for animated or hand-writing datasets (Pokemon and Kumapi) and $1.5$ for photo datasets (CelebA and Butterfly) as shown in Appendix B.
+(a) Degradations across different CFGs.
+(b) t-SNE of images.
+(c) FID $\downarrow$.
+Figure 3: (a) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chain of Diffusion.
+CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down.
+More images in Appendix B.
+(b) t-SNE plot visualizes how generated images evolve from the original distribution (black) and shows distinct paths for three CFG scales (Pokemon).
+Different CFG scales and iterations are differentiated with different colors and transparency.
+Arrows indicate how the distributions of generated images move for different CFG scales.
+While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (7.5 and 1.0) deviate fast, indicating image degradation.
+(c) Quantitative comparison for FID $\downarrow$ (Pokemon).
+CFG $2.5$ achieves the most robust performance.
+CFG $1.0$ degrades from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in (a).
 Observation 3: High CFG scales cause high-frequency degradation and low CFG scales cause low-frequency degradation.
-CFG scale does not only affect the speed of model collapse, but also the pattern of model collapse.
-As shown in Figure 22(b), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to low-frequency degradation.
-On the other hand, for CFG $7.5$ how images degrade looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated.
-The t-SNE plot in Figure 33(a) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales.
+The CFG scale affects affect not only the speed of model collapse, but also the pattern of model collapse.
+As shown in Figure 33(a), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to as low-frequency degradation.
+On the other hand, for CFG $7.5$ the degradation pattern looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated.
+The t-SNE plot in Figure 33(b) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales.
 These patterns were consistent in all four datasets and detailed results can be found in Appendix B.
-We further analyze this phenomenon by examining the latent space.
-First, we show a histogram of the final latent vectors before decoding in Figure 44(a).
-For a CFG of $1.0$, the latent distribution quickly evolves into a Gaussian-like shape, with its variance decreasing over iterations.
-This behavior aligns with previous work [5, 12, 19], which theoretically predicted that the self-consuming loop would crop out the tails of the distribution, reducing output diversity until it collapses to a single mode.
-We believe that these narrowing distributions in the latent space lead to blurrier and more homogeneous outputs in the image space.
-For a CFG of $7.5$, the evolution of the latent distribution is completely opposite, creating longer tails and approaching a more uniform distribution over space.
-For a CFG scale of $2.5$, which has the best reusability among the three, the latent distribution is better preserved over iterations.
+In Section 4, we detail how different patterns of power spectra in high-frequency regions from different CFG scales can be understood using the framework from genetic biology.
+Implications of our observations.
+Our extensive investigations reveal that a high CFG of $7.5$, a common choice to generate visually appealing images, significantly increases collapse rate to achieve slightly better $\rm{FID}_{1}$.
+Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [16].
+While they reported a monotonic increase in collapse rate as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$.
+It shows an intriguing trade-off between perceptual quality and reusability.
+This suggests that developers concerning reusability of images can substantially improve future generations by carefully choosing CFG.
+4 Quantitative Trait Models for Fully-Synthetic Training Loops
+This section introduces a novel perspective for understanding the model collapse in generative models by drawing parallels between genetic biology and the Chain of Diffusion.
+Distinguishable iterations in the Chain of Diffusion—where each iteration is clearly separable with no duplicated individuals, and the current iteration originates from the previous one—mirror genetic processes involving successive iterations of parents and offspring.
+By applying quantitative trait modeling from statistical genetics, we provide a framework to describe how images evolve across iterations in the Chain of Diffusion.
+We begin by introducing quantitative trait modeling and its underlying mathematical assumptions.
+Leveraging these assumptions, we derive a theorem demonstrating that the mean trait value exhibits linear drift while the variance stabilizes over time.
+We then show that this model can successfully capture the three key behaviors observed in Section 3: high-frequency degradation, low-frequency degradation, and optimal CFG scale performance.
+This suggests that different CFG scales in the Chain of Diffusion can be viewed as varying selection strategies, with power spectra corresponding to quantitative traits.
+We use the term “iteration” instead of “generation” for genetic generation to avoid confusion with image generation.
+(a) Selection mechanism.
+(b) Power spectra of images vs. Mean of phenotypes.
+Figure 4: (a) Selection mechanism with two-sided truncation.
+$r_{1}$ and $r_{2}$ ratios of samples are truncated from the left and right tails, and the remaining $r=1-r_{1}-r_{2}$ ratio of samples is selected.
+(b) Power spectra for different CFGs (Pokemon) and quantitative trait modeling results of different selection strategies.
+Directional selections with truncation can effectively explain our observations in Section 3: the behaviors of high- and low-frequency degradation and optimal CFG scale.
+Detailed settings are provided in Appendix D.2.
+4.1 Quantitative Trait Modeling
+Quantitative trait modeling in statistical genetics explores the evolution of quantitative phenotypes (e.g., height, weight, or color), which are determined by multiple genetic and environmental factors in complex ways.
+These phenotypes are typically assumed to follow a Gaussian distribution, with discrete iterations where parents and offspring are distinguishable ($t$- and $t+1$-th iterations are clearly separable).
+Let the distribution of phenotypes at the $t$-th iteration be denoted as $\mathcal{N}(\mu_{t},\sigma_{P,t}^{2})$ where $\mu_{t}$ and $\sigma_{P,t}^{2}$ are the mean and variance of quantitative phenotypes.
+The phenotypic variance $\sigma_{P,t}^{2}$ is the sum of the (additive) genetic variance $\sigma_{G,t}^{2}$ and the environmental variance $\sigma_{E}^{2}$, i.e., $\sigma_{P,t}^{2}=\sigma_{G,t}^{2}+\sigma_{E}^{2}$ [38] [^6] Genetic variance is composed of additive, dominance, and interaction variance.
+Here, we only consider additive variance, which is a common assumption in the field..
+When selection occurs in each iteration, whether natural (e.g., faster animals surviving predators) or artificial (e.g., breeding livestock for higher milk production), it affects the distribution of the effective population that influences the next iteration.
+We consider directional selection with truncation, as shown in Figure 44(a), where a ratio of $r$ samples is selected by truncating $r_{1}$ from the left and $r_{2}$ from the right side of the distribution.
+Here, $r+r_{1}+r_{2}=1$ and larger phenotype values are preferred when $r_{1}>r_{2}$.
+The (narrow-sense) heritability [^7] The heritability is narrow-sense when the genetic variance is restricted to additive variance., which is defined as the proportion of phenotypic variance attributable to additive genetic factors, can also be represented using the Breeder’s Eqn.
+[39] as:
+$h_{t}^{2}=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}}=\frac{\sigma_{G,t}^{2}}{\sigma_{G,t}^{2}+\sigma_{E}^{2}}=\frac{\mu_{t+1}-\mu_{t}}{\mu_{t}^{\prime}-\mu_{t}},$ (5)
+where the mean phenotype of the next iteration $\mu_{t+1}$ is represented using the mean phenotype of selected individuals $\mu_{t}^{\prime}$, the mean phenotype of the entire population at current iteration $\mu_{t}$, and heritability $h_{t}^{2}$.
+Additionally, we assume the genetic variance for the next iteration is determined by the variance of selected individuals $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}$.
+We prove the behaviors of mean and variance of phenotypes under this setting:
+Theorem 1
+Suppose the distributions of phenotypes follow Gaussian distribution and directional selection truncates individuals on both sides with ratios $r_{1}$ and $r_{2}$.
+Then mean of phenotypes asymptotically increases (decreases) by $\frac{c_{1}c_{2}}{\sqrt{1-c_{2}}}\sigma_{E}$ per iteration and the variance converges to $\frac{1}{1-c_{2}}\sigma_{E}^{2}$ when $r_{1}>r_{2}$ ($r_{2}<r_{1}$), where $c_{1}$ and $c_{2}$ are constants that depend on $r_{1}$ and $r_{2}$.
+The proof of Theorem 1 is provided in Appendix D.1.
+4.2 Explaining the Chain of Diffusion with Quantitative Trait Modeling
+Just as phenotypes evolve through complex functions of hidden genotypes across multiple iterations in quantitative trait modeling, various continuous features of images evolve similarly during the Chain of Diffusion.
+A key discovery in our work is that the high-frequency power spectra are crucial phenotypes that evolve over iterations, with the CFG scales acting as different selection mechanisms.
+These power spectra are computed using 2D Fourier transforms, focusing on frequencies above a certain threshold to capture high-frequency components.
+Different CFG scales during generation correspond to different selection strategies: a high CFG selects individuals in the right tail of the distribution, favoring more detailed features with reduced diversity, while a low CFG selects from the left tail.
+Figure 44(b) shows the evolution of power spectra in the Chain of Diffusion (in solid line) and the phenotype mean modeled with Equation  (9) (in dotted lines).
+The initial mean $\mu_{0}$ and genetic deviation $\sigma_{G,0}$ in the simulation are set to match those of the original image set at iteration 00.
+Three different configurations of ratios, $r_{1}$ and $r_{2}$, successfully model the power spectra distribution: CFG $7.5$ is modeled as selecting from upper $5$% of samples, CFG $2.5$ selects $50$% of the samples slightly favoring higher frequency, and CFG $1.0$ results from selecting the lower $30$%.
+We can see that this modeling can capture the evolution of power spectra very accurately over 6 iterations of Chain of Diffuson.
+Details about power spectra computation and simulation parameters are provided in Appendix D.2.2.
+4.3 Statistical Genetics as a Unifying Lens to Model Collapse
+Quantitative trait modeling, a tool borrowed from statistical genetics, serves as a solid theoretical framework for interpreting our experimental results.
+As genetic processes and the self-consuming training of generative models share a lot of similarities, we find that other concepts in mathematical genetics also have close ties to existing theoretical work on model collapse.
+For instance, the traditional Wright-Fisher model [19] describes how traits evolve from one generation to the next in a finite population.
+Due to randomness in the sampling process, the composition of traits within a finite population drifts over time and eventually collapses to a single phenotype.
+The Wright-Fisher model and its continuous variants [40] are exactly equivalent to the collapse behavior modeled with simple categorical or Gaussian distributions in [16, 3, 6, 22].
+Given its widespread use in genetics, many extensions of the Wright-Fisher model remain an active area of research, such as including different mechanisms of selection [41, 42] or mutations [43].
+Beyond the current work and existing theoretical studies on model collapse, we believe that statistical genetics offers a unified perspective to understand model collapse by reducing the complex dynamics of the self-consuming loop of generative model training to a small number of parameters, from which we can gain valuable insights.
+5 Reusable Image Generation with ReDiFine
+In the previous sections, we have discovered a significant role of CFG in model collapse and that an appropriate choice of CFG can effectively mitigate the collapse.
+However, the optimal CFG value that minimizes the collapse rate is different for each dataset (e.g., CFG 1.5 for CelebA, CFG 2.5 for Pokemon).
+There is no efficient way of finding an optimal CFG other than performing a grid search over a wide range of values and iteratively fine-tuning the diffusion model to evaluate the collapse rate for each configuration.
+In practical scenarios, it is unlikely that end users who are not machine learning experts will go through such a process during finetuning, just to prevent a potential model collapse that can happen in the next generation.
+This raises the question: How can we design a finetuning and generation strategy that is user-friendly and can effectively slow down model collapse?
+To address this question, we again draw inspiration from the evolution process in nature.
+In biology, mutations naturally counteract genetic drift and preserve diversity.
+Furthermore, selection in nature is often a soft process rather than a hard truncation, as illustrated in Figure 44(a).
+This soft selection allows for the inclusion of outliers, thereby maintaining the overall genetic diversity.
+We connect these biological inspirations with two strategies: condition drop finetuning to introduce more randomness and CFG scheduling during generation to transform hard selection to a softer one [^8] We show a modified quantitative trait modeling result with these two modifications in Appendix D.2 and show that it captures the ReDiFine experimental results effectively.
+In this section, we focus on showing the image generation results with ReDiFine.
+We refer the curious readers to the appendix..
+We propose Reusable Diffusion Finetuning (ReDiFine) method that incorporates these two ideas and show that it is highly effective at mitigating model collapse without the need for any finetuning.
 […]
-Implications of our observations.
-Through extensive investigations, we gained a comprehensive understanding of CFG scales’ effects on the Chain of Diffusion.
-Even though a high CFG scale of $7.5$ is a common choice to generate visually appealing images, it significantly sacrifices reusability to achieve slightly better FID.
-Moreover, it is noteworthy that CLIP score increases for the early three iterations and drops rapidly in the latter when the CFG scale is $7.5$ (Figure 33(b)).
-This shows that Stable Diffusion tends to generate images favored by CLIP encoder, which can contain artifacts easily amplified by further finetuning.
-Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [12] and they provide FID plots with varying levels of sampling bias by changing the CFG scale.
-While they reported a monotonic decrease in reusability as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$.
-It shows an intriguing trade-off between perceptual quality and reusability.
-This suggests that if developers of diffusion models are mindful of the reusability metric, can improve future generations of images substantially by carefully choosing CFG.
-4 ReDiFine
-In the previous section, we showed that choosing a good CFG scale can considerably improve reusability in the Chain of Diffusion.
-However, such optimal CFG scales vary from dataset to dataset.
-To mitigate model collapse solely with CFG tuning, we have to do multiple rounds of finetuning and sampling to evaluate the reusability-FID trade-off.
-This is too computationally burdensome to be a practical solution.
-In this section, we propose ReDiFine, Reusable Diffusion Finetuning, which can mitigate model collapse without any parameter tuning.
-ReDiFine combines two ideas—condition drop finetuning and CFG scheduling—and effectively improves the reusability-FID trade-off in all four datasets (comparable to optimal CFG scales).
-4.1 Condition drop finetuning.
-We propose condition drop finetuning, which randomly drops the text condition during finetuning to update both conditional and unconditional scores.
-Even though condition drop was suggested in the original CFG paper [32], it is not commonly done during finetuning since the first-iteration image qualities are outstanding without updating unconditional scores.
-However, the small difference between conditional and unconditional scores can accumulate over iterations and create a large gap as the iteration progresses in the Chain of Diffusion if we do not explicitly finetune the unconditional scores as shown in Figure 44(b).
-Inspired by this observation, we drop text conditions with probability $0.2$ during finetuning.
-We illustrate the effect of condition drop with CFG $7.5$ in Figure 44(b).
-Note that the norm of Diff does not grow as fast even with a high CFG of $7.5$.
-While this prevents a potential cause of the high-frequency degradation, the Diff value still spikes up in iteration 6.
-We conjecture that this is because the $D_{5}$ already contains severe artifacts that conditional scores have to grow substantially to fit the data.
-4.2 CFG Scheduling.
-Based on the analysis showing that a high CFG scale amplifies the discrepancy between the condition and uncondition scores during iterative fine-tuning (see Figure 44(b)), we propose gradually reducing the CFG scale during diffusion steps.
-This CFG scheduling mechanism mitigates the negative impacts of overemphasizing the conditional score in later stages while preserving the initial conditional information in the early stages.
-Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as
-$s=s_{0}\times e^{-\alpha\times t/T}$ (6)
-where $T$ is $30$ and $s_{0}$ is $7.5$. $\alpha$ denotes the degree of decrease and we use $2$ throughout this paper. Comparison to different $\alpha$ and linear decreasing can be found in Appendix E.
-Our method gradually decreases the CFG scale during diffusion steps, leveraging the benefits of high CFG scale in the early stage and low CFG scale in the later stage.
-This adaptive approach aligns with observations from [34] that different steps contribute uniquely to the generation process.
-In addition, our CFG scheduling takes advantage of using a small CFG scale without the need for extensive hyperparameter searches for each dataset, providing a flexible and robust choice for the Chain of Diffusion.
-5 Reusable Image Generation
+CFG Scheduling.
+We propose gradually reducing the CFG scale during diffusion steps to mitigate the impacts of overemphasizing the conditional score in later stages, which can lead to high-frequency degradation.
+Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as $s=s_{0}\times e^{-\alpha\times t/T}$, where $s_{0}$ is the initial CFG scale and $\alpha$ is the rate of exponential decay.
+This scheduling approach is consistent with findings by [44], which suggest that different diffusion steps contribute uniquely to the generation process.
+High CFG lets us capture high-level semantic information accurately during the initial diffusion steps while lower CFG in later steps can prevent generating unnecessary high-frequency details.
 Figure 5: ReDiFine effectively mitigates image degradation in the Chain of Diffusion.
@@
 (d) Butterfly.
-Figure 6: ReDiFine effectively mitigates Reusability-FID trade-off across all four datasets.
-While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low reusability and FID at the same time (lower is better).
-While CFG scheduling alone works well for Pokemon dataset, it is still beneficial to have condition drop finetuning for CelebA-1k and Kumapi datasets.
-Note that the differences in FID are relatively smaller than those in reusability, supporting the necessity to evaluate reusability in the Chain of Diffusion.
-In this section, we present generated images and qualitative metrics for ReDiFine.
-We use the CFG scale $s_{0}=7.5$ for all datasets and experiments except for the ablation study.
-Note that ReDiFine can be further improved with hyperparameter tuning, e.g., CFG scale, condition drop probability, and CFG scheduling function.
-However, our focus in this work is to demonstrate that ReDiFine can robustly stabilize iterative finetuning regardless of datasets and hyperparameters.
-This robustness provides a great advantage when applying our ReDiFine to finetune Stable Diffusion and generate images using any personally collected image sets.
-Visual comparison.
-Comparing Figure 5 with Figure 11(b), ReDiFine effectively slows down and mitigates the image degradation in the Chain of Diffusion compared to using the default CFG scale $7.5$ across four datasets.
-In addition, ReDiFine successfully achieves the same level of high-quality image generation in the Chain of Diffusion compared to using the optimal CFG scales for each dataset, which requires extensive hyperparameter tuning.
-This is demonstrated by visually comparing the generated images in Figure 5 with the ones generated using the optimal CFG scales for Pokemon and CelebA dataset in Figure 22(b).
-Additional comparisons for Kumapi and Butterfly datasets can be found in Appendix B.
-Mitigate Reusability-FID trade-off.
-In addition to the visual comparison, we also provide quantitative results in Figure 6.
-We can see that ReDiFine consistently helps mitigate Reusability-FID trade-off in the Chain of Diffusion across all four datasets.
-In contrast, a fixed CFG scale that works best on one dataset fails on others.
-For example, the optimal CFG scale $2.5$ for Pokemon dataset achieves high (poor) ${\rm Reusability}$ for CelebA-1k and Butterfly datasets.
-Moreover, the optimal CFG scale $1.5$ for CelebA-1k and Butterfly datasets results in the high (poor) ${\rm FID}_{1}$ for Pokemon dataset.
-In addition to its ability to generalize across different datasets, ReDiFine demonstrates robustness when initialized with different CFG scales.
-Illustrated in Figure 66(a), ReDiFine achieves comparable performance with CFG scales of $5.0$ and $10.0$ as CFG $=7.5$.
-Ablation study.
-We provide an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine.
-As shown in Figure 6, CFG scheduling alone achieves comparable reusability in the Pokemon dataset, but condition drop finetuning is essential for CelebA-1k and Kumapi.
-Combining these two, ReDiFine consistently achieves the best trade-off between ${\rm Reusability}$ and ${\rm FID}_{1}$ across all four datasets.
-Additionally, the differences in ${\rm FID}_{1}$ are relatively small—using different CFG scales can generate high-quality images in the first iteration except for very low CFG scales, e.g., CFG = $1.0$.
-In contrast, the variance of ${\rm Reusability}$ is much larger.
-This supports the necessity of evaluating our proposed ${\rm Reusability}$ metric for the Chain of Diffusion.
-Additional results can be found in Appendix E.
-More iterations.
-In addition to the comparison over 6 iterations, Figure 7 illustrates the evolution of image degradation over 12 iterations.
-Notably, as fine-tuning iterations increase, high-frequency degradation emerges with the optimal CFG, highlighting the necessity for additional hyperparameter search to sustain image quality across longer diffusion lengths.
-Conversely, ReDiFine preserves image quality without high-frequency degradation.
-Figure 7: Comparison of the optimal CFG scale and ReDiFine on 12 iterations for Pokemon dataset.
-(a) Histogram of latent values.
-(b) Norm of differences.
-(c) Power spectrum of latents.
-Figure 8: Comparison of latent analysis at the 6th iteration for ReDiFine and baselines.
-(a) ReDiFine preserves the latent distribution.
-(b) Conditional and unconditional score differences do not surge in the $6$-th iteration, showing dramatic improvements from baselines.
-(c) ReDiFine maintains similar power spectrum to CFG scale $2.5$ in the $6$-th iteration.
-Latent analysis.
-Figure 8 demonstrates that ReDiFine can successfully alleviate the symptoms we observed in the latent space in Section 3.
-When we plot latent histograms, we can notice that ReDiFine preserves the original latent distribution at the 6th iteration as compared to other baselines with different CFG values.
-More importantly, the conditional-unconditional score differences (Diff) are kept at bay all the way until the 6th iteration, as opposed to the baseline CFG $7.5$, showing a severe discrepancy between the unconditional and conditional scores.
-Additionally, the power spectrum of latents in the $6$th iteration increases or decreases for CFG scales $1.0$ and $7.5$ while ReDiFine and CFG scale $2.5$ preserve density distribution.
-More analysis and comparison for all iterations are provided in Appendix F.1.
-(a) Autocorrelation.
-(b) Average power spectra.
-Figure 9: The fingerprints of the original training set and synthetic sets show clear differences, and ReDiFine produces fingerprints similar to CFG scale 2.5.
-(a) Autocorrelation of image fingerprints.
-Horizontal and vertical lines gradually disappear for CFG scales $1.0$ and $7.5$ while they are maintained for CFG scale $2.5$ and ReDiFine.
-(b) Average power spectra of images.
-Central regions are amplified or diminished for CFG scales $1.0$ and $7.5$, demonstrating low and high-frequency degradation.
-Frequency fingerprint.
-Several works [35, 36] aim to identify fingerprints of synthetic images.
-High-quality synthetic images from different generative models have clearly distinct fingerprints, showing the potential to be used for synthetic image detection.
-We analyze fingerprints of synthetic images for different CFG scales and iterations, and compare them to fingerprints of the original training set.
-Both power spectra and autocorrelation show clear differences between the original training set and synthetic images as shown in Figure 9.
-Moreover, fingerprints of synthetic images from ReDiFine are similar to those of images from CFG scale $2.5$, while other CFG scales ($1.0$ and $7.5$) make fingerprints different from the first iteration as iterations proceed.
-Horizontal and vertical lines in autocorrelation gradually disappear and central regions in power spectra vary for further iterations.
-Also, the varying central regions in power spectra imply that low frequency features increase and decrease for CFG scale $1.0$ and $7.5$, respectively, aligning with visual inspections.
-Generating images with fingerprints similar to those of the original real images can be an interesting future direction to reduce the degradation in the Chain of Diffusion.
-Additional analysis with power spectrum density of images is provided in Appendix F.2.
-6 Conclusion and Discussion
-Motivated by the prominent usage of diffusion models by artists and the trend of the Internet being filled with synthetic data, we investigate how image quality degrades as we feed diffusion-generated synthetic images to finetune another diffusion model.
-In our self-consuming chain of diffusion finetuning setting, we observe image degradation in all four datasets we tested.
-We reveal that CFG plays a crucial role in the trade-off between image reusability and its perceptual quality in the first iteration.
-We propose a novel idea of generating reusable iamges and propose the ReDiFine strategy that combines two small techniques (condition drop finetuning and CFG scheduling) that can be easily implemented in the diffusion finetuning pipeline with no requirement for additional parameter tuning.
+Figure 6: ReDiFine effectively mitigates collapse-FID trade-off across all four datasets.
+While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low collapse rate and FID at the same time (lower is better).
+Note that the differences in FID are relatively smaller than those in collapse rate, supporting the necessity to evaluate collapse rate in the Chain of Diffusion.
+ReDeFine Results.
+We present the generated images and quantitative metrics for ReDiFine.
+We use the initial CFG scale $s_{0}=7.5$, diffusion steps $T=30$, decay rate $\alpha=2$, and condition drop probability $0.2$ for all of our experiments, except in the robustness comparison and ablation study.
+While ReDiFine can be further optimized through hyperparameter tuning (e.g., changing condition drop probability), our goal is to demonstrate that ReDiFine robustly mitigates model collapse with these default parameter settings for all datasets.
+In Figure 5, we can clearly see that ReDiFine significantly improves the image quality at later iterations compared to the baseline (Figure 11(b)) for all four datasets.
+In addition to the visual comparison, Figure 6 quantitatively shows the collapse-FID trade-off of ReDiFine.
+In all four datasets, ReDiFine shows substantially lower collapse rate (y-axis), compared to the baseline case (CFG=7.5).
+Furthermore, the performance of ReDiFine is close to the optimal Pareto curve spanned by different CFG scales, achieving similar performance as the optimal CFG values across all four datasets, demonstrating its effectiveness as a universal solution for mitigating model collapse.
+In contrast, using a fixed CFG scale that works well for one dataset often fails on others: CFG $2.5$ is optimal for Pokemon but performs poorly for CelebA-1k and Butterfly, and CFG $1.5$ is optimal for CelebA but performs poorly for Pokemon and Kumapi.
+We also compare ReDiFine with real image mixing, which is suggested as a mitigation strategy in several previous papers [16, 3, 8, 23].
+In Figure 7, we can observe that even infusing the dataset with 90% real images, it shows severe image degradation within six iterations.
+By visual comparison, we can see that Pikachu images in Figure 5 have less degradation than those with real image mixing.
+Furthermore, applying ReDiFine to these mixed datasets, as shown in Figure 7, significantly improves image quality compared to the baseline.
+This demonstrates that ReDiFine is an effective solution for various settings and can be used alongside other mitigation strategies such as real data mixing.
+Figure 7: Comparison of ReDiFine and baseline (CFG scale 7.5) with training sets mixed with original and synthetic images from the previous iteration.
+While mixing real images as much as $90$% still results in model collapse for the baseline (as shown in the first row), the ability of ReDiFine to mitigate model collapse is even more highlighted by mixing real images (second row).
+Ablation study & Further analyses.
+We conducted an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine.
+We plot the results of using only condition drop and CFG scheduling in green triangles on Figure 6.
+In Pokemon, using only one of those strategies outperforms ReDiFine, but in all other datasets, using only one strategy shows higher collapse rate than ReDiFine.
+Especially in CelebA, using either one of them showed significantly worse performance than ReDiFine.
+These results suggest that combining both condition drop and CFG scheduling builds robustness to the method, making ReDiFine effective across all tested datasets.
+We conducted further analyses on ReDiFine, examining the distribution of latent features, the evolution of the norm of Diff, the power spectra using 2D Fourier transforms, and forensic fingerprints based on prior studies [45, 46].
+Our analysis shows that ReDiFine effectively preserves the latent distribution and the norms of Diff over six iterations, with forensic fingerprints closely resembling those of the optimal CFG case.
+Detailed results are provided in Appendix G.
+Reusable image generation as a responsible AI practice.
+In this section, we demonstrated that ReDiFine significantly slows the collapse rate without sacrificing image quality, similar to how fair classifiers reduce bias without compromising accuracy [47].
+Adopting ReDiFine is a responsible, environment-conscious practice to prevent polluting the world with images that might look visually appealing, but totally unusable to improve future AI models.
+6 Conclusion
+The influx of AI-generated data into the world is inevitable and training sets that consist of synthetic data will be part of the AI development pipeline.
+In this paper, we empirically and theoretically studied the scenario of finetuning a model with its own generated data, where a gradual degradation called “model collapse” happens.
+We (1) identify the most impactful factor through comprehensive empirical investigations, (2) develop a novel theoretical perspective inspired by statistical genetics to explain model collapse, and (3) propose an effective mitigation strategy that generates reusable images for future training.
 We started this paper with a question: can current AI models learn from their own output and improve themselves?
-This is not only an imminent question as synthetic data rapidly outnumbers human-generated data [37, 38], but also a long-term existential question as it relates to reaching AI singularity—the point where AI learns from its own output to continuously enhance its intelligence and far surpasses human intelligence.
 Our paper shows a glimpse that widely-used text-to-image models are not ready to improve from their own creation quite yet.
-However, we believe that there are many open questions on how we develop AI in a world dominated by AI data.
-We present one angle by emphasizing the importance of generating reusable data for future AI developers.
-Another pathway could be developing a new learning algorithm that can discern real and synthetic data and take a different training technique for each.
+While we presented one solution focused on generating reusable data, many open directions remain, such as developing algorithms that can distinguish between real and synthetic data and apply different learning techniques accordingly.
 Supplementary Material
@@
 A.1 Hyperparameters
-We use Stable Diffusion v1.5 [3] as a pretrained text-to-image diffusion model and apply LoRA [21] to fine-tune it at each iteration.
-We utilize ft-MSE [39] as a fixed VAE to project images into latent space.
-We augment images with horizontal flip, and no additional augmentation is used.
-Our implementation is based on [26], with the environment built on PyTorch v2.2.2 [40] and its accompanying torchvision 0.17.2, utilizing CUDA 12.4 on NVIDIA A-100 and L40S GPUs.
-All default hyperparameters used throughout our paper can be found in Table 2.
+We finetune Stable Diffusion v1.5 [24] using LoRA [25] at each iteration, with ft-MSE [48] as a fixed VAE to project images into latent space.
+Horizontal flip is the only image augmentation applied.
+Our implementation follows [26] and is built on PyTorch v2.2.2 [49], with torchvision 0.17.2, running on CUDA 12.4 using NVIDIA A-100 and L40S GPUs.
+All default hyperparameters are listed in Table 1.
 Table 1: Default hyperparameters used for the Chain of Diffusion.
@@
 A.2 Datasets
-We use four image datasets to demonstrate that the degradation occurs universally: Pokemon [22], Kumapi [23], Butterfly [24], and CelebA-1k [25] to investigate various domains including animation, handwriting, and real pictures.
+We use four image datasets to demonstrate the universal nature of degradation: Pokemon [27], Kumapi [28], Butterfly [29], and CelebA-1k [30], covering animation, handwriting, and real images.
 All images are resized to $512\times 512$ pixels.
-BLIP captioner [41] and Waifu Diffusion v1.4 tagger [42] are used to generate text prompts.
-Samples images and prompts can be found in Table 3.
-A.2.1 Pokemon
-The Pokemon dataset [22] contains $1008$ images corresponding to the Pokemon indexing from No.1 to No.1008.
-Prompts are constructed by concatenating captions from Waifu Diffusion tagger and BLIP captioner.
-We restrict the length of BLIP captions between $50$ and $75$.
-Table 3: Samples images and prompts for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets.
+Text prompts are generated using BLIP captioner [50] and Waifu Diffusion v1.4 tagger [51].
+Sample images and prompts can listed in Table 2.
+Pokemon.
+The Pokemon dataset [27] contains $1008$ images indexed by number, with prompts combining BLIP captions (length $50$-$75$ words) and Waifu Diffusion tagger.
+Table 2: Sample images and prompts for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets.
 Pokemon
@@
 [Uncaptioned image]
-A.2.2 CelebA-1k
-We subsample the first $1000$ images from CelebA dataset [25] to construct CelebA-1k dataset. Prompts are generated using BLIP captioner with length between $25$ and $50$.
-A.2.3 Kumapi
-The Kumapi dataset [23] consists of $391$ images with handwriting styles. Prompts are generated using Waifu tagger, and we manually modify some prompts that do not reflect the images well.
-A.2.4 Butterfly
-Butterfly dataset [24] consists of $832$ images with $8$ species categories.
-We combine captions generated using the BLIP captioner and descriptions for species from the dataset.
-The length of captions generated by the BLIP captioner is between $5$ and $75$.
+CelebA-1k.
+CelebA-1k is a subsample of $1000$ images from CelebA [30], with BLIP captions ($25$-$50$ words).
+Kumapi.
+Kumapi [28] includes $391$ handwriting-style images, with Waifu-generated prompts and manual adjustments.
+Butterfly.
+Butterfly [29] includes $832$ images across $8$ species, using BLIP captions (length $50$-$75$ words) and species descriptions.
 Appendix B Different CFG Scales
-This section presents images from Chains of Diffusion for different CFG scales on four datasets outlined above.
-Figure 10, 11, 12, and 13 are for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets, respectively.
-For each dataset, results for $5$ different CFG scales are provided to cover high, low, and medium CFG scales.
-We found $2.5$ to be the optimal medium CFG scale for Pokemon and Kumapi datasets, and $1.5$ for CelebA-1k and Butterfly datasets.
-With CFG scales lower than the optimal value, low-frequency degradations are observed throughout the four datasets we demonstrate.
-On the other hand, high-frequency degradations with saturated color and repetitive patterns are universally observed with CFG scales higher than the optimal value.
-Notably, the optimal value for Pokemon and Kumapi causes the Chain of Diffusion for CelebA-1k and Butterfly datasets to degrade severely.
-Figure 10: Chain of Diffusion with different CFG scales for Pokemon dataset. CFG scale $2.5$ is the optimal value we found.
-Figure 11: Chain of Diffusion with different CFG scales for CelebA-1k dataset. CFG scale $1.5$ is the optimal value we found.
-Figure 12: Chain of Diffusion with different CFG scales for Kumapi dataset. CFG scale $2.5$ is the optimal value we found.
-Figure 13: Chain of Diffusion with different CFG scales for Butterfly dataset. CFG scale $1.5$ is the optimal value we found.
-Appendix C Hyperparameter investigations to unveil the most impactful factor of degradation
-In this section, we present experiment results to reveal the most impactful factor of degradation observed in the Chain of Diffusion.
-We conduct extensive investigations to check the effects of each hyperparameter outlined in Table 1.
-Based on the default hyperparameter setting in Table 2, we change each candidate hyperparameter and investigate whether any of them can meaningfully mitigate or affect the degradation.
-CFG scale is $7.5$ for all cases in this section.
+This section presents images from Chain of Diffusion at various CFG scales for the four datasets.
+Figure 8, 9, 10, and 11 corresponds to Pokemon, CelebA-1k, Kumapi, and Butterfly, respectively.
+Each dataset shows results for five CFG scales, covering high, medium, and low values.
+The optimal medium CFG scale is $2.5$ for Pokemon and Kumapi, and $1.5$ for CelebA-1k and Butterfly.
+Lower-than-optimal CFG scales lead to low-frequency degradations, while higher-than-optimal scales result in high-frequency degradations with saturated colors and repetitive patterns.
+Notably, the optimal CFG for Pokemon and Kumapi causes severe degradation in CelebA-1k and Butterfly.
+Figure 8: Chain of Diffusion for Pokemon at various CFG scales. The optimal CFG scale is $2.5$.
+Figure 9: Chain of Diffusion for CelebA-1k at various CFG scales. The optimal CFG scale is $1.5$.
+Figure 10: Chain of Diffusion for Kumapi at various CFG scales. The optimal CFG scale is $2.5$.
+Figure 11: Chain of Diffusion for Butterfly at various CFG scales. The optimal CFG scale is $1.5$.
+Appendix C Hyperparameter investigations to unveil the most significant factor of degradation
+This section presents experimental results identifying the most significant factors contributing to degradation in the Chain of Diffusion.
+We systematically vary each hyperparameter from Figure 2, using the default settings from Table 1, to assess their impact on degradation.
+The CFG scale is fixed at $7.5$ for all cases.
 C.1 Training set size
-We used CelebA-1k dataset to investigate how the size of the training set (both $D_{0}$ and $D_{k}$) affects the degradation observed in the Chain of Diffusion.
-We chose this dataset as we can increase or decrease the training set size easily by subsampling from a larger original dataset.
-We changed the size of the training set to $100$, $250$, $500$, and $2000$.
-Degradations are observed regardless of the size of the training set, as shown in Figure 14.
-Degradations occur at earlier iterations when a smaller number of images are used, but it is hardly distinguished in $6$th iteration.
-We fix the number of parameter updates.
-Figure 14: Chain of Diffusion with varying training set size on CelebA-1k dataset.
-Degradations begin to occur faster with smaller sizes, but all of them results in unrecognizable degree of degradations in $6$th iteration.
+We used CelebA dataset [30] to examine how training set size (both $D_{0}$ and $D_{k}$) impacts degradation in the Chain of Diffusion.
+By subsampling, we adjusted the training set to $100$, $250$, $500$, and $2000$ images.
+Degradation occurs regardless of dataset size, as shown in Figure 12, but appears earlier with smaller sets.
+By the $6$th iteration, images degrade severely for all cases.
+The number of parameter updates was kept constant across all dataset sizes.
+Figure 12: Chain of Diffusion on CelebA dataset with varying training set sizes. Degradation occurs faster with smaller sets, but all result in severe degradation by the $6$th iteration.
 C.2 Number of images per prompt
-Generating more than one image per prompt to increase the diversity of the training set is a straightforward method to mitigate the degradation in the Chain of Diffusion.
-We conducted an experiment to investigate how increasing the training set size by generating more images affects the degradation.
-Figure 15 demonstrates that $5$ times larger images cannot mitigate the degradation.
-It seems the degradation is slowed down by one iteration ($6$th iteration for the default case and $5$th iteration for the target case look similar), but at the cost of high computations, which makes it an infeasible solution.
-Figure 15: Chain of Diffusion when more than one images are generated per prompt on Pokemon dataset.
-Increasing the training set size by generating more than one image per prompt cannot mitigate the degradation.
+Generating multiple images per prompt is a simple way to increase training set diversity and can be considered as a solution to mitigate degradation in the Chain of Diffusion.
+We tested this by generating $5$ times more images.
+As shown in Figure 13, while degradation is slightly delayed (by one iteration), it remains unmitigated, and the high computational cost makes this approach impractical.
+Figure 13: Chain of Diffusion on Pokemon dataset with multiple images generated per prompt. Increasing the training set size in this way does not mitigate degradation.
 C.3 Mixing real images to synthetic sets
-Many previous works suggested augmenting the training set with real images to mitigate the degradation observed in iterative training.
-We investigated if mixing images from the original training set to a synthetic set at each iteration can relieve the degradation.
-Images are randomly sampled at each iteration and replaced by corresponding real images in the original training set.
-Figure 16 and 17 demonstrate how the degradation in the Chain of Diffusion differs when $50\%$ and $90\%$ images at each iteration are replaced for Pokemon and CelebA-1k datasets, respectively.
-It is noteworthy that $10\%$ is far enough to cause the degradation, and $50\%$ rarely slows down the degradation.
-It is interesting that CelebA-1k dataset is much more vulnerable to degradation.
-Figure 16: Chain of Diffusion when real images are randomly sampled and replace synthetic images in every iteration on Pokemon dataset.
-$50\%$ of real images rarely slow down the degradation, and $10\%$ is enough to commence the degradation.
-Figure 17: Chain of Diffusion when real images are randomly sampled and replace synthetic images in every iteration on CelebA-10k dataset.
-The degradation is rarely slowed down even when only $10\%$ of synthetic images are used for finetuning.
-Natural images undergo more serious degradation compared to animation images in Figure 16.
+Many previous works suggest augmenting the training set with real images to mitigate degradation during iterative training.
+We investigated whether mixing images from the original training set into the synthetic set at each iteration could alleviate this issue.
+At each iteration, images are randomly replaced with corresponding real images.
+Figure 14 and 15 show how degradation in the Chain of Diffusion varies when $50$% and $90$% of images are replaced for Pokemon and CelebA-1k datasets, respectively.
+Notably, even $10$% synthetic images are sufficient to induce degradation, and $50$% replacement rarely slows it down.
+CelebA-1k dataset appears to be significantly more susceptible to degradation.
+Figure 14: Chain of Diffusion on Pokemon dataset with real images randomly replacing synthetic images at each iteration.
+A $50$% replacement rarely slows degradation, while $10$% synthetic images are sufficient to initiate it.
+Figure 15: Chain of Diffusion on CelebA-1k dataset with real images randomly replacing synthetic images at each iteration.
+A $50$% replacement rarely slows degradation, while $10$% synthetic images are sufficient to initiate it.
+It suffers from more severe degradation than Pokemon dataset as compared with Figure 14.
 C.4 Prompt Set
-We hypothesized that the descriptiveness of prompts can affect the degradation of the Chain of Diffusion.
-Accordingly, we examined different sets of prompts for Pokemon and CelebA-1k datasets.
-The default prompt set used for Pokemon dataset is the concatenations of Waifu and BLIP captions.
-BLIP captions are generated to be between $50$ and $75$.
-We investigated how Waifu prompt set alone and different lengths of BLIP captions can affect the degradation in Figure 18.
-It is interesting that slightly different styles of high-frequency degradation are observed and short prompt set can relieve repetitive patterns at the cost of decreased diversity.
-Different lengths of BLIP captions in CelebA-1k dataset demonstrate similar degrees of degradation, but insufficient or excessive descriptiveness of prompt makes images to be different from the original ones as shown in Figure 19.
-Figure 18: Chain of Diffusion when different prompts are used on Pokemon dataset.
-The default prompts are concatenations of BLIP caption of length between $50$ and $75$, and Waifu caption.
-We compare the Chain of Diffusion with Waifu captions, BLIP captions (used for default prompts), and short BLIP captions.
-Short captions are generated to have lengths smaller than $25$.
-Figure 19: Chain of Diffusion when different prompts are used on CelebA-1k dataset.
-The default prompts have lengths between $25$ and $50$.
-We compare the Chain of Diffusion with the lengths of prompts larger than $50$ (Long) and smaller than $25$ (Short).
+We hypothesized that the descriptiveness of prompts influences degradation in the Chain of Diffusion.
+We tested various prompt sets for Pokemon and CelebA-1k datasets.
+For Pokemon dataset, the default prompt set consists of concatenated Waifu and BLIP captions, with BLIP captions ranging from $50$ to $75$ words.
+Figure 16 shows how using only Waifu prompts and varying BLIP caption lengths affect degradation.
+Notably, different styles of high-frequency degradation were observed; shorter prompts reduced repetitive patterns but decreased diversity.
+In CelebA-1k dataset, varying BLIP caption lengths resulted in similar degradation levels, but prompts that were either insufficiently or excessively descriptive caused the images to deviate from the originals, as shown in Figure 17.
+Figure 16: Chain of Diffusion on Pokemon dataset with different prompts.
+The default prompts are concatenations of BLIP captions ($50$-$75$ words) and Waifu captions.
+We compare the Chain of Diffusion using default captions, Waifu captions, short (less than $25$ words) and long ($50$-$75$ words) BLIP captions.
+Figure 17: Chain of Diffusion on CelebA-1k dataset with different prompts.
+The default prompts range from $25$ to $50$ words.
+We compare the Chain of Diffusion using longer prompts (over $50$ words) and shorter prompts (under $25$ words).
 C.5 U-Net and Text-Encoder
-Figure 20 shows the Chains of Diffusion when one of U-Net or text encoder is finetuned.
-Similar degradation happens when the text encoder is not updated (second row), and the degradation pattern is different when the U-Net of Stable Diffusion is not updated since image qualities that models can generate do not change in this case.
-However the contents of images are not preserved as text encoder is updated.
-Figure 20: Chain of Diffusion when one of U-net and text encoder is finetuned on Pokemon dataset.
+Figure 18 illustrates the Chain of Diffusion with either the U-Net or text encoder finetuned.
+When the text encoder is not updated (second row), similar degradation occurs.
+However, the degradation pattern changes when the U-Net is not updated, as the model’s ability to generate images remains unchanged.
+In contrast, updating the text encoder results in a loss of image content preservation.
+Figure 18: Chain of Diffusion on Pokemon dataset with either the U-Net or text encoder finetuned.
 C.6 Number of diffusion steps
-We examined if the number of diffusion steps during generation is insufficient and causes the degradation. Figure 21 demonstrates that the number of diffusion steps cannot improve the Chain.
+We investigated whether an insufficient number of diffusion steps during generation contributes to degradation.
+Figure 19 shows that increasing the number of diffusion steps does not enhance the Chain of Diffusion.
 C.7 Number of Epochs
-We also tested if the training is insufficient or excessive for our default setting.
-However, images from the first iterations look similar in quality, as shown in Figure 21, and the Chains result in similar degradations.
-We finetune models for $100$ epochs as the default setting since loss values still decrease in $50$ epochs.
-Figure 21: Chain of Diffusion when the number of diffusion steps and training epochs differ on Pokemon dataset.
-Both more diffusion steps and different training epochs cannot mitigate the degradation, and similar patterns are observed.
+We also examined whether insufficient or excessive training affects our default setting.
+As shown in Figure 19, images from the initial iterations exhibit similar quality, resulting in comparable degradations.
+We set the default finetuning to $100$ epochs since loss values continue to decrease after $50$ epochs.
+Figure 19: Chain of Diffusion on Pokemon dataset with varying diffusion steps and training epochs.
+Both increased diffusion steps and differing training epochs fail to mitigate degradation, resulting in similar patterns.
 C.8 Learning Rate
-Similarly, we examined in Figure 22 if finetuning is properly done with our default setting.
-We multiplied the learning rate for U-Net and text encoder by $10$ and $0.1$.
-Images of the first iterations indicate that the default values are appropriate for finetuning, and the degradation constantly occurs for different learning rate.
-Figure 22: Chain of Diffusion when learning rate differs and Gaussian noise is added to the original training set on Pokemon dataset.
+Similarly, we assessed finetuning adequacy in Figure 20 by adjusting the learning rates for the U-Net and text encoder by x10101010 and x0.10.10.10.1.
+Images from the initial iterations show that the default values are suitable for finetuning.
+Although the styles are different, degradation consistently occurs across different learning rates.
+Figure 20: Chain of Diffusion on Pokemon dataset with varying learning rates and added Gaussian noise to the original training set.
 C.9 CLIP Skip
-We examined if CLIP skip hyperparameter can affect the degradation.
-CLIP skip determines what intermediate feature from CLIP text encoder is used as text embedding for conditional generation.
-Smaller value means the feature closer to the output of CLIP text encoder is used, and larger values for features closer to the input text.
-Figure 23 shows that this hyperparameter rarely changes the degradation patterns.
-Figure 23: Chain of Diffusion when different CLIP skip hyperparameter is used on Pokemon dataset. CLIP skip hyperparameter has negligible effect to the degradation.
+We investigated the impact of the CLIP skip hyperparameter on degradation.
+The CLIP skip determines which intermediate feature from the CLIP text encoder is used as the text embedding for conditional generation, with smaller values selecting features closer to the output and larger values selecting those nearer to the input text.
+As shown in Figure 21, this hyperparameter has minimal effect on degradation patterns.
+Figure 21: Chain of Diffusion on Pokemon dataset using different CLIP skip hyperparameters. The CLIP skip hyperparameter shows a negligible effect on degradation.
 C.10 Adding Gaussian noise to the original training set
-We inspected if there are any differences between real and synthetic images, and if they can be cause the degradation.
-Random Gaussian noise is added to the original training set $D_{0}$.
-Figure 22 shows that the characteristics of the original training set have some effects, but the degradation still happens.
-This aligns with our observations that the degradations are universal to real, animation, and hand-written images.
+We examined whether differences between real and synthetic images contribute to degradation by adding random Gaussian noise to the original training set $D_{0}$.
+Figure 20 illustrates that while the characteristics of the original training set have some effect, degradation still occurs.
+This supports our findings that degradations are universal across real, animation, and handwritten images.
 C.11 Stable Diffusion XL
-We investigated how image degradation occurs for different Stable Diffusion models.
-We note here that optimal hyperparameters for finetuning SDXL using LoRA have not been investigated enough yet, so we use the same hyperparameters to Stable Diffusion v1.5 and it can be suboptimal.
-Due to the space complexity, we reduce batch size to $2$.
-We also use the same resolution $512\times 512$ for this model as the first iteration images are impressively high-quality.
+We investigated image degradation for a different Stable Diffusion model, noting that the optimal hyperparameters for finetuning SDXL using LoRA are not well established.
+Consequently, we applied the same hyperparameters used for Stable Diffusion v1.5, which may be suboptimal.
+To manage space complexity, we reduced the batch size to $2$ and maintained a resolution of $512\times 512$, as the first iteration images exhibit impressive quality.
+Results are presented in Figure 22.
 Figure 22: Chain of Diffusion of SDXL on Pokemon dataset.
-Optimal hyperparameters for finetuning SDXL have not been investigated enough yet, so our experiments are largely based on hyperparameters for Stable Diffusion v1.5.
-We note that images for Pikachu are provided to align with other cases, but SDXL has a huge bias in generating Pikachu well while it fails for other indices.
-Nonetheless, there are clear changes with different CFG scales.
-Appendix D ReDiFine
-Figure 25, 26, 27, and 28 show how robust ReDiFine is to different CFG scales. It successfully mitigates the high-frequency degradations for an extensive range of CFG scales.
+Due to insufficient investigation into optimal hyperparameters for finetuning SDXL, our experiments largely rely on those from Stable Diffusion v1.5.
+Appendix D Quantitative Trait Modeling
+D.1 Proof of Theorem 1
+Proof. The difference between the successive means in quantitative trait modeling is given by:
+$\Delta\mu_{t}=\mu_{t+1}-\mu_{t}=h_{t}^{2}(\mu_{t}^{\prime}-\mu_{t}).$ (6)
+When Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$ is truncated on both sides with $r_{1}$ and $r_{2}$ ratios, the mean and variance of truncated Gaussian distribution are expressed as:
+$\mu_{trun}=\mu-\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}\sigma,$ (7)
+$\sigma_{trun}^{2}=(1-\frac{\beta\varphi(\beta)-\alpha\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}-(\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)})^{2})\sigma^{2},$ (8)
+where $\varphi$ and $\Phi$ are the probability density function (PDF) and cumulative distribution function (CDF) of the standard normal distribution, and $\alpha=\Phi^{-1}(r_{1})$ and $\beta=\Phi^{-1}(1-r_{2})$.
+Accordingly, given $\mu=\mu_{t}$ and $\sigma^{2}=\sigma_{P,t}^{2}$ with $\mu_{t}^{\prime}=\mu_{trun}$ and $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}=\sigma_{trun}^{2}$, we have
+$\Delta\mu_{t}=h_{t}^{2}(\mu_{t}^{\prime}-\mu_{t})=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}}c_{1}\sigma_{P,t}=c_{1}\frac{\sigma_{P,t-1}^{\prime 2}}{\sigma_{P,t}^{2}}\sigma_{P,t}=c_{1}c_{2}\frac{\sigma_{P,t-1}^{2}}{\sigma_{P,t}^{2}}\sigma_{P,t},$ (9)
+where $c_{1}=|\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}|$ and $c_{2}=1-\frac{\beta\varphi(\beta)-\alpha\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}-(\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)})^{2}$ when $r_{1}>r_{2}$.
+On the other hand, the mean phenotypes decreases when $r_{1}<r_{2}$ as:
+$\Delta\mu_{t}=-c_{1}c_{2}\frac{\sigma_{P,t-1}^{2}}{\sigma_{P,t}^{2}}\sigma_{P,t}.$ (10)
+Furthermore, the phenotype variance converges over time as:
+$\sigma_{P,t}^{2}=\sigma_{G,t}^{2}+\sigma_{E}^{2}=\sigma_{P,t-1}^{\prime 2}+\sigma_{E}^{2}=c_{2}\sigma_{P,t-1}^{2}+\sigma_{E}^{2}.$ (11)
+$\sigma_{P,t}^{2}-\frac{\sigma_{E}^{2}}{1-c_{2}}=c_{2}(\sigma_{P,t-1}^{2}-\frac{\sigma_{E}^{2}}{1-c_{2}})=\cdots=c_{2}^{t}(\sigma_{P,0}^{2}-\frac{\sigma_{E}^{2}}{1-c_{2}}).$ (12)
+As a result, the phenotype variance converges to $\frac{\sigma_{E}^{2}}{1-c_{2}}$ for $10<c_{2}<$ 1 (the variance of truncated distribution is smaller than the variance of the original distribution), and the mean asymptotically increases (decreases) by $\frac{c_{1}c_{2}}{\sqrt{1-c_{2}}}\sigma_{E}$ per iteration when $r_{1}>r_{2}$ ($r_{2}>r_{1}$).
+$\square$
+D.2 Simulation setup
+Our simulation aims to demonstrate that our experimental results can be modeled using quantitative trait modeling.
+We show that the radial sum of power spectra of images is one of the phenotypes explained by our theoretical analysis.
+D.2.1 Computing the radial sum of power spectra.
+The power spectra of images are computed as the square of the magnitude of the 2D Fourier transform.
+Here, we demonstrate how to compute the radial sum of power spectra in order to use the sum of radial power spectra above a certain threshold.
+Given a set of images $\{I_{i}\}$ where $i\in\{1,2,\cdots,N\}$, the 2D Discrete Fourier Transform (DFT) of an image $I_{i}$ of size $M\times N$ is computed as:
+$F_{i}(u,v)=\sum^{M-1}_{x=0}\sum^{N-1}_{y=0}I_{i}(x,y)e^{-2\pi j(\frac{ux}{M}+\frac{vy}{N})}$ (13)
+where $F_{i}(u,v)$ represents the frequency component at coordinates $(u,v)$.
+The power spectrum of an image $I_{i}$ at $(u,v)$ is the square of the magnitude of its Fourier transform $P_{i}(u,v)=|F_{i}(u,v)|^{2}$.
+We compute the radial sum of the power spectra using a norm of each frequency component $(u,v)$ as $r(u,v)=\sqrt{(\frac{u}{M})^{2}+(\frac{v}{N})^{2}}$.
+Given the threshold frequency $\tau$, we compute the radial sum of the high-frequency power spectra of an image $I_{i}$ as
+$S_{i}=\sum^{M-1}_{u=0}\sum^{N-1}_{v=0}P_{i}(u,v)\cdot{\rm I}(r(u,v)>\tau)$ (14)
+where ${\rm I}(\cdot)$ is an indicator function.
+We compute the sum of high-frequency components because low-frequency components tend to be noisy and use the threshold frequency of $0.02$.
+Then, the total sum of power spectra for all images can be written as:
+$S=\sum^{N}_{i=0}S_{i}=\sum^{N}_{i=0}\sum^{M-1}_{u=0}\sum^{N-1}_{v=0}P_{i}(u,v)\cdot{\rm I}(r(u,v)>\tau).$ (15)
+Practically, we shift the Fourier transform maps so that the frequency norm larger than $\frac{1}{\sqrt{2}}$ is considered to be in the opposite direction.
+Code for detailed implementation comes from the official code of [45].
+Figure 23: Results comparing the simulation for mutations and power spectra of images generated by ReDiFine (with CFG scale $7.5$).
+Power spectra and simulation are plotted in solid and dotted lines, respectively.
+Our modifications to heritability and selection process successfully demonstrate changes occur to images by ReDiFine.
+D.2.2 Simulation parameters matching different CFG scales.
+We simulate different selection strategies using truncation ratios $r_{1}$ and $r_{2}$.
+To model high, medium, and low CFG scales from our experiments, we apply $(0.025,0.675)$, $(0.5,0.09)$, and $(0.95,0.0002)$ for $(r_{1},r_{2})$, respectively, which correspond to CFG scales of $7.5$, $2.5$, and $1.0$.
+For initial values in the simulation, we compute the mean ($0.027$) and standard deviation ($0.056$) from the original training set (iteration 00), using them as the initial values for mean and genetic standard deviation for our simulation.
+We set the environmental standard deviation to $0.25$.
+D.2.3 ReDiFine and mutations.
+ReDiFine combines condition drop finetuning and CFG scheduling, inspired by the mutation mechanism that compensates for distribution shift and maintains genetic distributions in population genetics.
+We apply two modifications to our theoretical analysis of Section 4—adding mutation variance to heritability and smoothing truncations—to simulate the effects of ReDiFine in the Chain of Diffusion.
+Specifically, we add the mutation standard deviation $\sigma_{M}$ of $0.1$ to heritability as:
+$h_{t}^{2}=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}+\sigma_{M}^{2}}=\frac{\sigma_{G,t}^{2}}{\sigma_{G,t}^{2}+\sigma_{E}^{2}+\sigma_{M}^{2}},$ (16)
+representing the randomness added to each iteration due to mutations.
+This influences the effects of previous iteration to current iteration, which reflect finetuning.
+Moreover, we apply exponential tails to truncations instead of cut-off thresholds where samples outside the truncation area can be randomly selected with exponential distribution ($e^{-\alpha d(x)}$ where $d(x)$ is a distance to truncation zone and we use $\alpha=0.1$).
+This modified selection simulates the effect of CFG scheduling during image generation.
+Figure 23 shows the effects of two modifications to our simulation results, and they closely align with the power spectra of images generated by ReDiFine.
+This demonstrates that the effects of ReDiFine can be understood as interference similar to mutations in population genetics.
+This suggests further research on model collapse motivated from other fields like biology.
+Appendix E ReDiFine
+Figure 24, 25, 26, and 27 show how robust ReDiFine is to different CFG scales. It successfully mitigates the high-frequency degradations for a wide range of CFG scales.
 Figure 24: Chain of Diffusion of ReDiFine with different CFG scales on Pokemon dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
@@
 Figure 27: Chain of Diffusion of ReDiFine with different CFG scales on Butterfly dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
-In addition to visual inspections for images generated using our ReDiFine, we also compare the quantitative results of ReDiFine to baselines with different CFG scales on FID, CLIP, and SFD.
-Results are shown in Figure 29.
-For different datasets and metrics, ReDiFine achieves comparable performance to the optimal CFG scale ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly).
+In addition to visual inspections for images generated using ReDiFine, we compare the quantitative results of ReDiFine to baselines with different CFG scales across FID and CLIP score.
+Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
+${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (17)
+to evaluate each iteration $k$.
+SFD overcomes the problem of FID being sensitive to the number of images to compare.
+DiNOv2 features are used to compute SFD.
+The results, shown in Figure 28, demonstrate that ReDiFine achieves performance comparable to the optimal CFG scales ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly) across different datasets and metrics.
 (a) Pokemon, CLIP.
@@
 Figure 28: Quantitative results of ReDiFine and baselines (different CFG scales).
-Appendix E Ablation study
+Appendix F Ablation study
 This section provides an ablation study to understand how condition drop finetuning and CFG scheduling contribute to the success of ReDiFine.
-E.1 Condition drop finetuning
+F.1 Condition drop finetuning
 We conducted an ablation study to understand how the probability of dropping text embedding during finetuning affects the image quality in the Chain of Diffusion.
-Based on the probability used for training Stable Diffusion ($0.1$ or $0.2$), we tested $0.1$, $0.2$, and $0.4$.
-For both Pokemon and CelebA-1k datasets, a probability of 0.2 seems to work the best, as shown in Figure 30 and Figure 31, respectively.
+We examine $0.1$, $0.2$, and $0.4$ as Stable Diffusion is trained using $0.1$ or $0.2$.
+For both Pokemon and CelebA-1k datasets, a probability of 0.2 works the best, as shown in Figure 29 and Figure 30, respectively.
 Interestingly, condition drop finetuning helps to mitigate the color saturation problem, but its effect decreases with a higher probability.
@@
 Figure 30: Chain of Diffusion with condition drop finetuning on CelebA-1k dataset.
-E.2 CFG scheduling
-We also tested how different decreasing strategies for CFG scale affect image degradation in the Chain of Diffusion.
-We used two different exponential orders for exponential decreasing, and also compared a linear decreasing strategy.
-Figure 32 shows that CFG scheduling is effective for Pokemon dataset, generating high-quality images comparable to those of ReDiFine.
-However, it fails to improve image quality on CelebA-1k dataset, as shown in Figure 33.
-This demonstrates the necessity of condition drop finetuning to improve the Chain of Diffusion universally for various datasets.
+F.2 CFG scheduling
+We also evaluated how different CFG scale decreasing strategies impact image degradation in the Chain of Diffusion.
+We experimented with two different exponential decay rates and compared them with a linear decreasing strategy.
+Figure 31 demonstrates that CFG scheduling is effective for Pokemon dataset, generating high-quality images comparable to those generated by ReDiFine.
+However, as shown in Figure 32, it fails to enhance image quality on CelebA-1k dataset.
+This highlights the necessity of condition drop finetuning for achieving universal improvements in the Chain of Diffusion across various datasets.
 Figure 31: Chain of Diffusion with CFG scheduling on Pokemon dataset.
 Figure 32: Chain of Diffusion with CFG scheduling on CelebA-1k dataset.
-Appendix F Analysis
-In this section, we provide various analyses of images generated through the Chain of Diffusion.
-Specifically, we provide frequency analysis of latents during diffusion steps and fingerprints of synthetic images for forensic analysis.
-F.1 Latent analysis
-(a) Histogram of latent vectors.
-(b) Diffs.
-(c) Power spectrum density.
-Figure 34: Fingerprints of the original training set and synthetic sets for Pokemon dataset.
-There are clear differences between fingerprints of natural images (Original) and synthetic images.
-ReDiFine makes similar fingerprints to CFG scale $2.5$.
-(a) Autocorrelation.
-Horizontal and vertical lines gradually disappear for CFG scales $1.0$ and $7.5$ while they are maintained for CFG scale $2.5$ and ReDiFine.
-(b) Average power spectra.
-Central regions are amplified or diminished for CFG scales $1.0$ and $7.5$, demonstrating low and high-frequency degradation.
-Figure 34 demonstrates how latent distribution evolves for different iterations and diffusion steps on Pokemon dataset.
-First, the histogram of the final latent vectors before decoding is shown in Figure 3434(a).
-For a CFG of $1.0$, the latent distribution quickly evolves into a Gaussian-like shape, with its variance decreasing over iterations.
-This behavior aligns with previous work [5, 12, 19], which theoretically predicted that the self-consuming loop would crop out the tails of the distribution, reducing output diversity until it collapses to a single mode.
-We believe that these narrowing distributions in the latent space lead to blurrier and more homogeneous outputs in the image space.
-The evolution of the latent distribution is completely opposite for $7.5$, creating longer tails and approaching a more uniform distribution over space.
-For a CFG scale of $2.5$, which has the best reusability among the three, the latent distribution is better preserved over iterations.
-ReDiFine better preserves the histogram from the first to the last iteration, achieving both first iteration fidelity and reusability.
+Appendix G Analysis
+In this section, we present a series of analyses of images generated through the Chain of Diffusion.
+Specifically, we examine the distribution of latent values and the differences between conditional and unconditional scores.
+Additionally, we analyze the power spectra of the images using 2D Fourier transforms and explore fingerprints through forensic analysis [45, 46].
+(a) Probability density function (PDF) of values of latent vectors at the last diffusion step.
+(b) Norm of differences between conditional and unconditional scores during diffusion steps.
+Figure 33: Histogram of latent values and Diffs during diffusion steps for Pokemon dataset.
+(a) Latent distribution shrinks over iteration for low CFG and expands with high CFG.
+Larger values in latent vectors are more likely to occur with high CFG, gradually increasing the tail of the distribution.
+(b) Differences between conditional and unconditional scores increase as the training set is more degraded.
+Especially, high differences in the later diffusion steps can be a cause of high-frequency degradation.
+G.1 Latent analysis
+Figure $3$(a) illustrates how the distribution of latent values evolves across different iterations.
+The histograms show the final latent vectors before decoding into pixel space, comparing various CFG scales and ReDiFine.
+For a CFG of $1.0$, the latent distribution rapidly converges into a Gaussian-like shape, with its variance shrinking over iterations.
+This behavior is consistent with previous work [3, 16, 11], which theoretically predicted that the self-consuming loop progressively trims the tails of the distribution, reducing output diversity until it collapses to a single mode.
+We hypothesize that this narrowing in the latent space leads to blurrier, more homogeneous outputs in pixel space.
+Conversely, at a CFG scale of $7.5$, the latent distribution develops longer tails and tends toward a more uniform spread across space.
+A CFG scale of $2.5$, which demonstrates the best reusability among the three, better preserves the latent distribution over iterations.
+ReDiFine further enhances this preservation, maintaining the histogram from the first to the last iteration, thus achieving both high fidelity in the first iteration and better reusability.
+G.2 Differences between conditional and unconditional scores
 […]
-Figure 3434(c) demonstrates how the power spectrum density is maintained by ReDiFine.
-While high and low CFG scales cannot preserve the power spectrum of the first iteration as the Chain of Diffusion proceeds, ReDiFine can maintain the density distribution during iterations.
-This prevents images from being saturated to high or low-frequency degradation.
-Moreover, the power spectrum is more similar to the original training set compared to CFG scale $2.5$.
-In summary, it is shown that ReDiFine can successfully alleviate the symptoms we observed in the latent space for baselines in Section 3.
-We can notice that ReDiFine preserves the original latent distribution at the $6$th iteration as compared to other baselines with different CFG values.
-More importantly, the conditional-unconditional score differences (Diff) are kept at bay all the way until the 6th iteration, as opposed to the baseline CFG $7.5$, showing a severe discrepancy between the unconditional and conditional scores.
-F.2 Fingerprints for forensic analysis
+G.3 Power spectra of 2D Fourier transforms
+Figure 34 demonstrates the radial and angular spectrum power density of both the original and synthetic images.
+It is evident that ReDiFine closely maintains the radial spectrum power density of the original training set, whereas even a CFG scale $2.5$ falls short.
+Additionally, ReDiFine demonstrates stable angular spectra throughout the Chain of Diffusion, even though they differ from those of the original training set.
+Pokemon dataset is used for power spectra analysis.
 (a) Radial spectrum power density.
@@
 Power density of generated images by ReDiFine remains during the iterations while baselines cannot maintain angular distribution.
-Fingerprints of synthetic images have distinct characteristics that differ according to the types of generative models, architectures, and training images.
-This provides unique patterns that can be used for synthetic image detection.
-Figure 9 shows how fingerprints of the original training set and synthetic images differ and how fingerprints of synthetic images change in the Chain of Diffusion.
-We use $1008$ images from Pokemon dataset.
-Following visual inspection, ReDiFine makes similar fingerprints to CFG scale $2.5$.
-Moreover, fingerprints for ReDiFine and baseline with CFG scale $2.5$ are maintained through the Chain of Diffusion.
-On the other hand, low or high CFG scales remain different patterns to the synthetic images as iterations proceed.
-Fingerprints gradually change for low or high CFG scales, showing the effects of high or low-frequency degradation.
-Figure 35 demonstrates the radial and angular spectrum power density of the original and synthetic images.
-It is clearly observed that ReDiFine maintains a similar radial spectrum power density compared to the original training set, while even CFG scale $2.5$ fails.
-Moreover, it shows stable angular spectrums during the Chain of Diffusion despite it being distinct from that of the original training set.
+G.4 Fingerprints for forensic analysis
+(a) Autocorrelation.
+(b) Average power spectra.
+Figure 35: The fingerprints of the original training set and synthetic sets show clear differences, and ReDiFine produces fingerprints similar to CFG scale 2.5.
+(a) Autocorrelation of image fingerprints.
+Horizontal and vertical lines gradually disappear for CFG scales $1.0$ and $7.5$ while they are maintained for CFG scale $2.5$ and ReDiFine.
+(b) Average power spectra of images.
+Central regions are amplified or diminished for CFG scales $1.0$ and $7.5$, demonstrating low and high-frequency degradation.
+Several works [45, 46] aim to identify fingerprints of synthetic images.
+High-quality synthetic images from different generative models have clearly distinct fingerprints, showing the potential to be used for synthetic image detection.
+We analyze fingerprints of synthetic images for different CFG scales and iterations, and compare them to fingerprints of the original training set.
+Both autocorrelation and average power spectra show clear differences between the original training set and synthetic images, as shown in Figure 35.
+Moreover, how the fingerprints of synthetic images evolve throughout the Chain of Diffusion differ for ReDiFine and different CFG scales.
+Specifically, fingerprints of synthetic images from ReDiFine are similar to those of images from CFG scale $2.5$, while other CFG scales ($1.0$ and $7.5$) make fingerprints different from the first iteration as iterations proceed.
+Horizontal and vertical lines in autocorrelation gradually disappear and central regions in power spectra vary for further iterations.
+Also, the varying central regions in power spectra imply that low frequency features increase and decrease for CFG scale $1.0$ and $7.5$, respectively, aligning with visual inspections.
+Generating images with fingerprints similar to those of the original real images can be an interesting future direction to reduce the degradation in the Chain of Diffusion.
+Pokemon dataset is used for fingerprint analysis.

### v3

@@
 Abstract
-The success of generative models has reached a unique threshold where their outputs are indistinguishable from real data, leading to the inevitable contamination of future data collection pipelines with synthetic data.
-While their potential to generate infinite samples initially offers promise for reducing data collection costs and addressing challenges in data-scarce fields, the severe degradation in performance has been observed when iterative loops of training and generation occur—known as “model collapse.” This paper explores a practical scenario in which a pretrained text-to-image diffusion model is finetuned using synthetic images generated from a previous iteration, a process we refer to as the “Chain of Diffusion.” We first demonstrate the significant degradation in image quality caused by this iterative process and identify the key factor driving this decline through rigorous empirical investigations.
-Drawing an analogy between the Chain of Diffusion and biological evolution, we then introduce a novel theoretical analysis based on quantitative trait modeling.
+Model collapse, the severe degradation of generative models when iteratively trained on their own outputs, has gained significant attention in recent years.
+This paper examines Chain of Diffusion, where a pretrained text-to-image diffusion model is finetuned on its own generated images.
+We demonstrate that severe image quality degradation was universal and identify CFG scale as the key factor impacting this model collapse.
+Drawing on an analogy between the Chain of Diffusion and biological evolution, we then introduce a novel theoretical analysis based on quantitative trait modeling from statistical genetics.
 Our theoretical analysis aligns with empirical observations of the generated images in the Chain of Diffusion.
 Finally, we propose Reusable Diffusion Finetuning (ReDiFine), a simple yet effective strategy inspired by genetic mutations.
-ReDiFine mitigates model collapse without requiring any hyperparameter tuning, making it a plug-and-play solution for reusable image generation.
+It operates robustly across various scenarios without requiring any hyperparameter tuning, making it a plug-and-play solution for reusable image generation.
 1 Introduction
 Can state-of-the-art AI models learn from their own outputs and improve themselves?
-As generative AI models (e.g., GPT, Diffusion) now churn out uncountable synthetic texts and images, this question piqued the curiosity of many researchers in the past couple of years.
-While some show positive results of self-improving [1, 2], most report an undesirable “model collapse”—a phenomenon where a model’s performance degrades when it undergoes multiple cycles of training with the self-generated data [3, 4, 5, 6, 7, 8, 9, 10].
-When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text with many repetitions [11], and its linguistic diversity declines rapidly [12, 13]; image models also show quality degradation [14, 15] and loss of diversity [16, 17].
+As generative AI models (e.g., GPT, Diffusion) now churn out uncountable synthetic texts and images, this question piqued curiosity from many researchers in the past couple of years.
+While some show positive results of self-improving [1, 2], most report an undesirable “model collapse”—a phenomenon where a model’s performance degrades when it goes through multiple cycles of training with the self-generated data [3, 4, 5, 6, 7, 8, 9, 10].
+When large language models (LLMs) are trained with their own outputs, it begins to produce low-quality text that has a lot of repetitions [11], and its linguistic diversity declines rapidly [12, 13]; image models also show quality degradation [14, 15] and loss of diversity [16, 17].
 The goal of this paper is to investigate model collapse in the practical scenario of finetuning pretrained text-to-image diffusion models.
-An end user of text-to-image diffusion models often wants to finetune the latest model to generate images with a very specific style (e.g., creating new characters in the style of Pokémon).
-In fact, hundreds of new finetuned diffusion models are regularly uploaded on platforms like CivitAI [^1] https://civitai.com/, each designed to produce different styles of images.
+End users often finetune the latest model to generate images with a specific style (e.g., creating characters in the style of Pokémon).
+In fact, hundreds of new finetuned models trained on different styles and tons of generated images are regularly uploaded on platforms like CivitAI [^1] https://civitai.com/.
 When users scrape the Internet to collect images in the style they want, it becomes almost inevitable that synthetic images will be included in their datasets.
-This is because the number of real images is limited, while synthetic images can be generated in massive quantities and dominate online sources.
-As a result, users will feel compelled to include more synthetic images in their datasets to meet the data demands of these ever-growing, data-hungry models.
-To this end, we conduct a thorough investigation into how various hyperparameters commonly used during diffusion finetuning (e.g., learning rate, diffusion steps, prompts) impact model collapse.
-From our extensive empirical analysis, we make a crucial observation: the classifier-free guidance (CFG) scale is the most significant factor influencing the rate of model collapse.
-Moreover, we observe a fascinating phenomenon that the direction of degradation varies with different CFG values—low CFG results in low-frequency degradation in images, while high CFG leads to high-frequency degradation.
-The critical role of CFG in determining both the rate and direction of model collapse is a novel insight previously unrecognized in the literature.
-To gain a deeper theoretical understanding of this phenomenon, we draw upon the concept of quantitative trait modeling from statistical genetics.
-Unlike existing studies that attribute model collapse to limited sample sizes and conclude that it leads to zero variance [16, 3, 6], we propose that the underlying cause is a truncation-based selection process modulated by CFG.
-This theoretical model accurately describes our experimental observations.
+To this end, we conduct a thorough empirical investigation into how various hyperparameters commonly used during diffusion finetuning (e.g., learning rate, diffusion steps, prompts) impact model collapse.
+We make a crucial observation: the classifier-free guidance (CFG) scale is the most impactful factor affecting the rate of model collapse.
+Moreover, we observe a fascinating phenomenon that the direction of degradation varies—low CFG results in low-frequency degradation in images, while high CFG leads to high-frequency degradation.
+The critical role of CFG in both the rate and direction of model collapse is a novel insight previously unrecognized in the literature.
+To gain a deeper theoretical understanding, we draw upon the concept of quantitative trait modeling from statistical genetics.
+Unlike existing studies that attribute model collapse to limited sample sizes and conclude it leads to zero variance [3, 6, 16], we show that a truncation-based selection results in mean drift of quantitative features.
+We found the high-frequency power spectra of generated images as one of the features that accurately follow our theoretical modeling.
 Moreover, we show that other theoretical work on model collapse can also be connected to statistical genetics models, such as random genetic drift and its variants [18, 19, 20, 21].
 This fresh angle of drawing parallels with statistical genetics offers a new framework to understand the mechanism of model collapse.
-Finally, inspired by our empirical results and theoretical insights, we introduce the Reusable Diffusion Finetuning (ReDiFine) method.
-Our experiments show that selecting an appropriate CFG can significantly slow down model collapse.
-However, finding the optimal CFG is computationally expensive, requiring numerous iterations of finetuning.
-To address this, we propose ReDiFine as a plug-and-play solution that achieves a comparable reduction in collapse rate without the need for any hyperparameter tuning.
-The default ReDiFine setting effectively mitigates model collapse across all four datasets we tested.
-Moreover, ReDiFine is effective at mitigating model collapse not only in a worst-case scenario of having a fully synthetic dataset, but also in a more practical case, where we have a mix of real and synthetic images in the training set.
-Our contributions can be summarized as follows:
-  • We performed thorough and extensive empirical investigations into model collapse when we finetune a diffusion model with diffusion-generated images.
-We tested an exhaustive list of experimental parameters on four datasets (two digital art & two natural images) to identify what affects model collapse.
-Our analysis reveals that CFG scale is the most influential factor that not only controls the rate of model collapse but also dictates the type of image degradation (Section 3).
+While we have shown the crucial role of CFG in mitigating model collapse both empirically and theoretically, searching for the optimal CFG scale over iterations is computationally expensive and cumbersome to implement in practice.
+We propose a simple solution that bypasses CFG tuning, called Reusable Diffusion Finetuning (ReDiFine).
+It introduces small modifications in the finetuning and generation—condition drop and CFG scheduling—and we demonstrate that ReDiFine achieves performance comparable to the optimal CFG scale across all hyperparameter settings we tested.
+Our contributions can be summarized as follow:
+  • We conducted a comprehensive investigation of model collapse when finetuning a diffusion model on its own outputs, testing a wide range of parameters on four datasets (two digital art and two natural images).
+Our results show that CFG scale is the most critical factor, governing both the rate of model collapse and the type of image degradation (Section 3).
   • We provide a novel theoretical analysis of model collapse based on quantitative trait modeling that can accurately predict how power spectra of generated images evolve over iterations (Section 4).
-  • We propose a simple yet effective strategy to mitigate model collapse, named ReDiFine, which combines condition drop finetuning and CFG scheduling.
-Across all four datasets, ReDiFine successfully generates more “reusable” images, which can be reused to finetune a diffusion model without causing a severe model collapse (Section 5).
+  • We propose ReDiFine, a simple yet effective strategy to achieve a near-optimal reusability-fidelity trade-off without any hyperparameter tuning.
+By combining condition drop finetuning and CFG scheduling, ReDiFine operates robustly across various scenarios—such as synthetic-real mixed datasets and varying initial CFG settings—across all four datasets we evaluated (Section 5).
 2 Related Work
 The self-consuming training loop and the associated phenomenon known as “model collapse” have become significant areas of study in the past two years [3, 4, 5, 7, 8, 9, 10, 12, 13, 14, 15, 16].
-Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [6], has been observed in both language and image generative models.
+Model collapse, defined as “a degenerative process affecting generations of learned generative models, where generated data end up polluting the training set of the next generation of models” in [6], has been observed in both language and image generative models [^2] The definition of model collapse is expanding [22], but we focus on a fully self-consuming training loop [16]..
 Empirical studies on LLMs [13] reveal that linguistic diversity collapses, especially in high-entropy tasks [12], although this can be mitigated with data accumulation [2].
 In image generation, several works [3, 10, 14, 15, 16, 17] note image degradation when diffusion models are recursively trained with self-generated data.
-We conduct extensive empirical experiments to identify the most significant factor causing model collapse in text-to-image diffusion models.
+We conduct extensive experiments to reveal the key factor causing model collapse in text-to-image diffusion models.
 Our findings reveal that while model collapse is universally observed across various datasets and scenarios, it manifests in three distinct types of image degradation.
+Theoretical studies on model collapse largely focused on diversity reduction, typically framed as either decreasing covariance in continuous domains [3, 16, 23] or shrinking support in discrete domains [9, 11], which aligns with our observations on decreased recall.
+The finite number of generated samples has been identified as a key cause of model collapse [3, 8, 11, 23], while others [9, 16, 24] emphasize that sampling bias reduces the effective distribution.
+There also exists the study about recursive stability of generative models in the self-consuming training loops [25].
+In contrast, we present a novel framework based on quantitative trait modeling, providing a fresh direction based on mean drift induced by the selection process.
+We further demonstrate that many existing theories of model collapse can be understood through extensions of classical statistical genetics.
+Most existing works echo the importance of incorporating a large proportion of real data [3, 8, 16, 24, 25] or accumulating data over iterations [2] to mitigate model collapse.
+[4] instead suggests a self-correcting self-consuming loop using an expert model, a physics simulator for human motion generation, to correct synthetic outputs.
+Several concurrent works seek to address model collapse through their theoretical insights.
+[26] studies the effects of synthetic data in scaling laws, showing how the scales of models affect model collapse behaviors.
+Moreover, [27] and [28] emphasize the importance of data verification in preventing performance degradation, while [29] introduces a novel token editing strategy for data curation in LLM tasks.
+On the other hand, [30] and [31] propose model-specific solutions leveraging real data to counteract the negative effects of synthetic data in image-generative models.
+In our work, we propose an alternative solution through reusable image generation, and show that it can generate high-quality images comparable to the optimal hyperparameter without extensive search.
+3 Model Collapse in Self-Consuming Chain of Diffusion Finetuning
 (a) Self-Consuming Chain of Diffusion Finetuning.
@@
 Chain of diffusion begins with the original real dataset $D_{0}$.
-(b) Image degradation universally occurs during Chain of Diffusion across four datasets.
-As the Chain of Diffusion progresses, the severity of image degradation intensifies, exhibiting universal patterns of highly saturated images containing repetitive high-frequency patterns.
-This consistently holds across four datasets and $10$ scenarios that we comprehensively investigate in Section 3.2.
-Theoretical studies on model collapse have largely focused on the reduction of diversity, typically framed as either decreasing covariance in continuous domains [16, 6, 3] or shrinking support in discrete domains [11, 9].
-The finite number of generated samples per iteration has been identified as a key cause of model collapse by several authors [3, 6, 22, 11, 8], while others [16, 23, 9] emphasize that sampling bias reduces the generative model’s effective distribution.
-In contrast, we present a novel theoretical framework based on quantitative trait modeling, shifting the focus from variance reduction to mean drift induced by the selection process.
-We further demonstrate that many existing theories of model collapse can be understood as extensions of classical results from statistical genetics.
-Regarding mitigation strategies for model collapse, existing works echo the importance of incorporating a large proportion of real data throughout the training loop [16, 3, 8, 23] or accumulating additional data [2].
-The only mitigation strategy beyond altering the training data composition is proposed by [4], who suggests a self-correcting self-consuming loop using an expert model to correct synthetic outputs.
-While this approach is demonstrated in human motion generation with a physics simulator, having an expert model may not be feasible and could be too costly for many real-world applications.
-In our work, we propose an alternative solution through reusable image generation, and show that it can mitigate model collapse more effectively than mixing more real data.
-3 Model Collapse in Self-Consuming Chain of Diffusion
+(b) Image degradation universally occurs across multiple domains, and low and high CFG scales lead to blurry images and high-frequency degradation, respectively.
+As the Chain of Diffusion progresses, the severity of image degradation intensifies, which holds consistently across four datasets and $11$ scenarios in Section 3.2.
+CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down.
+More images in Appendix C.
 3.1 Problem setting & Experimental Setup
@@
 Given a pretrained generative model $M_{0}$, an original training image set $D_{0}=\{x_{0,i}|i\in[0,N-1]\}$, and a prompt set $P=\{y_{i}|i\in[0,N-1]\}$, where $N$ is the number of total images in the dataset, each image $x_{0,i}$ is paired with a corresponding text prompt $y_{i}$.
-$M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a fully synthetic loop [16] for finetuning a pretrained generative model.
+$M_{k+1}$ is a model finetuned from $M_{0}$ using the generated image set $D_{k}=\{x_{k,i}|i\in[0,N-1]\}$ and the prompt set $P$, which simulates a fully synthetic loop [16].
 Then, $M_{k+1}$ generates a set of images $D_{k+1}$ for the next iteration using the prompt set $P$:
@@
 $D_{k+1}=\text{Generate}(M_{k+1},P).$ (2)
-During the Chain of Diffusion, $M_{0}$ and $P$ are fixed across all the iterations.
-To maintain the same size of training dataset, we generate one image per text prompt for all iterations.
-The overall pipeline of the Chain of Diffusion is shown in Figure 11(a).
-Figure 2: Left: Description of $10$ potential factors (excluding CFG) that we examined as candidate sources for model collapse.
+During the Chain of Diffusion, $M_{0}$ and $P$ are fixed across iterations, and one image is generated per one prompt to maintain the dataset size for all iterations.
+The overall pipeline is shown in Figure 1(a).
+We note that this setting is consistent with [6, 8, 12, 13, 16].
+Model and datasets.
+We use Stable Diffusion v1.5 [32] as the pretrained model $M_{0}$ and apply LoRA [33] to finetune $M_{0}$ at each iteration.
+We build our implementation on [34] and experiment on four datasets: Pokemon [35], Kumapi [36], Butterfly [37], and CelebA-1k [38] to investigate various domains including animation, handwriting, and real pictures.
+$M_{0}$ is finetuned for 100 epochs during each iteration.
+More details can be found in Appendix B.1 and B.2.
+Evaluation metrics.
+We use Frechet Inception Distance (FID) [39] to measure image quality.
+Following [22], we use DiNOv2 [40] as a feature extractor since it is more consistent with our visual inspection than Inception-V3 network [41].
+With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 as the FID score.
+Additional experiments with other metrics (CLIP score [42], Sample-wise Feature Distance (SFD), and recall) are presented in Appendix F.5.
+We propose a new metric to quantify the degradation of generated images. We define collapse rate as the performance degeneration per iteration in the Chain of Diffusion:
+collapse $\text{collapse rate}=\Delta{\rm FID}=\frac{{\rm FID}_{K}-{\rm FID}_{1}}{K-1},$ (3)
+where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set.
+We use FID as a performance metric here, but this can be any other performance metric of interest.
+Note that a low collapse rate indicates more reusable images since the model does not degrade much.
+We have $K=6$ for the rest of the paper.
+Figure 2: Left: Description of $11$ potential factors (excluding CFG) that we examined as candidate sources for model collapse.
 All experiments were conducted on Pokemon except for the dataset size.
 For dataset size, we use CelebA since its original dataset is bigger, and we can subsample $500$, $1000$, and $2000$ images.
-For prompt, we concatenate prompts with different lengths.44footnotemark: 4 Right: All hyperparameter settings other than changing CFG show a high collapse rate greater than $100$, i.e., FID score increases by $\sim$2x in 6 iterations, indicating severe image degradation.
-The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the collapse rate defined in Equation (3).
-For both $\text{FID}_{1}$ and collapse rate, lower is better.55footnotemark: 5
-Model and datasets.
-We use Stable Diffusion v1.5 [24] as the pretrained model $M_{0}$ and apply LoRA [25] to finetune the Stable Diffusion at each iteration.
-We build our implementation on [26] and perform experiments on four datasets: Pokemon [27], Kumapi [28], Butterfly [29], and CelebA-1k [30] to investigate various domains including animation, handwriting, and real pictures.
-During each iteration, we finetune the pretrained Stable Diffusion $M_{0}$ for 100 epochs.
-More implementation details can be found in Appendix A.1 and A.2.
-Evaluation metrics.
-We use Frechet Inception Distance (FID) [31] to measure image fidelity.
-Following [32], we use DiNOv2 [33] as a feature extractor for FID since it is more consistent with our visual inspection than Inception-V3 network [34].
-With a slight abuse of terms, we will still refer to the Frechet distance with DiNOv2 features as the FID score.
-Quantitative results for additional metrics including CLIP score [35] can be found in Appendix E.
-In addition to evaluation metrics for generative models presented above, we propose a new metric to quantify the reusability of images generated in the Chain of Diffusion.
-We define collapse rate as the performance degeneration per iteration in the Chain of Diffusion:
-collapse $\text{collapse rate}=\Delta{\rm FID}=\frac{{\rm FID}_{K}-{\rm FID}_{1}}{K-1},$ (3)
-where ${\rm FID}_{k}$ stands for the FID between $k$-th iteration set and the original training set.
-We use FID as a performance metric here, but this can be CLIP, or any other performance metric of interest.
-Note that a low collapse rate indicates more reusable images since it means that the model does not degrade much during the Chain of Diffusion and we have $K=6$ in the rest of the paper.
+For prompt, we concatenate prompts with different lengths.
+Center: All hyperparameter settings other than changing CFG show a high collapse rate greater than 1.0, i.e., FID score increases by $\sim$2x in 6 iterations, indicating severe image degradation.
+The x-axis represents $\text{FID}_{1}$, quantifying the generation performance at the first iteration and the y-axis represents the collapse rate defined in Eqn.
+3.
+For both $\text{FID}_{1}$ and collapse rate, lower is better.
+Right: Quantitative comparison for FID $\downarrow$ (Pokemon).
+CFG $2.5$ achieves the most robust performance.
+CFG $1.0$ degrades from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in Figure 1(b).
 3.2 Model Collapse in the Chain of Diffusion
@@
 Observation 1: Model collapse is universal in the Chain of Diffusion.
-We observe significant image degradation in all four datasets (see Figure 11(b)) in the Chain of Diffusion.
-The quality of generated images begins to clearly deteriorate in the third iteration and drops even more rapidly once the visible degradation emerges, reaching an unrecognizable level in two or three additional iterations.
-We investigated a variety of different scenarios (summarized in Figure 2) to see if this degradation is an anomaly of specific hyperparameter settings or if it is a ubiquitous phenomenon.
-We tested various dataset sizes for $D_{0}$ and $D_{k}$, increasing the size of synthetic datasets $D_{k}$ by generating more than one image per prompt, mixing real images from $D_{0}$ to $D_{k}$, changing the descriptiveness of prompts, freezing U-Net or text encoder during finetuning, and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip).
-We also tested adding a small Gaussian noise in each image in the original set $D_{0}$ to see if having small random perturbations can improve reusability, and investigated if the degradation happens for larger Stable Diffusion.
+We observe significant image degradation in all four datasets in the Chain of Diffusion (see Figure 1(b) and Appendix C).
+The quality begins to deteriorate in the third iteration and drops rapidly once the degradation emerges, reaching an unrecognizable level in two or three additional iterations.
+Quantitative evaluation (FID, CLIP score, SFD, and recall) also indicates this degradation (Appendix F.5).
+We observed significant drops in recall in all experiments, indicating diversity reduction discussed in several recent works.
+We then investigated a variety of different scenarios (summarized in Figure 2) to see if this degradation is an anomaly of specific hyperparameters or a ubiquitous phenomenon.
+We tested various dataset sizes for $D_{0}$ and $D_{k}$, increasing the size of $D_{k}$ by generating more than one image per prompt, mixing real images of $D_{0}$ to $D_{k}$, changing the descriptiveness of prompts [^3] C, W, S, and L for Combine, Waifu, Short, and Long, respectively.
+We concatenate BLIP and Waifu captions as default setting, referred to as Combine.
+More details can be found in Appendix D.4., freezing U-Net or text encoder, and various other hyperparameters (# sampling steps, # epochs, learning rate, and CLIP skip).
+We also added a small Gaussian noise in each image in the original set $D_{0}$ to test if small random perturbations can improve reusability, or merged all four datasets into one to see if the extended domain delays model collapse (Cross-domain data).
 In all settings we tested, image degradation was universally present and very fast.
-We plot the results for Pokemon on Figure 2 where y-axis is collapse rate and x-axis is the ${\rm FID}_{1}$ (better when closer to the origin).
-While adding noise to images and mixing $90\%$ real images to every iteration as proposed by [16, 3, 8, 23, 2] show the lowest collapse rate, they still exhibit significant degradation as shown in quantitative values (FID score has been doubled in $6$ iterations).
-Visual inspection for images is provided in Appendix C.
+We plot the trade-off for Pokemon on Figure 2 where y-axis is collapse rate and x-axis is the ${\rm FID}_{1}$ (better when closer to the origin).
+While adding noise to images and mixing $90\%$ real images to every iteration as proposed by [2, 3, 8, 16, 24] show the lowest collapse rate, they still exhibit significant degradation (FID score has been doubled in $6$ iterations) [^4] We do not display scenarios that change the training dataset size, such as Img/Prompt and Dataset Size, as varying sizes differ FID scales.
+We observe similar image degradation for them, presented in the Appendix D..
+Moreover, this degradation is unavoidable for Stable Diffusion XL (Appendix D.11).
+We further examined different settings from the literature where all previously generated images are accumulated (data accumulation [2], Appendix D.12) or the same model is iteratively finetuned (iterative retraining [3, 15], Appendix F.4), and consistently observed model collapse.
+(a) t-SNE of images.
+(b) Selection mechanism.
+(c) Power spectra of images vs. Mean of phenotypes.
+Figure 3: (a) t-SNE plot visualizes how generated images evolve from the original distribution (black) and shows distinct paths for three CFG scales (Pokemon).
+Different CFG scales and iterations are differentiated with different colors and transparency.
+Arrows indicate how the distributions of generated images move for different CFG scales.
+While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (1.0 and 7.5) deviate fast, indicating image degradation.
+(b) Selection mechanism with two-sided truncation.
+$r_{1}$ and $r_{2}$ ratios of samples are truncated from the left and right tails, and the remaining $r=1-r_{1}-r_{2}$ ratio of samples is selected.
+(c) Power spectra for different CFGs (Pokemon) and quantitative trait modeling results align well for different selection strategies.
+Directional selections with truncation can effectively explain our observations in Section 3: the behaviors of high- and low-frequency degradation and optimal CFG scale.
+Detailed settings are provided in Appendix E.3.
 Observation 2: CFG is the most significant factor that impacts the model collapse.
-Throughout all our experiments, classifier-free guidance (CFG) had the biggest impact in the speed of model collapse.
-CFG scale was first introduced in [36] to modulate the strength between unconditional and conditional scores at each diffusion step as follows:
+Throughout all our experiments, classifier-free guidance (CFG) had the biggest impact on the speed of model collapse.
+CFG scale was first introduced in [43] to modulate the balance between unconditional score and conditional score at each diffusion step as follows:
+Total Score=Unconditional Score
++CFG(Conditional ScoreUnconditional Score)
+Total ScoreUnconditional Score
+CFGConditional ScoreUnconditional Score
 […]
-A higher CFG means that we emphasize the conditional score for the given prompt more, which pushes the generation to align better with the prompt and often leads to higher-fidelity images.
+High CFG emphasizes the conditional score for the given prompt, which pushes the generation to align better with the prompt and often leads to higher-fidelity images.
 On the other hand, lower CFG places less weight on the conditional score and provides more diversity in generated images.
 For those familiar with temperature sampling [44], CFG plays a similar role as temperature, which adjusts the trade-off between fidelity and diversity.
-In Figure 2, we observe that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$ on x-axis), which is expected from our understanding of CFG.
-A more surprising part is that this comes at the cost of a worse collapse rate (an increase on the y-axis).
-Additionally, when the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and increased CFG worsens both ${\rm FID}_{1}$ and collapse rate.
-Similarly, when the CFG scale is too low—below $2.0$—the improvement in the collapse rate plateaus and both ${\rm FID}_{1}$ and collapse rate begin to increase.
-There is an optimal region of CFG values (near $2.5$, specific to Pokemon), where we achieve low collapse rate while maintaining a good quality in the first iteration as well.
-Moreover, Figure 33(c) presents FID for Pokemon to demonstrate how different CFG scales affect the performance during iterations.
-CFG scale $2.5$ achieves the most robust performance for Pokemon across iterations.
-Interestingly, optimal CFG scales differ for different styles: $2.5$ for animated or hand-writing datasets (Pokemon and Kumapi) and $1.5$ for photo datasets (CelebA and Butterfly) as shown in Appendix B.
-(a) Degradations across different CFGs.
-(b) t-SNE of images.
-(c) FID $\downarrow$.
-Figure 3: (a) Low CFG leads to blurry images and high CFG leads to high-frequency degradation in the Chain of Diffusion.
-CFG $2.5$ for Pokemon and $1.5$ for CelebA exhibit an ideal middle ground where both types of degradation slow down.
-More images in Appendix B.
-(b) t-SNE plot visualizes how generated images evolve from the original distribution (black) and shows distinct paths for three CFG scales (Pokemon).
-Different CFG scales and iterations are differentiated with different colors and transparency.
-Arrows indicate how the distributions of generated images move for different CFG scales.
-While CFG 2.5 (blue) stays near the original images (black), high and low CFG scales (7.5 and 1.0) deviate fast, indicating image degradation.
-(c) Quantitative comparison for FID $\downarrow$ (Pokemon).
-CFG $2.5$ achieves the most robust performance.
-CFG $1.0$ degrades from the beginning of the chain while CFG $7.5$ begins to degrade in the third iteration, which aligns with the visual inspection in (a).
+Figure 2 shows that as we increase the CFG scale, the image quality in the first iteration improves (smaller ${\rm FID}_{1}$ on x-axis), which is expected from our understanding of CFG.
+Surprisingly, this comes at the cost of a worse collapse rate (an increase on the y-axis).
+When the CFG scale is as high as $7.5$ or $10.0$, the improvement in ${\rm FID}_{1}$ plateaus, and it worsens both ${\rm FID}_{1}$ and collapse rate.
+Similarly, when the CFG scale is too low—below $2.0$—the improvement in collapse rate plateaus, and both ${\rm FID}_{1}$ and collapse rate begin to increase.
+There is an optimal region of CFG values (near $2.5$, specific to Pokemon), where we achieve a low collapse rate while maintaining low ${\rm FID}_{1}$ as well.
+Figure 2 presents FID for Pokemon to demonstrate how different CFG scales affect performance.
+CFG scale $2.5$ achieves the most robust performance for all metrics.
+Interestingly, optimal CFG scales differ for different styles: $2.5$ for animated or hand-writing datasets (Pokemon and Kumapi) and $1.5$ for photo datasets (CelebA and Butterfly) as shown in Appendix C.
 Observation 3: High CFG scales cause high-frequency degradation and low CFG scales cause low-frequency degradation.
-The CFG scale affects affect not only the speed of model collapse, but also the pattern of model collapse.
-As shown in Figure 33(a), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to as low-frequency degradation.
-On the other hand, for CFG $7.5$ the degradation pattern looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated.
-The t-SNE plot in Figure 33(b) clearly demonstrates that the distribution shift over iterations follows distinct paths for high, low, and medium CFG scales.
-These patterns were consistent in all four datasets and detailed results can be found in Appendix B.
-In Section 4, we detail how different patterns of power spectra in high-frequency regions from different CFG scales can be understood using the framework from genetic biology.
+CFG scale does not only affect the speed of model collapse, but also the pattern of model collapse.
+As shown in Figure 1(b), CFG $1.0$ makes the images progressively more blurry in the Chain of Diffusion, eventually collapsing to images without any structure, which we refer to low-frequency degradation.
+On the other hand, for CFG $7.5$ how images degrade looks completely different: some features start to be emphasized excessively, repetitive patterns begin to appear, and the overall color distribution becomes saturated.
+The t-SNE plot in Figure 3(a) clearly demonstrates distinct paths for high, low, and medium CFG scales over iterations.
+They were consistent in all four datasets (Appendix C).
+In Section 4, we detail how different patterns of high-frequency power spectra from different CFG scales can be understood using the framework from genetic biology.
 Implications of our observations.
-Our extensive investigations reveal that a high CFG of $7.5$, a common choice to generate visually appealing images, significantly increases collapse rate to achieve slightly better $\rm{FID}_{1}$.
+We showed that a high CFG of $7.5$, a common choice to generate visually appealing images, significantly increases collapse rate to achieve slightly better $\rm{FID}_{1}$.
 Sampling for maximizing the perceptual quality was coined as ‘sampling bias’ in [16].
-While they reported a monotonic increase in collapse rate as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not entirely monotonic when we look at a wider range of CFG scales from $1.0$ to $10.0$.
-It shows an intriguing trade-off between perceptual quality and reusability.
-This suggests that developers concerning reusability of images can substantially improve future generations by carefully choosing CFG.
+While they reported a monotonic increase in collapse rate as CFG increased from $1.0$ to $2.0$, we show that the holistic picture is not monotonic when we look CFG scales from $1.0$ to $10.0$.
+It shows an intriguing trade-off between perceptual quality and reusability, suggesting developers to consider reusability of images to achieve substantially improved future generations by carefully choosing CFG.
 4 Quantitative Trait Models for Fully-Synthetic Training Loops
-This section introduces a novel perspective for understanding the model collapse in generative models by drawing parallels between genetic biology and the Chain of Diffusion.
-Distinguishable iterations in the Chain of Diffusion—where each iteration is clearly separable with no duplicated individuals, and the current iteration originates from the previous one—mirror genetic processes involving successive iterations of parents and offspring.
-By applying quantitative trait modeling from statistical genetics, we provide a framework to describe how images evolve across iterations in the Chain of Diffusion.
-We begin by introducing quantitative trait modeling and its underlying mathematical assumptions.
-Leveraging these assumptions, we derive a theorem demonstrating that the mean trait value exhibits linear drift while the variance stabilizes over time.
-We then show that this model can successfully capture the three key behaviors observed in Section 3: high-frequency degradation, low-frequency degradation, and optimal CFG scale performance.
-This suggests that different CFG scales in the Chain of Diffusion can be viewed as varying selection strategies, with power spectra corresponding to quantitative traits.
+This section introduces a novel perspective to understand model collapse in generative models by drawing parallels between genetic biology and the Chain of Diffusion.
+Distinguishable iterations in the Chain of Diffusion—where each iteration is separable with no duplicated individual, and the current iteration originates from the previous one—mirror genetic processes involving successive iterations of parents and offspring.
 We use the term “iteration” instead of “generation” for genetic generation to avoid confusion with image generation.
-(a) Selection mechanism.
-(b) Power spectra of images vs. Mean of phenotypes.
-Figure 4: (a) Selection mechanism with two-sided truncation.
-$r_{1}$ and $r_{2}$ ratios of samples are truncated from the left and right tails, and the remaining $r=1-r_{1}-r_{2}$ ratio of samples is selected.
-(b) Power spectra for different CFGs (Pokemon) and quantitative trait modeling results of different selection strategies.
-Directional selections with truncation can effectively explain our observations in Section 3: the behaviors of high- and low-frequency degradation and optimal CFG scale.
-Detailed settings are provided in Appendix D.2.
+We begin by explaining how “diversity collapse” of language models in self-consuming training loops [9, 11] could be explained by the classic Wright-Fisher model from statistical genetics.
+Then we introduce Quantitative Trait modeling and its underlying mathematical assumptions.
+Based on them, we derive a theorem showing that the mean trait value exhibits linear drift, while the variance stabilizes over time.
+This model can successfully capture the three key behaviors observed in Section 3: high-frequency degradation, low-frequency degradation, and optimal CFG scale.
+It suggests that different CFG scales in the Chain of Diffusion can be viewed as varying selection strategies, with power spectra corresponding to quantitative traits.
+4.1 Connection to Statistical Genetics
+The Wright-Fisher (WF) model [19] describes how allele (genetic variants) frequencies evolve in a population over iterations, assuming no selection or mutation.
+In each iteration, the entire population is replaced, with the next iteration formed through random sampling with replacement—a process known as genetic drift.
+When the population size is finite, the WF model predicts that one allele will eventually reach fixation, while others are lost.
+We provide further explanations in the Appendix E.1.
+Language models (LMs) follow the same basic idea, with binary alleles replaced by a vocabulary of discrete words.
+LMs estimate word frequencies from a training corpus, and in self-consuming training loops, new texts are generated by the current LM and used to train the next.
+This acts analogously to genetic drift in large-scale, but finite, generations.
+Rare words may gradually disappear from the corpus—an effect that is irreversible.
+Recent works on diversity collapse in LMs [9, 11] can be interpreted as fixation by drift, in line with the WF model.
+In contrast, image generation operates in a near-continuous space of pixel combinations (e.g., $(256^{3})^{512\times 512}$ for $512\times 512$ RGB images), where no output is truly lost—unseen combinations can still be generated.
+In this setting, the process is typically modeled with Gaussian distributions [3, 6, 16, 23], and analogs of genetic drift and fixation emerge through the continuous variants of the WF model [45].
+Given its widespread use in genetics, the WF model has been extended to incorporate selection [46, 47] or mutations [48], further enriching the analogy.
+These connections establish a strong conceptual correspondence between statistical genetics and model collapse—one that, to our knowledge, has been largely overlooked.
+Our work highlights this parallel and demonstrates how it can serve as a unifying theoretical perspective.
 4.2 Quantitative Trait Modeling
-Quantitative trait modeling in statistical genetics explores the evolution of quantitative phenotypes (e.g., height, weight, or color), which are determined by multiple genetic and environmental factors in complex ways.
-These phenotypes are typically assumed to follow a Gaussian distribution, with discrete iterations where parents and offspring are distinguishable ($t$- and $t+1$-th iterations are clearly separable).
+Prior works can be viewed through the lens of genetic drift, where the persistence or disappearance of certain word tokens is driven largely by chance.
+In contrast, we adopt a Quantitative Trait (QT) modeling, incorporating a form of natural selection that reflects human preferences, to explain the evolution of feature distributions and the onset of model collapse.
+QT modeling in statistical genetics explores the evolution of quantitative phenotypes (e.g., height, weight, or color), which are influenced by many genetic and environmental factors.
+These traits are typically modeled as Gaussian-distributed variables and evolve across discrete iterations, where parent and offspring populations are distinguishable ($t$- and $t+1$-th iterations are clearly separable).
 Let the distribution of phenotypes at the $t$-th iteration be denoted as $\mathcal{N}(\mu_{t},\sigma_{P,t}^{2})$ where $\mu_{t}$ and $\sigma_{P,t}^{2}$ are the mean and variance of quantitative phenotypes.
@@
 Here, we only consider additive variance, which is a common assumption in the field..
-When selection occurs in each iteration, whether natural (e.g., faster animals surviving predators) or artificial (e.g., breeding livestock for higher milk production), it affects the distribution of the effective population that influences the next iteration.
-We consider directional selection with truncation, as shown in Figure 44(a), where a ratio of $r$ samples is selected by truncating $r_{1}$ from the left and $r_{2}$ from the right side of the distribution.
-Here, $r+r_{1}+r_{2}=1$ and larger phenotype values are preferred when $r_{1}>r_{2}$.
+When selection occurs, whether natural (e.g., faster animals surviving predators) or artificial (e.g., breeding livestock for higher milk production), it affects the distribution of the effective population that influences the next iteration.
+We consider directional selection with truncation as shown in Figure 3(b), where $r$ ratio of samples is selected by truncating $r_{1}$ from the left and $r_{2}$ from the right side of the distribution.
+Here, $r+r_{1}+r_{2}=1$ and larger phenotypes are preferred when $r_{1}>r_{2}$.
 The (narrow-sense) heritability [^6] The heritability is narrow-sense when the genetic variance is restricted to additive variance., which is defined as the proportion of phenotypic variance attributable to additive genetic factors, can also be represented using the Breeder’s Eqn.
@@
 $h_{t}^{2}=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}}=\frac{\sigma_{G,t}^{2}}{\sigma_{G,t}^{2}+\sigma_{E}^{2}}=\frac{\mu_{t+1}-\mu_{t}}{\mu_{t}^{\prime}-\mu_{t}},$ (5)
-where the mean phenotype of the next iteration $\mu_{t+1}$ is represented using the mean phenotype of selected individuals $\mu_{t}^{\prime}$, the mean phenotype of the entire population at current iteration $\mu_{t}$, and heritability $h_{t}^{2}$.
-Additionally, we assume the genetic variance for the next iteration is determined by the variance of selected individuals $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}$.
+where the mean phenotype of the next iteration $\mu_{t+1}$ is represented using the mean phenotype of selected individuals $\mu_{t}^{\prime}$, the mean phenotype of the entire population $\mu_{t}$, and heritability $h_{t}^{2}$.
+We assume the genetic variance for the next iteration is determined by the variance of selected individuals $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}$.
 We prove the behaviors of mean and variance of phenotypes under this setting:
-Theorem 1
+Theorem 4.1.
 Suppose the distributions of phenotypes follow Gaussian distribution and directional selection truncates individuals on both sides with ratios $r_{1}$ and $r_{2}$.
 Then mean of phenotypes asymptotically increases (decreases) by $\frac{c_{1}c_{2}}{\sqrt{1-c_{2}}}\sigma_{E}$ per iteration and the variance converges to $\frac{1}{1-c_{2}}\sigma_{E}^{2}$ when $r_{1}>r_{2}$ ($r_{2}<r_{1}$), where $c_{1}$ and $c_{2}$ are constants that depend on $r_{1}$ and $r_{2}$.
-The proof of Theorem 1 is provided in Appendix D.1.
+The proof of Theorem 4.1 is provided in Appendix E.2.
 4.3 Explaining the Chain of Diffusion with Quantitative Trait Modeling
-Just as phenotypes evolve through complex functions of hidden genotypes across multiple iterations in quantitative trait modeling, various continuous features of images evolve similarly during the Chain of Diffusion.
-A key discovery in our work is that the high-frequency power spectra are crucial phenotypes that evolve over iterations, with the CFG scales acting as different selection mechanisms.
-These power spectra are computed using 2D Fourier transforms, focusing on frequencies above a certain threshold to capture high-frequency components.
-Different CFG scales during generation correspond to different selection strategies: a high CFG selects individuals in the right tail of the distribution, favoring more detailed features with reduced diversity, while a low CFG selects from the left tail.
-Figure 44(b) shows the evolution of power spectra in the Chain of Diffusion (in solid line) and the phenotype mean modeled with Equation  (9) (in dotted lines).
-The initial mean $\mu_{0}$ and genetic deviation $\sigma_{G,0}$ in the simulation are set to match those of the original image set at iteration 00.
-Three different configurations of ratios, $r_{1}$ and $r_{2}$, successfully model the power spectra distribution: CFG $7.5$ is modeled as selecting from upper $5$% of samples, CFG $2.5$ selects $50$% of the samples slightly favoring higher frequency, and CFG $1.0$ results from selecting the lower $30$%.
-We can see that this modeling can capture the evolution of power spectra very accurately over 6 iterations of Chain of Diffuson.
-Details about power spectra computation and simulation parameters are provided in Appendix D.2.2.
-4.3 Statistical Genetics as a Unifying Lens to Model Collapse
-Quantitative trait modeling, a tool borrowed from statistical genetics, serves as a solid theoretical framework for interpreting our experimental results.
-As genetic processes and the self-consuming training of generative models share a lot of similarities, we find that other concepts in mathematical genetics also have close ties to existing theoretical work on model collapse.
-For instance, the traditional Wright-Fisher model [19] describes how traits evolve from one generation to the next in a finite population.
-Due to randomness in the sampling process, the composition of traits within a finite population drifts over time and eventually collapses to a single phenotype.
-The Wright-Fisher model and its continuous variants [40] are exactly equivalent to the collapse behavior modeled with simple categorical or Gaussian distributions in [16, 3, 6, 22].
-Given its widespread use in genetics, many extensions of the Wright-Fisher model remain an active area of research, such as including different mechanisms of selection [41, 42] or mutations [43].
-Beyond the current work and existing theoretical studies on model collapse, we believe that statistical genetics offers a unified perspective to understand model collapse by reducing the complex dynamics of the self-consuming loop of generative model training to a small number of parameters, from which we can gain valuable insights.
+Just as phenotypes evolve through hidden genotypes in QT modeling, image features similarly evolve in the Chain of Diffusion.
+We identify high-frequency power spectra as a crucial phenotype, influenced by CFG scale as a selection mechanism.
+Using 2D Fourier transforms, we analyze high-frequency components above a certain threshold.
+Different CFG scales correspond to different selection strategies: a high CFG selects individuals in the right tail of the distribution, favoring more detailed features with reduced diversity, while a low CFG selects from the left tail.
+Figure 3(c) compares the evolution of power spectra in the Chain of Diffusion (solid line) with the phenotype mean modeled by Theorem 4.1 (dotted lines).
+The simulation parameters, including initial mean $\mu_{0}$ and genetic deviation $\sigma_{G,0}$, are set to match the original image set at iteration 00.
+Three ratio configurations, $r_{1}$ and $r_{2}$, effectively model the power spectra distribution: CFG $7.5$ is modeled as selecting the top $5$%, CFG $2.5$ selects $50$% of the samples slightly favoring higher frequency, and CFG $1.0$ selected the lower $30$%.
+This modeling accurately captures power spectra evolution over six iterations of the Chain of Diffusion.
+More details can be found in Appendix E.3.2.
+4.4 Further Discussions
+Observable features of generated images are influenced by a complex interplay of factors like random noise, text signals, and various hyperparameters.
+The influence of CFG is thus multifaceted—analogous to how a single genotype gives rise to multiple phenotypes (e.g., red hair and freckles).
+Among many possible phenotypes affected by CFG, we identified one with a clear and consistent trend over iterations.
+It shows how the theoretical framework on mean-drift from QT modeling can interpret our empirical results, revealing that key ideas from statistical genetics naturally align with model collapse.
+Beyond the scope of our current findings, we believe that statistical genetics offers a unified perspective about model collapse by reducing complex dynamics of the self-consuming training loops to a few parameters, from which valuable insights can be drawn.
 5 Reusable Image Generation with ReDiFine
-In the previous sections, we have discovered a significant role of CFG in model collapse and that an appropriate choice of CFG can effectively mitigate the collapse.
-However, the optimal CFG value that minimizes the collapse rate is different for each dataset (e.g., CFG 1.5 for CelebA, CFG 2.5 for Pokemon).
-There is no efficient way of finding an optimal CFG other than performing a grid search over a wide range of values and iteratively fine-tuning the diffusion model to evaluate the collapse rate for each configuration.
-In practical scenarios, it is unlikely that end users who are not machine learning experts will go through such a process during finetuning, just to prevent a potential model collapse that can happen in the next generation.
-This raises the question: How can we design a finetuning and generation strategy that is user-friendly and can effectively slow down model collapse?
-To address this question, we again draw inspiration from the evolution process in nature.
-In biology, mutations naturally counteract genetic drift and preserve diversity.
-Furthermore, selection in nature is often a soft process rather than a hard truncation, as illustrated in Figure 44(a).
+In the previous sections, we have discovered a significant role of CFG in model collapse and that a good choice of CFG can mitigate the collapse while preserving the first iteration FID.
+However, the optimal CFG value differs for each dataset (e.g., 1.5 for CelebA, 2.5 for Pokemon), making iterative finetuning for evaluating the collapse rate on each configuration inevitable to find an optimal CFG.
+In practice, it is unlikely that non-expert end users will go through such a process to prevent a potential model collapse.
+This raises the question: How can we design a user-friendly finetuning and generation strategy that can slow down model collapse without CFG tuning?
+Figure 4: ReDiFine significantly improves the baseline without extensive hyperparameter tuning in the Chain of Diffusion.
+It performs robustly across multiple datasets, resolving high-frequency degradation observed in the baseline with the default configuration.
+To address this question, we again draw inspiration from the evolution process in nature where mutations naturally counteract genetic drift and preserve diversity.
+Furthermore, selection in nature is often a soft process rather than a hard truncation illustrated in Figure 3(b).
 This soft selection allows for the inclusion of outliers, thereby maintaining the overall genetic diversity.
-We connect these biological inspirations with two strategies: condition drop finetuning to introduce more randomness and CFG scheduling during generation to transform hard selection to a softer one [^8] We show a modified quantitative trait modeling result with these two modifications in Appendix D.2 and show that it captures the ReDiFine experimental results effectively.
-In this section, we focus on showing the image generation results with ReDiFine.
-We refer the curious readers to the appendix..
-We propose Reusable Diffusion Finetuning (ReDiFine) method that incorporates these two ideas and show that it is highly effective at mitigating model collapse without the need for any finetuning.
+We connect these biological inspirations with two strategies: condition drop finetuning to include more randomness and CFG scheduling during generation to transform hard selection to a softer one [^7] We show a modified quantitative trait modeling result with these two modifications in Appendix E.3 and show that it captures the ReDiFine experimental results effectively..
+We propose Reusable Diffusion Finetuning (ReDiFine) which integrates these two ideas and achieves a collapse rate similar to the optimal CFG, producing reusable images with minimal extra effort.
 […]
+(a) Pokemon.
+(b) CelebA-1k.
+(c) Kumapi.
+(d) Butterfly.
+Figure 5: ReDiFine performs comparably to the optimal CFG scale, outperforming the baseline (CFG=7.5) on collapse-FID trade-off across all four datasets.
+While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low collapse rate and FID at the same time (lower is better).
+Note that the differences in FID are relatively smaller than those in collapse rate, supporting the necessity to evaluate collapse rate in the Chain of Diffusion.
 CFG Scheduling.
@@
 This scheduling approach is consistent with findings by [51], which suggest that different diffusion steps contribute uniquely to the generation process.
-High CFG lets us capture high-level semantic information accurately during the initial diffusion steps while lower CFG in later steps can prevent generating unnecessary high-frequency details.
-Figure 5: ReDiFine effectively mitigates image degradation in the Chain of Diffusion.
-ReDiFine successfully preserves the characteristics and features without further dataset-specific hyperparameter search.
-Artifacts observed in high-frequency degradation do not exist for all four datasets.
-(a) Pokemon.
-(b) CelebA-1k.
-(c) Kumapi.
-(d) Butterfly.
-Figure 6: ReDiFine effectively mitigates collapse-FID trade-off across all four datasets.
-While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low collapse rate and FID at the same time (lower is better).
-Note that the differences in FID are relatively smaller than those in collapse rate, supporting the necessity to evaluate collapse rate in the Chain of Diffusion.
+We note here that ReDiFine adopts exponential decrease of CFG scale for simplicity, but any advanced CFG scheduling methods [52, 53, 54] can be explored in future research.
 ReDeFine Results.
-We present the generated images and quantitative metrics for ReDiFine.
-We use the initial CFG scale $s_{0}=7.5$, diffusion steps $T=30$, decay rate $\alpha=2$, and condition drop probability $0.2$ for all of our experiments, except in the robustness comparison and ablation study.
-While ReDiFine can be further optimized through hyperparameter tuning (e.g., changing condition drop probability), our goal is to demonstrate that ReDiFine robustly mitigates model collapse with these default parameter settings for all datasets.
-In Figure 5, we can clearly see that ReDiFine significantly improves the image quality at later iterations compared to the baseline (Figure 11(b)) for all four datasets.
+We present the generated images and quantitative metrics for ReDiFine, showing its robust performance without hyperparameter tuning in all four datasets.
+We use the default initial CFG scale $s_{0}=7.5$, decay rate $\alpha=2$, and condition drop probability $0.2$, except in the robustness comparison and ablation study.
+Figure 4 shows that ReDiFine generates significantly better images compared to the baselines (CFG=7.5), and performs comparably to the optimal CFG scales.
 In addition to the visual comparison, Figure 5 quantitatively shows the collapse-FID trade-off of ReDiFine.
 In all four datasets, ReDiFine shows substantially lower collapse rate (y-axis), compared to the baseline case (CFG=7.5).
-Furthermore, the performance of ReDiFine is close to the optimal Pareto curve spanned by different CFG scales, achieving similar performance as the optimal CFG values across all four datasets, demonstrating its effectiveness as a universal solution for mitigating model collapse.
-In contrast, using a fixed CFG scale that works well for one dataset often fails on others: CFG $2.5$ is optimal for Pokemon but performs poorly for CelebA-1k and Butterfly, and CFG $1.5$ is optimal for CelebA but performs poorly for Pokemon and Kumapi.
-We also compare ReDiFine with real image mixing, which is suggested as a mitigation strategy in several previous papers [16, 3, 8, 23].
-In Figure 7, we can observe that even infusing the dataset with 90% real images, it shows severe image degradation within six iterations.
-By visual comparison, we can see that Pikachu images in Figure 5 have less degradation than those with real image mixing.
-Furthermore, applying ReDiFine to these mixed datasets, as shown in Figure 7, significantly improves image quality compared to the baseline.
-This demonstrates that ReDiFine is an effective solution for various settings and can be used alongside other mitigation strategies such as real data mixing.
-Figure 7: Comparison of ReDiFine and baseline (CFG scale 7.5) with training sets mixed with original and synthetic images from the previous iteration.
-While mixing real images as much as $90$% still results in model collapse for the baseline (as shown in the first row), the ability of ReDiFine to mitigate model collapse is even more highlighted by mixing real images (second row).
+Furthermore, the performance of ReDiFine is close to the optimal Pareto curve spanned by different CFG scales, achieving similar performance as the optimal CFG values, demonstrating its effectiveness as a universal and user-friendly solution.
+In contrast, using a fixed CFG scale that works well for one dataset often fails on others: CFG $2.5$ is optimal for Pokemon but performs poorly for CelebA-1k and Butterfly, and CFG $1.5$ is optimal for CelebA-1k but performs poorly for Pokemon and Kumapi.
+Further experimental results of ReDiFine for six additional iterations, cross-domain data, and iterative retraining can be found in Appendix F.
 Ablation study & Further analyses.
-We conducted an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine.
-We plot the results of using only condition drop and CFG scheduling in green triangles on Figure 6.
-In Pokemon, using only one of those strategies outperforms ReDiFine, but in all other datasets, using only one strategy shows higher collapse rate than ReDiFine.
-Especially in CelebA, using either one of them showed significantly worse performance than ReDiFine.
-These results suggest that combining both condition drop and CFG scheduling builds robustness to the method, making ReDiFine effective across all tested datasets.
+We conducted an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to ReDiFine.
+As shown in triangles on Figure 5, using only one of those strategies outperforms ReDiFine in Pokemon, but in all other datasets, using only one strategy shows higher collapse rate than ReDiFine.
+Especially in CelebA-1k, using either one of them showed significantly worse performance than ReDiFine.
+It suggests that combining condition drop and CFG scheduling builds robustness to the method, making ReDiFine effective across all tested datasets.
 We conducted further analyses on ReDiFine, examining the distribution of latent features, the evolution of the norm of Diff, the power spectra using 2D Fourier transforms, and forensic fingerprints based on prior studies [55, 56].
 Our analysis shows that ReDiFine effectively preserves the latent distribution and the norms of Diff over six iterations, with forensic fingerprints closely resembling those of the optimal CFG case.
-Detailed results are provided in Appendix G.
-Reusable image generation as a responsible AI practice.
-In this section, we demonstrated that ReDiFine significantly slows the collapse rate without sacrificing image quality, similar to how fair classifiers reduce bias without compromising accuracy [47].
-Adopting ReDiFine is a responsible, environment-conscious practice to prevent polluting the world with images that might look visually appealing, but totally unusable to improve future AI models.
+Detailed results are provided in Appendix H.
 6 Conclusion
 The influx of AI-generated data into the world is inevitable and training sets that consist of synthetic data will be part of the AI development pipeline.
-In this paper, we empirically and theoretically studied the scenario of finetuning a model with its own generated data, where a gradual degradation called “model collapse” happens.
-We (1) identify the most impactful factor through comprehensive empirical investigations, (2) develop a novel theoretical perspective inspired by statistical genetics to explain model collapse, and (3) propose an effective mitigation strategy that generates reusable images for future training.
+In this paper, we empirically and theoretically studied the scenario of finetuning a generative model with its own generated data, where a gradual degradation called “model collapse” happens.
+We (1) identify the most impactful factor through comprehensive empirical investigations, (2) develop a novel theoretical perspective inspired by statistical genetics to explain model collapse, and (3) propose ReDiFine strategy for diffusion finetuning to slow down image quality degradation in model collapse.
 We started this paper with a question: can current AI models learn from their own output and improve themselves?
@@
 While we presented one solution focused on generating reusable data, many open directions remain, such as developing algorithms that can distinguish between real and synthetic data and apply different learning techniques accordingly.
-Supplementary Material
-Appendix A Experimental setup
-A.1 Hyperparameters
+We further discuss the scopes and limitations of this work in Appendix A.
+Appendix A Scope of Work, Limitations and Broader Impacts
+A.1 Scope and limitations
+In this paper, we provide three main contributions: significant observations from extensive experiments, a novel theoretical framework inspired by statistical genetics, and an effective mitigation strategy for model collapse.
+We detail the boundary and scope of our work in the literature to better understand our contributions.
+First, our empirical observations are based on self-consuming training loops and finetuning of Stable Diffusion v1.5.
+Model collapse is a universal phenomenon observed in recent machine learning training, as high-quality generated contents are now available across multiple domains.
+We target diffusion finetuning with LoRA as it is one of the most practical interests in the field, where tons of finetuned models and images are being released to the Internet every day.
+While we already included the effects of mixing real images, iterative retraining, and domain accumulation to extend the setting from fully self-consuming training loops, we remain open to further investigations about full training, different finetuning methods, and architectures for future work.
+Next, our theoretical framework focuses on introducing an interesting analogy between two fields to open a new area for discussion.
+We showed that QT modeling from statistical genetics provides a powerful tool to understand model collapse based on the structural analogy between self-consuming training loops and genetic processes.
+Further extensions based on advanced methods from statistical genetics would be an important future direction in investigating the model collapse of generative models.
+Specifically, finding connections between these advanced theoretical methods with self-consuming training loops augmented with real data or accumulated data would be an exciting future direction to extend this analogy.
+Finally, the proposed ReDiFine is intended to be an efficient solution that can avoid multiple evaluations of the Chain of Diffusion while achieving performance comparable to that of the optimal CFG scale.
+It could perform comparably to the optimal CFG on four different datasets, which is a promising achievement as the first solution specific to diffusion finetuning.
+Investigating a more powerful method specific for diffusion finetuning would be a promising direction for future research, but we aim to introduce a simple method without additional real images or hyperparameter tuning that demonstrates the concept and importance of generating reusable images.
+A.2 Broader Impacts
+Generating reusable images for further training or finetuning holds the potential to significantly enhance the quality, stability, and robustness of future generative models.
+By promoting continuity across iterations, this approach can improve long-term performance and reduce degradation in self-consuming loops.
+However, as generative models become more powerful and self-reinforcing, there is an increasing risk that such advancements could be misused to replicate, mimic, or infringe upon copyrighted assets, licensed material, or the creative work of artists—especially when the boundary between generated and human-made content becomes less clear.
+We strongly emphasize that any use of generated content for further training must be carried out with a deep understanding of the rights and intent associated with the original data.
+Respecting intellectual property, license agreements, and the creative labor of individuals is essential to maintaining ethical standards in machine learning research.
+While our work aims to improve the technical reliability of generative models, we advocate for responsible usage and encourage the community to develop frameworks and norms that protect creators while fostering innovation.
+Appendix B Experimental setup
+B.1 Hyperparameters
 We finetune Stable Diffusion v1.5 [32] using LoRA [33] at each iteration, with ft-MSE [57] as a fixed VAE to project images into latent space.
@@
 Our implementation follows [34] and is built on PyTorch v2.2.2 [58], with torchvision 0.17.2, running on CUDA 12.4 using NVIDIA A-100 and L40S GPUs.
+Each finetuning and generation takes about 2 hours for  1k images.
 All default hyperparameters are listed in Table 1.
@@
 Number of Images per Prompt $1$
-A.2 Datasets
+B.2 Datasets
 We use four image datasets to demonstrate the universal nature of degradation: Pokemon [35], Kumapi [36], Butterfly [37], and CelebA-1k [38], covering animation, handwriting, and real images.
@@
 Butterfly.
-Butterfly [29] includes $832$ images across $8$ species, using BLIP captions (length $50$-$75$ words) and species descriptions.
-Appendix B Different CFG Scales
+Butterfly [37] includes $832$ real picture images across $8$ species, using BLIP captions (length $50$-$75$ words) and species descriptions.
+B.3 Evaluation metrics
+FID
+The Fréchet Inception Distance (FID) [39] is a widely used metric for evaluating the quality of generative models.
+It measures the similarity between the distributions of generated images and real images by comparing their feature representations extracted from a pretrained Inception network.
+Following [22], we use DiNOv2 feature extractor [40] instead of Inception, as it aligns better with human evaluation for high-quality images.
+CLIP score
+The CLIP score is a metric used to evaluate the alignment between an image and a text description, leveraging the pretrained CLIP (Contrastive Language-Image Pre-training) model [42].
+CLIP is designed to encode both images and text into a shared embedding space, where semantically related image-text pairs have high cosine similarity.
+Based on this embedding space, the CLIP score measures the text-alignment capability of text-to-image generative models by computing cosine similarities of generated images and text descriptions used for generations.
+SFD
+We suggest and compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the same text prompt as the fidelity metric applicable for text-to-image generation.
+It computes text-prompt-wise similarities of the real image and generated image, more directly comparing two pairs of images while other metrics compares the overall distribution of datasets.
+We calculate
+${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (6)
+to evaluate each iteration $k$. SFD overcomes the problem of FID being sensitive to the number of images to compare.
+Recall
+As a key metric for evaluating the diversity of generative models, recall measures the fraction of the real data distribution that is captured by the generated data distribution.
+Following the method proposed by [61], we utilize the DiNOv2 feature extractor and set the number of neighbors to 5 for our evaluations.
+Appendix C Different CFG Scales
 This section presents images from Chain of Diffusion at various CFG scales for the four datasets.
-Figure 8, 9, 10, and 11 corresponds to Pokemon, CelebA-1k, Kumapi, and Butterfly, respectively.
+Figure 6, 7, 8, and 9 corresponds to Pokemon, CelebA-1k, Kumapi, and Butterfly, respectively.
 Each dataset shows results for five CFG scales, covering high, medium, and low values.
@@
 Figure 9: Chain of Diffusion for Butterfly at various CFG scales. The optimal CFG scale is $1.5$.
-Appendix C Hyperparameter investigations to unveil the most significant factor of degradation
+Appendix D Hyperparameter investigations to unveil the most significant factor of degradation
 This section presents experimental results identifying the most significant factors contributing to degradation in the Chain of Diffusion.
@@
 The CFG scale is fixed at $7.5$ for all cases.
-C.1 Training set size
+D.1 Training set size
 We used CelebA dataset [38] to examine how training set size (both $D_{0}$ and $D_{k}$) impacts degradation in the Chain of Diffusion.
@@
 Figure 10: Chain of Diffusion on CelebA dataset with varying training set sizes. Degradation occurs faster with smaller sets, but all result in severe degradation by the $6$th iteration.
-C.2 Number of images per prompt
+D.2 Number of images per prompt
 Generating multiple images per prompt is a simple way to increase training set diversity and can be considered as a solution to mitigate degradation in the Chain of Diffusion.
@@
 Figure 11: Chain of Diffusion on Pokemon dataset with multiple images generated per prompt. Increasing the training set size in this way does not mitigate degradation.
-C.3 Mixing real images to synthetic sets
+D.3 Mixing real images to synthetic sets
 Many previous works suggest augmenting the training set with real images to mitigate degradation during iterative training.
@@
 At each iteration, images are randomly replaced with corresponding real images.
-Figure 14 and 15 show how degradation in the Chain of Diffusion varies when $50$% and $90$% of images are replaced for Pokemon and CelebA-1k datasets, respectively.
-Notably, even $10$% synthetic images are sufficient to induce degradation, and $50$% replacement rarely slows it down.
+Figure 12 and 13 show how degradation in the Chain of Diffusion varies when $50$%, $90$%, $95$%, $98$% and $99$% of images are replaced for Pokemon and CelebA-1k datasets, respectively.
+Notably, even $5$% synthetic images are sufficient to induce degradation, and $50$% replacement rarely slows it down.
 CelebA-1k dataset appears to be significantly more susceptible to degradation.
+Importantly, using only $5\%$ and $2\%$ synthetic images is sufficient to trigger model collapse in Pokemon and CelebA-1k datasets.
+This demonstrates the catastrophic impact of synthetic data on diffusion finetuning and raises the question of whether synthetic data detection can be a safe solution.
 Figure 12: Chain of Diffusion on Pokemon dataset with real images randomly replacing synthetic images at each iteration.
@@
 Figure 13: Chain of Diffusion on CelebA-1k dataset with real images randomly replacing synthetic images at each iteration.
-A $50$% replacement rarely slows degradation, while $10$% synthetic images are sufficient to initiate it.
+A $50$% replacement rarely slows degradation, while $5$% synthetic images are sufficient to initiate it.
 It suffers from more severe degradation than Pokemon dataset as compared with Figure 12.
-C.4 Prompt Set
+D.4 Prompt Set
 We hypothesized that the descriptiveness of prompts influences degradation in the Chain of Diffusion.
@@
 We compare the Chain of Diffusion using longer prompts (over $50$ words) and shorter prompts (under $25$ words).
-C.5 U-Net and Text-Encoder
+D.5 U-Net and Text-Encoder
 Figure 16 illustrates the Chain of Diffusion with either the U-Net or text encoder finetuned.
@@
 Figure 16: Chain of Diffusion on Pokemon dataset with either the U-Net or text encoder finetuned.
-C.6 Number of diffusion steps
+D.6 Number of diffusion steps
 We investigated whether an insufficient number of diffusion steps during generation contributes to degradation.
 Figure 17 shows that increasing the number of diffusion steps does not enhance the Chain of Diffusion.
-C.7 Number of Epochs
+D.7 Number of Epochs
 We also examined whether insufficient or excessive training affects our default setting.
@@
 Both increased diffusion steps and differing training epochs fail to mitigate degradation, resulting in similar patterns.
-C.8 Learning Rate
+D.8 Learning Rate
 Similarly, we assessed finetuning adequacy in Figure 18 by adjusting the learning rates for the U-Net and text encoder by x10101010 and x0.10.10.10.1.
@@
 Figure 18: Chain of Diffusion on Pokemon dataset with varying learning rates and added Gaussian noise to the original training set.
-C.9 CLIP Skip
+D.9 CLIP Skip
 We investigated the impact of the CLIP skip hyperparameter on degradation.
@@
 Figure 19: Chain of Diffusion on Pokemon dataset using different CLIP skip hyperparameters. The CLIP skip hyperparameter shows a negligible effect on degradation.
-C.10 Adding Gaussian noise to the original training set
+D.10 Adding Gaussian noise to the original training set
 We examined whether differences between real and synthetic images contribute to degradation by adding random Gaussian noise to the original training set $D_{0}$.
@@
 This supports our findings that degradations are universal across real, animation, and handwritten images.
-C.11 Stable Diffusion XL
+D.11 Stable Diffusion XL
 We investigated image degradation for a different Stable Diffusion model, noting that the optimal hyperparameters for finetuning SDXL using LoRA are not well established.
@@
 Results are presented in Figure 20.
-Figure 22: Chain of Diffusion of SDXL on Pokemon dataset.
+Figure 20: Chain of Diffusion of SDXL on four datasets.
 Due to insufficient investigation into optimal hyperparameters for finetuning SDXL, our experiments largely rely on those from Stable Diffusion v1.5.
-Appendix D Quantitative Trait Modeling
-D.1 Proof of Theorem 1
+D.12 Data Accumulation
+Data accumulation experiments aim to investigate whether concept overfitting and disappearing are major reasons for model collapse.
+The training set for iteration $t$ is the combination of all previously generated sets, including the original training set.
+The number of training epochs is controlled accordingly to maintain the total number of updates.
+Figure 21: Chain of Diffusion on Pokemon dataset when training set is accumulated from previous iterations. All concepts from previous iterations are preserved for finetuning.
+Figure 22: Chain of Diffusion on CelebA-1k dataset when training set is accumulated from previous iterations. All concepts from previous iterations are preserved for finetuning.
+Appendix E Quantitative Trait Modeling
+E.1 Wright-Fisher model - Genetic drift and fixation
+The Wright-Fisher model describes the evolution of discrete allele frequencies in a finite population.
+We assume distinguishable iterations (no overlapping individuals) and no selection or mutation.
+In each iteration, all individuals from the previous iteration (the parents) are replaced by new individuals (the offspring).
+The offspring are sampled from the parental population via random sampling with replacement; for simplicity, we assume a haploid model.
+Consider a population of constant size $N$ and two alleles, $A$ and $a$.
+If allele $A$ is present in $i$ individuals in the parent interaction, the probability that it appears in $j$ individuals in the offspring iteration is given by the binomial distribution:
+$P_{ij}=\binom{N}{j}({i\over N})^{j}(1-{i\over N})^{N-j},0\leq i,j\leq N.$ (7)
+This process defines a Markov chain, where the state in each iteration depends only on the state in the previous one.
+The Wright-Fisher model thus characterizes the stochastic dynamics of allele frequencies across iterations.
+Certain states in a Markov chain are absorbing—that is, once entered, they cannot be exited.
+In the Wright-Fisher model, the fixation or loss of an allele corresponds to such absorbing states.
+Once the population becomes entirely composed of a single allele, the other allele cannot reappear under the assumptions of no mutation or migration.
+Importantly, in a finite population, the process is guaranteed to eventually reach one of these two absorbing states.The probability that allele $A$ becomes fixed depends solely on its initial frequency in the population.
+E.2 Proof of Theorem 4.1
 Proof. The difference between the successive means in quantitative trait modeling is given by:
@@
 $\square$
-D.2 Simulation setup
+E.3 Simulation setup
 Our simulation aims to demonstrate that our experimental results can be modeled using quantitative trait modeling.
 We show that the radial sum of power spectra of images is one of the phenotypes explained by our theoretical analysis.
-D.2.1 Computing the radial sum of power spectra.
+E.3.1 Computing the radial sum of power spectra.
 The power spectra of images are computed as the square of the magnitude of the 2D Fourier transform.
@@
 Our modifications to heritability and selection process successfully demonstrate changes occur to images by ReDiFine.
-D.2.2 Simulation parameters matching different CFG scales.
+E.3.2 Simulation parameters matching different CFG scales.
 We simulate different selection strategies using truncation ratios $r_{1}$ and $r_{2}$.
@@
 We set the environmental standard deviation to $0.25$.
-D.2.3 ReDiFine and mutations.
-ReDiFine combines condition drop finetuning and CFG scheduling, inspired by the mutation mechanism that compensates for distribution shift and maintains genetic distributions in population genetics.
+E.3.3 ReDiFine and mutations.
+ReDiFine combines condition drop finetuning and CFG scheduling, drawing inspiration from genetic mutation mechanisms that counter distributional shifts and help maintain genetic diversity in population genetics.
+In the Chain of Diffusion, training and generation can be viewed analogously to reproduction and natural survival.
+ReDiFine modifies both stages to preserve image quality across iterations.
+The two core components—condition drop finetuning and CFG scheduling—are rooted in biological analogies.
+Condition drop finetuning, akin to mutation, introduces controlled variability to ensure that the offspring distribution (the next iteration) remains aligned with the parent distribution (the previous iteration).
+This mitigates over-specialization (generated images becoming too uniform or isolated) and prevents degeneration under strong selection pressures.
+CFG scheduling, inspired by natural selection, dynamically adjusts the CFG scale during generation, promoting diversity by allowing broader phenotypic expression while retaining directed evolution toward desired traits.
+By integrating these biologically inspired strategies, ReDiFine improves the stability of the Chain of Diffusion and prevents catastrophic degradation across iterations.
 We apply two modifications to our theoretical analysis of Section 4—adding mutation variance to heritability and smoothing truncations—to simulate the effects of ReDiFine in the Chain of Diffusion.
@@
 This suggests further research on model collapse motivated from other fields like biology.
-Appendix E ReDiFine
+Appendix F ReDiFine
+F.1 Visual Inspections
 Figure 24, 25, 26, and 27 show how robust ReDiFine is to different CFG scales. It successfully mitigates the high-frequency degradations for a wide range of CFG scales.
@@
 Figure 27: Chain of Diffusion of ReDiFine with different CFG scales on Butterfly dataset. ReDiFine successfully achieves robust image qualities for varying CFG scales.
-In addition to visual inspections for images generated using ReDiFine, we compare the quantitative results of ReDiFine to baselines with different CFG scales across FID and CLIP score.
-Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
-${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (17)
-to evaluate each iteration $k$.
-SFD overcomes the problem of FID being sensitive to the number of images to compare.
-DiNOv2 features are used to compute SFD.
+F.2 Further iterations
+We conduct additional experiments to compare the baseline with the optimal CFG scale and ReDiFine over extended iterations.
+As shown in Figure 28, ReDiFine consistently generates images of similar quality up to 12 iterations, whereas the optimally tuned CFG scale fails to sustain image quality.
+This decline suggests that repeated hyperparameter searches are necessary to identify suitable CFG scales for subsequent iterations.
+Such an approach becomes increasingly impractical as the number of iterations grows, highlighting the limitations of relying on the optimal CFG scale to mitigate model collapse.
+Figure 28: Chain of Diffusion of baseline with the optimal CFG scale and ReDiFine for more iterations.
+The generation quality of ReDiFine is preserved for additional iterations while optimally found CFG scale $2.5$ fails to maintain the image qualities.
+Similar high-frequency degradation is observed.
+F.3 Cross-domain Data
+We additionally conduct experiments when the training dataset is the cross-domain set of four datasets (Pokemon, CelebA-1k, Kumapi, and Butterfly) to investigate whether having a broader range of concepts impact model collapse.
+The merged dataset serves as the original training set, while the combined captions are used for image generation at each iteration.
+The results, presented in Figure 29, show that despite the increased number of images and the inclusion of diverse concepts and domains, model collapse persists at both low and high CFG scales.
+Moreover, no single CFG scale (e.g., $1.5$ or $2.5$) can consistently produce high-quality, reusable images across all datasets, highlighting the limitations of relying on an optimal CFG scale for diverse domains.
+In contrast, ReDiFine leverages the increased conceptual diversity in the original training set, generating more reliable images across all datasets.
+To ensure a fair comparison, we control the number of training epochs to maintain consistent updates across experiments.
+Figure 29: Chain of Diffusion with the multi-domain set of datasets (Pokemon, CelebA-1k, Kumapi, and Butterfly) used for finetuning.
+Model collapse remains evident at low and high CFG scales.
+A single CFG scale (e.g., $1.5$ or $2.5$) fails to achieve optimal performance across both the Pokemon and CelebA-1k datasets.
+In contrast, ReDiFine successfully generates high-quality, reusable images simultaneously.
+While images for Kumapi and Butterfly datasets are not displayed due to space constraints, they are included in the finetuning process along with the other datasets.
+F.4 Iterative Retraining
+Some prior works on model degeneration examine scenarios in which a single model is continually trained on synthetic data it has generated.
+To adapt our Chain of Diffusion framework to this setting, we consider a setup where the same model is finetuned iteratively across multiple iterations.
+At each iteration, the model generates a fixed number of images using a predefined prompt set, and these generated images are then used to further finetune the model.
+Figure 30 demonstrates how images degrade under this setting across different CFG scales and ReDiFine.
+While severe degradation is observed for low and high CFG scales, the optimal CFG scale and ReDiFine are able to mitigate model collapse, generating high-quality images.
+This indicates that the effect of ReDiFine is maintained even when a single model is continually finetuned.
+Figure 30: Chain of Diffusion where a single model is continually finetuned across multiple iterations.
+Model collapse is observed consistently at CFG scales of 1.0 and 7.5, while the baseline with a CFG scale of 2.5 and ReDiFine effectively mitigate model collapse.
+This demonstrates that model collapse is a universal phenomenon across different settings, and the effect of ReDiFine is effective robustly.
+F.5 Quantitative results
+In addition to visual inspections for images generated using ReDiFine, we compare the quantitative results of ReDiFine to baselines with different CFG scales across FID, CLIP score, SFD, and recall.
+Details about each metric are provided in Appendix B.3.
+DiNOv2 features are used to compute FID, recall, and SFD.
+We follow [61] to compute recall and set the number of neighbors for computing recall $5$.
 The results, shown in Figure 31, demonstrate that ReDiFine achieves performance comparable to the optimal CFG scales ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly) across different datasets and metrics.
@@
 (l) Butterfly, SFD.
+(m) Pokemon, recall.
+(n) CelebA-1k, recall.
+(o) Kumapi, recall.
+(p) Butterfly, recall.
 Figure 31: Quantitative results of ReDiFine and baselines (different CFG scales).
-Appendix F Ablation study
+Appendix G Ablation study
 This section provides an ablation study to understand how condition drop finetuning and CFG scheduling contribute to the success of ReDiFine.
-F.1 Condition drop finetuning
+G.1 Condition drop finetuning
 We conducted an ablation study to understand how the probability of dropping text embedding during finetuning affects the image quality in the Chain of Diffusion.
@@
 Figure 33: Chain of Diffusion with condition drop finetuning on CelebA-1k dataset.
-F.2 CFG scheduling
+G.2 CFG scheduling
 We also evaluated how different CFG scale decreasing strategies impact image degradation in the Chain of Diffusion.
@@
 Figure 35: Chain of Diffusion with CFG scheduling on CelebA-1k dataset.
-Appendix G Analysis
+Appendix H Analysis
 In this section, we present a series of analyses of images generated through the Chain of Diffusion.
@@
 Especially, high differences in the later diffusion steps can be a cause of high-frequency degradation.
-G.1 Latent analysis
-Figure $3$(a) illustrates how the distribution of latent values evolves across different iterations.
+H.1 Latent analysis
+Figure 36(a) illustrates how the distribution of latent values evolves across different iterations.
 The histograms show the final latent vectors before decoding into pixel space, comparing various CFG scales and ReDiFine.
@@
 ReDiFine further enhances this preservation, maintaining the histogram from the first to the last iteration, thus achieving both high fidelity in the first iteration and better reusability.
-G.2 Differences between conditional and unconditional scores
+H.2 Differences between conditional and unconditional scores
 […]
-G.3 Power spectra of 2D Fourier transforms
+H.3 Power spectra of 2D Fourier transforms
 Figure 37 demonstrates the radial and angular spectrum power density of both the original and synthetic images.
@@
 Power density of generated images by ReDiFine remains during the iterations while baselines cannot maintain angular distribution.
-G.4 Fingerprints for forensic analysis
+H.4 Fingerprints for forensic analysis
 (a) Autocorrelation.